### v1

# Large deviations for the smallest eigenvalue of a deformed GOE with an outlier

Abstract.
We establish a large deviation principle for the smallest eigenvalue of a random matrix model composed of the sum of a GOE matrix and a diagonal matrix with an outlier. Our result generalizes and unifies previously studied cases.
1. Introduction
1.1. Setting of the problem
In this paper, we study the asymptotic behavior of the smallest eigenvalue of a sequence of random matrices $\{X_{N}\}_{N\in\mathbb{N}}$ where $X_{N}$ is given by the sum of two $N\times N$ matrices: a matrix $G_{N}$ taken from the Gaussian Orthogonal Ensemble (GOE) of variance $t$, and a deterministic matrix $B_{N}$ whose empirical eigenvalue distribution converges weakly towards a probability measure $\nu$. We also allow the deterministic matrix to have an outlier $\Lambda$ below the support of $\nu$.
In this setting, the typical behavior of the spectrum of $X_{N}=G_{N}+B_{N}$ is well known. The empirical measure of the eigenvalues of $X_{N}$ converges towards the free convolution $\sigma_{t}\boxplus\nu$ where $\sigma_{t}$ is the semicircular distribution with variance $t$[VDN92, BV93]. This probability measure was described in detail in [Bia97]. If we denote by $G_{\mu}$ the Stieltjes transform of a probability measure $\mu$ on $\mathbb{R}$ given, for $z$ outside the support of $\mu$, by
$G_{\mu}(z)=\int\frac{\mathrm{d}\mu(x)}{z-x}\,,$
then, the Stieltjes transform $G_{\nu\boxplus\sigma_{t}}$ of $\sigma_{t}\boxplus\nu$ is defined as follows: Consider the mapping
$H_{\nu,t}:\left|\begin{array}[]{rcl}\overline{\Omega_{\nu,t}}&\longrightarrow&\mathbb{C}^{+}\cup\mathbb{R}\\ z&\longmapsto&z+tG_{\nu}(z)\\ \end{array},\right.$
with
$\Omega_{\nu,t}:=\{u+iv\in\mathbb{C}^{+}:v>v_{t}(u)\}$ (1.1)
where for all $u\in\mathbb{R}$,
$v_{t}(u):=\inf\Bigr{\{}v\in\mathbb{R}^{+}:\int\frac{\mathrm{d}\mu(\lambda)}{|\lambda-u|^{2}+v^{2}}\leq\frac{1}{t}\Bigr{\}}.$ (1.2)
Then, $H_{\nu,t}$ is invertible with inverse $\omega_{\nu,t}$ from $\mathbb{C}^{+}$ to $\overline{\Omega_{\nu,t}}$ and moreover
$G_{\nu\boxplus\sigma_{t}}(z)=G_{\nu}(\omega_{\nu,t}(z))\,.$
We denote by $\ell_{\nu,t}$ (resp. $\ell_{\nu}$) the left boundary point of the support of $\nu\boxplus\sigma_{t}$ (resp. of $\nu$). Note that $H_{\nu,t}$ is increasing on $({-\infty},\omega_{\nu,t}(\ell_{\nu,t}))$ and decreasing on $(\omega_{\nu,t}(\ell_{\nu,t}),\ell_{\nu})$, so that the equation $H_{\nu,t}(x)=\lambda$ has at most two solutions smaller than $\ell_{\nu}$, with the smallest being $\omega_{\nu,t}(\lambda)$.
When $-\Lambda$ is large enough, it is well-known that the smallest eigenvalue $\lambda_{1}(X_{N})$ of $X_{N}=G_{N}+B_{N}$ escapes the bulk, becoming smaller than $\ell_{\nu,t}$, a phenomenon called the Ben-Arous-Baik-Peche (BBP) transition. Throughout the paper, we denote
$\rho_{\nu,t}^{\Lambda}:=H_{\nu,t}(\Lambda)=\Lambda+tG_{\nu}(\Lambda)$
and
$\ell_{\nu,t}^{\Lambda}:=\begin{cases}\rho_{\nu,t}^{\Lambda}&\text{if $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$}\\ \ell_{\nu,t}&\text{if $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$}.\end{cases}$
The following convergence result was proved in [CDMFF11]:
Theorem 1.1.
The smallest eigenvalue of $X_{N}=G_{N}+B_{N}$ converges almost surely towards $\ell_{\nu,t}^{\Lambda}$ when $N$ goes to infinity. Moreover the map $\Lambda\in(-\infty,\ell_{\nu})\mapsto\ell_{\nu,t}^{\Lambda}$ and $\nu\mapsto\ell_{\nu,t}^{\Lambda}$ are continuous (if $\mathcal{P}(\mathbb{R})$ is endowed with the weak topology).
Our aim is to study the large deviations for the law of $\lambda_{1}(X_{N})$. Such a question was addressed by Ben McKenna [McK21] in the case where $B_{N}$ has no outlier, namely when $\Lambda=\ell_{\nu}$. The strategy used in [McK21] is based on a tilt by spherical integrals introduced in [GH20] and subsequently developed in diverse contexts in [GH22, CDG23, DHG24, GM20, AGH21]. However, this strategy can only be used to study large deviations below the limiting outlier $\rho_{\nu,t}^{\Lambda}$. On the other hand, [Mai07] studied the case where $\nu=\delta_{0}$, namely when $B_{N}$ has only one non trivial eigenvalue $\Lambda$. Large deviations were then obtained thanks to an exact estimation of the density due to the asymptotics of spherical integrals derived in [GM05]. In this paper, we prove a large deviation principle for $\lambda_{1}(X_{N})$, unifying the results of [Mai07] (where $\nu=0$) and [McK21] (where no outlier is permitted, namely $\Lambda=\ell_{\nu}$). Our approach is new and consists of estimating the density of the law of the smallest eigenvalue by exhibiting a fixed point functional equation for the large deviation rate function.
Our study is motivated by a problem arising in the theory of spin glasses. It is well known in the physics literature that the equilibrium properties of the $p$-spin glass model are related to the topology of the energy landscape of a random function called the Thouless-Anderson-Palmer (TAP) free energy [TAP77]. In the case of the pure Ising $p$-spin glass, the Hessian of the TAP free energy contains a projector term, which may produce a negative eigenvalue at low enough temperature [Ple99, BM80, PR04]. Therefore, computing the (annealed) number of local minima requires studying the large deviations of a random matrix of the form $G_{N}+D_{N}-\Lambda uu^{T}$, where $G_{N}$ is a GOE matrix, $D_{N}$ is a diagonal matrix, $\Lambda\in\mathbb{R}$, and $u\in\mathbb{R}^{N}$. This corresponds to the model considered in this paper.
Several other works have studied large deviation principles (LDPs) for the extreme eigenvalues of GOE matrices. The article [Mai07], which examined a rank-one deformation of the GOE was extended in [BGGM12] to finite rank deformations, and in [BG20], which proved an LDP for the joint distribution of extreme eigenvalues and their associated eigenvectors. By leveraging the computation of spherical integrals, [GH20] and [GM20] investigated the extreme eigenvalues of sharp subgaussian matrices and free convolution models. In [McK21], an LDP was obtained for $G_{N}+B_{N}$ where $B_{N}$ is a diagonal matrix without outliers with a rate function expressed variationally.
The smallest eigenvalue of Wigner matrices is highly sensitive to the entry distribution’s tail behavior so that even the speed of its large deviation principle depends on this tail. For distributions with tails heavier than Gaussian, an LDP was established in [Aug16], showing the speed and the rate function depend only on the tail decay rate. Sharp sub-Gaussian and general sub-Gaussian distributions have been widely studied. Sharp sub-Gaussian distributions have tails asymptotically equivalent to Gaussian, resulting in a rate function matching the one of the GOE [BC14, AGH21]. In contrast, general sub-Gaussian distributions may exhibit a transition to a non-universal rate function, indicating eigenvector localization phenomena [CDG23]. Recently, an LDP for the largest eigenvalue of random matrices with a variance profile was established, highlighting the impact of the variance profile on the rate function [Hus22, DHG24].
1.2. Main result
The GOE of size $N\times N$ of variance $t>0$ is the symmetric random matrix $(G_{ij})_{1\leq i,j\leq N}$ where for each $i\leq j$ and $k\leq l$ such that $(i,j)\neq(k,l)$, $G_{ij}$ and $G_{kl}$ are independent and such that for each $i\leq j$,
$\mathrm{Var}[G_{ij}]=\frac{t}{N}(1+1_{i=j}).$ (1.3)
We make the following assumptions on the sequence of matrices $(B_{N})_{N\in\mathbb{N}}$:
Assumptions 1.1.
Assume that $B_{N}\in\mathcal{S}_{N}(\mathbb{R})$ has eigenvalues $b_{1}^{N}\leq\cdots\leq b_{N}^{N}$ satisfying the following:
  1. (1)
    There exists a measure $\nu\in\mathcal{P}_{1}(\mathbb{R})$ with support bounded from below such that
    $\lim_{N\to\infty}d\Bigr{(}\frac{1}{N}\sum_{i=1}^{N}\delta_{b_{i}^{N}},\nu\Bigr{)}=0$ (1.4)
    where $d$ is a distance compatible with the weak topology such as the one defined in (1.13)
  2. (2)
    With $\ell_{\nu}$ the left boundary point of the support of $\nu$, we assume there exists $\Lambda\leq\ell_{\nu}$ finite so that
    $\lim_{N\to\infty}b_{1}^{N}=\Lambda<\ell_{\nu}$ (1.5)
  3. (3)
    $\lim_{N\to\infty}b_{2}^{N}=\ell_{\nu}.$ (1.6)
  4. (4)
    $\sup_{N}|b_{N}^{N}|<\infty$ and $\nu$ has finitely many atoms.
We denote by $\eta_{\nu}:[0,1]\mapsto\mathbb{R}$ an increasing function such that $\nu(\mathrm{d}x)=\eta_{\nu}\#1_{[0,1]}\mathrm{d}x$: it is the reciprocal of the partition function $\nu((-\infty,.])$.
Let $\Lambda\leq\ell_{\nu}$ and $\lambda\leq\ell_{\nu,t}$. Throughout the paper, when $\lambda\in[H_{\nu,t}(\ell_{\nu}),\ell_{\nu,t}]$, we denote $\omega_{\nu,t}^{*}(\lambda)$ the unique solution in $[\omega_{\nu,t}(\ell_{\nu,t}),\ell_{\nu})$ of the equation $H_{\nu,t}(\gamma)=\lambda$. Note that this condition is satisfied when $\lambda\geq H_{\nu,t}(\Lambda)$ and $\Lambda\in[\omega_{\nu,t}(\ell_{\nu,t}),\ell_{\nu}]$. Define
$\gamma(\Lambda,\lambda):=\begin{cases}\Lambda&\text{if $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$, or $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda\leq\rho_{\nu,t}^{\Lambda}$}\\ \omega^{*}_{\nu,t}(\lambda)&\text{otherwise}.\end{cases}$ (1.7)
For every probability measure $\nu$ on $\mathbb{R}$, we let $S_{\nu}$ be the logarithmic potential defined by
$S_{\nu}:x\in\mathbb{R}\setminus\mathrm{supp}(\nu)\mapsto-\int\log|x-\lambda|\mathrm{d}\nu(\lambda).$ (1.8)
We set $I_{\nu,t}^{\Lambda}$ to be infinite on $(\ell_{\nu,t},+\infty)$ and otherwise be given by
$I_{\nu,t}^{\Lambda}(\lambda)=\frac{1}{2}(S_{\nu}(\omega_{\nu,t}(\lambda))-S_{\nu}(\gamma(\Lambda,\lambda)))+\frac{1}{4t}((\lambda-\gamma(\Lambda,\lambda))^{2}-(\lambda-\omega_{\nu,t}(\lambda))^{2}).$ (1.9)
Theorem 1.2 (LDP for the smallest eigenvalue of Gaussian matrices).
Let $t>0$. Let $(B_{N})_{N\in\mathbb{N}}$ be a sequence of matrices of size $N\times N$ satisfying Assumptions 1.1. Let $\nu$ and $\Lambda$ be as in a (1.4) and (1.5). Let $G_{N}$ be a sequence of GOE matrices of size $N\times N$ and variance $t$.
Then, the smallest eigenvalue of $G_{N}+B_{N}$ satisfies a large deviation principle at speed $N$ with good rate function $I_{\nu,t}^{\Lambda}$. Moreover $I_{\nu,t}^{\Lambda}$ is strictly convex and attains its minimum value at $\ell_{\nu,t}^{\Lambda}$.
Remark 1.2.
Assumption (4) should not be necessary. Moreover, our approach extends readily to the case where $G_{N}$ is a GUE matrix instead of a GOE matrix.
Remark 1.3.
Our result can be applied when $B_{N}$ has no outlier by taking $\Lambda=\ell_{\nu}$, thus recovering the result of [McK21], and in the case where $\nu=\delta_{0}$, thus recovering the result of [Mai07]. This is detailed in Section 4.
1.3. Notation
  • We denote by $\mathcal{P}(\mathbb{R})$ the set of probability measures on $\mathbb{R}$.
  • For any $\nu\in\mathcal{P}(\mathbb{R})$, we denote by $\mathrm{supp}(\nu)$ the support of $\nu$ and $\ell_{\nu}$ the infimum of $\mathrm{supp}(\nu)$. For $t>0$ we denote $\ell_{\nu,t}:=\ell_{\nu\boxplus\sigma_{t}}$.
  • For any $\nu\in\mathcal{P}(\mathbb{R})$, we let $S_{\nu}$ be the logarithmic potential defined by
    $S_{\nu}:x\in\mathbb{R}\setminus\mathrm{supp}(\nu)\mapsto-\int\log|x-\lambda|\mathrm{d}\nu(\lambda).$ (1.10)
    We let +=$\mathbb{C}^{+}=\{z=x+iy:y>0\}$. Define for $\nu\in\mathcal{P}(\mathbb{R})$ with support $[a,b]$,
    $G_{\nu}:\mathbb{C}^{+}\cup\mathbb{R}\setminus[a,b]:z\mapsto\int\frac{\mathrm{d}\nu(x)}{z-x}.$ (1.11)
    We then let $G_{\nu}(a)=\lim_{y\uparrow a}G_{\nu}(y)$ and $G_{\nu}(b)=\lim_{y\downarrow b}G_{\nu}(y)$. One can show that $G_{\nu}|_{\mathbb{R}}$ is a bijection from $\mathbb{R}\setminus[a,b]$ to $(G_{\nu}(a),G_{\nu}(b))\setminus\{0\}$. We will write
    $K_{\nu}:(G_{\nu}(a),G_{\nu}(b))\setminus\{0\}\to\mathbb{R}\setminus[a,b]$ (1.12)
    its inverse.
  • For every $\nu\in\mathcal{P}(\mathbb{R})$, we let
    $\rho_{\nu,t}^{\Lambda}:=H_{\nu,t}(\Lambda)=\Lambda+tG_{\nu}(\Lambda)$
    and
    $\ell_{\nu,t}^{\Lambda}=\begin{cases}\rho_{\nu,t}^{\Lambda}&\text{if $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$}\\ \ell_{\nu,t}&\text{if $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$}.\end{cases}$
  • For every $\nu\in\mathcal{P}(\mathbb{R})$ and $\lambda\in[H_{\nu,t}(\ell_{\nu}),\ell_{\nu,t}]$, we denote $\omega_{\nu,t}^{*}(\lambda)$ the unique solution in $[\omega_{\nu,t}(\ell_{\nu,t}),\ell_{\nu})$ of the equation $H_{\nu,t}(\gamma)=\lambda$.
  • We denote $(e_{1},\ldots,e_{N})$ the canonical basis of $\mathbb{R}^{N}$.
  • We let $d$ be the bounded Lipschitz distance on probability measure on $\mathbb{R}$ defined by for all $\nu,\mu\in\mathcal{P}(\mathbb{R})$ by
    $d(\mu,\nu)=\sup\left|\int f\mathrm{d}(\mu-\nu)\right|,$ (1.13)
    where the supremum is taken over $1$-lipschitz functions $f:\mathbb{R}\to\mathbb{R}$ such that $\|f\|_{L^{\infty}}\leq 1$.
  • For each $p\geq 1$, we denote $\Delta^{p}$ the $p$-simplex
    $\Delta^{p}:=\{(y_{1},\ldots,y_{p})\in[0,1]^{p}:y_{1}+\cdots+y_{p}=1\}.$ (1.14)
1.4. Heuristics
We assume in these heuristics that the matrix $B_{N}$ has eigenvalues $\Lambda:=\eta_{0}<\eta_{1}<\cdots<\eta_{p}$ with multiplicities $1$, $N_{1},\ldots,N_{p}$, such that for each $i=1,\ldots,p$, $N_{i}/N$ converges to $\alpha_{i}>0$. We outline the arguments to derive the weak large deviation principle, namely to estimate the probability that the smallest eigenvalue $\lambda_{1}(X_{N})$ of $X_{N}=G_{N}+B_{N}$ is close to a given ${x\leq\ell_{\nu,t}}$. Let $\varepsilon>0$.
The probability that the smallest eigenvalue $\lambda_{1}(X_{N})$ of ${X_{N}}$ lies in $(x-\varepsilon,x+\varepsilon)$ is given by
$\mathbb{P}(\lambda_{1}(X_{N})\in(x-\varepsilon,x+\varepsilon))=\frac{1}{Z_{N}^{t}}\int_{\lambda_{1}(X_{N})\in(x-\varepsilon,x+\varepsilon)}e^{-\frac{N}{4t}\operatorname{Tr}((X_{N}-B_{N})^{2})}\,\mathrm{d}X_{N},$
where $\mathrm{d}X_{N}$ is the Lebesgue measure on symmetric matrices of size $N\times N$ and $Z_{N}^{t}$ is the normalizing constant. Let us more generally denote by $\lambda_{1}\leq\cdots\leq\lambda_{N}$ the eigenvalues of $X_{N}$ and ${(v_{1},\ldots,v_{N})}$ the associated unitary eigenvectors.
One can split the term in the density into
$\operatorname{Tr}((X_{N}-B_{N})^{2})=\lambda_{1}^{2}-2\lambda_{1}\langle v_{1},B_{N}v_{1}\rangle+2\langle v_{1},B_{N}^{2}v_{1}\rangle-\langle v_{1},B_{N}v_{1}\rangle^{2}+\operatorname{Tr}((p_{v_{1}}^{\perp}(X_{N}-B_{N})p_{v_{1}}^{\perp})^{2}),$
where $p_{v_{1}}^{\perp}$ stands for the projection onto the ortho-complement of $v_{1}$. Performing the change of variables $X_{N}\mapsto(\lambda_{1},\ldots,\lambda_{N},v_{1},O)$, we find a constant $C_{N}^{t}$ such that
$\int_{\lambda_{1}\in(x-\varepsilon,x+\varepsilon)}e^{-\frac{N}{4t}\operatorname{Tr}((X_{N}-B_{N})^{2})}\,\mathrm{d}X_{N}\\ =C_{N}^{t}\int_{\lambda_{1}\in(x-\varepsilon,x+\varepsilon)}\,\mathrm{d}\lambda_{1}\,\mathrm{d}v_{1}e^{-\frac{N}{4t}(\lambda_{1}^{2}-2\lambda_{1}\langle v_{1},B_{N}v_{1}\rangle+2\langle v_{1},B_{N}^{2}v_{1}\rangle-\langle v_{1},B_{N}v_{1}\rangle^{2})}\\ \times\prod_{j>1}|\lambda_{1}-\lambda_{j}|\int e^{-\frac{N}{4t}\operatorname{Tr}((OD(\lambda)O^{T}-p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp})^{2})}\prod_{2\leq i<j}|\lambda_{i}-\lambda_{j}|1_{\lambda_{1}\leq\ldots\leq\lambda_{N}}\,\mathrm{d}\lambda_{2}\cdots\,\mathrm{d}\lambda_{N}\,\mathrm{d}O,$
where $D(\lambda)=\mbox{diag}(\lambda_{2},\ldots,\lambda_{N})$ and $O$ follows the Haar measure on $v_{1}^{\perp}$. For every $\lambda\in\{\eta_{0},\ldots,\eta_{p}\}$, let us denote $E_{\lambda}$ the associated eigenspace of $B_{N}$. We observe that $\langle v_{1},B_{N}v_{1}\rangle$ and $\langle v_{1},B_{N}^{2}v_{1}\rangle$ only depend on ${Y_{N}:=\{\|v_{1}\|_{E_{\eta_{k}}}^{2},k=0,1,\ldots,p\}}$ so that the first term in the above density is a polynomial $L_{\nu,t}^{\Lambda}$ in $Y_{N}$ (defined in (2.16)).
Fix $Y=(y_{0},\ldots,y_{p})\in\Delta^{p+1}$, where $\Delta^{p+1}$ is the $(p+1)$-simplex (1.14). Restricting the domain of integration to vectors $v_{1}\in\mathbb{R}^{N}$ such that $Y_{N}$ is close to $Y$ and using the large deviation principle for $Y_{N}$ when $v_{1}$ follows the uniform law on the sphere, we roughly obtain, since the empirical measure of the eigenvalues of $B_{N}$ is close to $\nu:=\sum_{i=1}^{p}\alpha_{i}\delta_{\eta_{i}}$, the lower bound
$\liminf_{N\to\infty}\frac{1}{N}\log\frac{1}{Z_{N}^{t}}\int_{\lambda_{1}(X_{N})\in(x-\varepsilon,x+\varepsilon)}e^{-\frac{N}{4t}\operatorname{Tr}((X_{N}-B_{N})^{2})}\,\mathrm{d}X_{N}\gtrsim-J_{\nu,t}^{\Lambda}(x,Y)-\int\log|\lambda-x|\,\mathrm{d}(\nu\boxplus\sigma_{t})(\lambda)\\ +\frac{x^{2}}{4t}+C_{t}+\liminf_{N\to\infty}\frac{1}{N}\log\frac{1}{Z_{N-1}^{t(N-1)/N}}\int_{\lambda_{2}\geq x}e^{-\frac{N}{4t}\operatorname{Tr}((p_{v_{1}}^{\perp}(X_{N}-B_{N})p_{v_{1}}^{\perp})^{2})}\,\mathrm{d}X_{N},$ (1.15)
where $C_{t}$ is a finite constant and
$J_{\nu,t}^{\Lambda}:(\lambda,Y)\in(-\infty,\ell_{\nu,t}]\times\Delta^{p+1}\mapsto L_{\nu,t}^{\Lambda}(Y)-\frac{1}{2}\sum_{k=1}^{p}\alpha_{k}\log(y_{k}/\alpha_{k}).$
By Courant-Fischer’s theorem, when ${Y_{N}\simeq Y}$, the matrix $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$ has a unique outlier approximately equal to $\Lambda_{1}:=\Phi(\Lambda,Y)$, which is defined as the smallest solution of the equation
$\frac{y_{0}}{\Lambda-\Lambda_{1}}+\sum_{k=1}^{p}\frac{y_{k}}{\eta_{k}-\Lambda_{1}}=0.$
Moreover, the spectrum of $B_{N}^{\prime}=p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$ converges to ${\nu=\sum_{k=1}^{p}\alpha_{k}\delta_{\eta_{k}}}$. As a consequence, we see that the last term in (1.15) is also the probability of deviation of the smallest eigenvalue $\lambda_{2}$ of a $(N-1)\times(N-1)$ matrix given by the sum of a GOE matrix and a matrix $B_{N}^{\prime}$ with empirical measure of the eigenvalues converging to $\nu$ and with outlier $\Phi(\Lambda,Y)$. We set
$F_{\nu,t}^{-}(\Lambda,x):=-\lim_{\varepsilon\to 0}\liminf_{N\to\infty}\frac{1}{N}\log\mathbb{P}(\lambda_{1}(X_{N})\in(x-\varepsilon,x+\varepsilon)).$
As a consequence of the previous considerations, we obtain a functional inequality
$-F_{\nu,t}^{-}(\Lambda,x)\gtrsim\sup_{Y\in\Delta^{p+1}}\Bigl{\{}J_{\nu,t}^{\Lambda}(x,Y)-C_{t}-\int\log|\lambda-x|\,d(\nu\boxplus\sigma_{t})(\lambda)+\frac{x^{2}}{4t}-\inf_{y\geq x}F_{\nu,t}^{-}(\Phi(\Lambda,Y),y)\Bigr{\}},$
[…]
$-F_{\nu,t}^{+}(\Lambda,x)\lesssim\sup_{Y\in\Delta^{p+1}}\Bigl{\{}J_{\nu,t}^{\Lambda}(x,Y)-C_{t}-\int\log|\lambda-x|\,d(\nu\boxplus\sigma_{t})(\lambda)+\frac{x^{2}}{4t}-\inf_{y\geq x}F_{\nu,t}^{+}(\Phi(\Lambda,Y),y)\Bigr{\}}\,.$
We then solve the above functional inequalities. We first show that $F^{+}_{\nu,t}$ and $F^{-}_{\nu,t}$ vanish at ${\ell^{\Lambda}_{\nu,t}}$, are decreasing on $(-\infty,{\ell^{\Lambda}_{\nu,t}}]$ and increasing on $({\ell^{\Lambda}_{\nu,t}},\ell_{\nu,t}]$. Therefore $\inf_{y\geq x}F_{\nu,t}^{\pm}(\Phi(\Lambda,Y),y)$ vanish except when $x\geq{\ell^{\Phi(\Lambda,Y)}_{\nu,t}}$, namely when $\omega_{\nu,t}(x)\geq\Phi(\Lambda,Y)$. Then $\inf_{y\geq x}F_{\nu,t}^{\pm}(\Phi(\Lambda,Y),y)=F_{\nu,t}^{\pm}(\Phi(\Lambda,Y),x)$. We then prove that $F_{\nu,t}^{+},F_{\nu,t}^{-}$ are in fact equal to $I_{\nu,t}^{\Lambda}$ which satisfies the functional equation
$I_{\nu,t}^{\Lambda}(\lambda)=\inf_{Y\in\Delta^{p+1}}\Bigr{\{}J_{\nu,t}^{\Lambda}(\lambda,Y)-C_{t}-\int\log|\lambda-x|\mathrm{d}(\nu\boxplus\sigma_{t})(x)+\frac{\lambda^{2}}{4t}+I_{\nu,t}^{\Phi(\Lambda,Y)}(\lambda)1_{\omega_{\nu,t}(\lambda)\geq\Phi(\Lambda,Y)}\Bigr{\}}\,.$
This establishes the weak LDP which can easily be upgraded to a full LDP. We finally extend the result to general matrices that satisfy Assumptions 1.1. This is achieved by bounding $B_{N}$ from below and above by two matrices that have a finite number of eigenvalues, and by using monotonicity arguments.
Acknowledgments
We would like to thank Gérard Ben Arous for proposing the problem and for numerous useful and interesting discussions.
This project was partly supported by the ERC Project LDRAM : ERC-2019-ADG Project 884584
2. Large deviations in the pure point case
Let $B_{N}$ be a $N\times N$ diagonal matrix with eigenvalues $\Lambda$ with multiplicity $1$ and $\eta_{1},\ldots,\eta_{p}$ with multiplicity $N_{1},\cdots,N_{p}$ such that that for each $i\in\{1,\ldots,p\}$, $N_{i}/N$ converges towards $\alpha_{i}>0$. Let $G_{N}$ be a GOE matrix of size $N\times N$ and variance $t>0$. We assume that $\Lambda<\eta_{1}<\eta_{2}<\cdots<\eta_{p}$ and denote
$\nu:=\sum_{i=1}^{p}\alpha_{i}\delta_{\eta_{i}}\,.$ (2.1)
Let $\Delta^{p+1}$ be the $(p+1)$-simplex as defined in (1.14). We recall that $\nu\boxplus\sigma_{t}$ is the limiting distribution of the spectrum of $G_{N}+B_{N}$ and $\ell_{\nu,t}$ denotes the leftmost point in the support of $\nu\boxplus\sigma_{t}$. The goal of this section is to prove the following large deviation principle, which specializes Theorem 1.2 to the case where $B_{N}$ has a fixed number of distinct eigenvalues. Let $I_{\nu,t}^{\Lambda}$ be as defined in (1.9).
Proposition 2.1 (LDP in the pure point case).
Let $\nu$ be as in (2.1) and $\Lambda\leq\ell_{\nu}=\inf_{1\leq k\leq p}\eta_{k}$.
  1. (1)
    The map $I_{\nu,t}^{\Lambda}$ is a good rate function which is strictly convex and vanishes at ${\ell^{\Lambda}_{\nu,t}}$.
  2. (2)
    Let $\lambda_{1}$ be the smallest eigenvalue of $X_{N}=G_{N}+B_{N}$. Then the law of $\lambda_{1}$ satisfies a large deviation principle with speed $N$ and rate function $I_{\nu,t}^{\Lambda}$.
It is easy to see that $I_{\nu,t}^{\Lambda}$ is a good rate function. Indeed, it is continuous on $(-\infty,\ell_{\nu,t})$ since $S_{\nu}$ is continuous on $(-\infty,\ell_{\nu})$, $\omega_{\nu,t}$ is continuous and increasing on $(-\infty,\ell_{\nu,t})$, with values in $(-\infty,\omega_{\nu,t}(\ell_{\nu,t})]$, and $\omega_{\nu,t}^{*}$ is continuous with values in $[\omega_{\nu,t}(\ell_{\nu}),\ell_{\nu,t}]$. This shows that $I_{\nu,t}^{\Lambda}$ is continuous expect possibly at $\ell_{\nu,t}$ where it is lower semicontinuous. Moreover, $I_{\nu,t}^{\Lambda}(\ell^{\Lambda}_{\nu,t})=0$. Indeed, if $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$, then $\ell_{\nu,t}^{\Lambda}=\rho_{\nu,t}^{\Lambda}$ and $\omega_{\nu,t}(\rho^{\Lambda}_{\nu,t})=\Lambda=\gamma(\lambda,\Lambda)$. Besides, if $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$, $\ell_{\nu,t}^{\Lambda}=\ell_{\nu,t}$ and $\omega_{\nu,t}(\ell_{\nu,t})=\omega_{\nu,t}^{*}(\ell_{\nu,t})=\gamma(\lambda,\Lambda)$. Let us show that
$\lim_{\lambda\to-\infty}{I_{\nu,t}^{\Lambda}}(\lambda)=+\infty.$
As $\lambda\to-\infty$, we can notice that $\omega_{\nu,t}(\lambda)\sim\lambda$. It follows that
${I_{\nu,t}^{\Lambda}}(\lambda)=\frac{1}{2}S_{\nu}(\omega_{\nu,t}(\lambda))+\frac{\lambda^{2}}{2t}+o(\lambda^{2})=\frac{\lambda^{2}}{2t}+o(\lambda^{2})\underset{\lambda\to-\infty}{\longrightarrow}+\infty.$
Hence, the level sets of $I_{\nu,t}^{\Lambda}$ are compact. The convexity of $I_{\nu,t}^{\Lambda}$ is established in Lemma 3.1, along with the fact that the minimum is achieved at $\ell^{\Lambda}_{\nu,t}$. It follows in particular that $I_{\nu,t}^{\Lambda}$ is non-negative. Hence, $I_{\nu,t}^{\Lambda}$ is a convex good rate function.
Moreover, the fact that $\lambda_{1}$ is exponentially tight is clear since $B_{N}$ is uniformly bounded and the smallest eigenvalue of the GOE is well known to be exponentially tight [AGZ10, Section 2.6.2]. We more precisely find, see the proof of [GH20, Lemma 1.8], that there exists a finite constant $C$ such that for $N$ large enough,
$\mathbb{P}(\|X_{N}\|_{\infty}\geq K)\leq e^{-N(K-C)}.$ (2.2)
The main point of the proof, on which we will focus for the remainder of this section, is therefore to prove a weak large deviation principle, namely that for every $\lambda<\ell_{\nu,t}$,
$-I_{\nu,t}^{\Lambda}(\lambda)\leq\lim_{\delta\rightarrow 0}\liminf_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-\lambda|\leq\delta)\leq\lim_{\delta\rightarrow 0}\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-\lambda|\leq\delta)\leq-I_{\nu,t}^{\Lambda}(\lambda).$ (2.3)
Indeed the large deviation principle (2.3) is easy to prove for $x>\ell_{\nu,t}$. Indeed, by concentration of measure, we know that for every $\zeta>0$, there exists $C_{\zeta}>0$ such that for $N$ large enough,
$\mathbb{P}\Bigr{(}d\Bigr{(}\frac{1}{N}\sum\delta_{\lambda_{i}},\nu_{\alpha}\boxplus\sigma_{t}\Bigr{)}\geq\zeta\Bigr{)}\leq e^{-C_{\zeta/2}N^{2}}\,,$ (2.4)
where $d$ is the distance defined in (1.13) and $\nu=\nu_{\alpha}=\sum\alpha_{i}\delta_{\eta_{i}}$. For further use, we observe that (2.4) is uniform in $\alpha\in\Delta^{p}$. However, if $\lambda>\ell_{\nu,t}$, for $\varepsilon>0$ such that $\lambda>\ell_{\nu,t}+2\varepsilon$ there exists $\zeta>0$ small enough so that $\{|\lambda_{1}-\lambda|\leq\varepsilon\}\subset\{d\Bigr{(}\frac{1}{N}\sum\delta_{\lambda_{i}},\nu_{\alpha}\boxplus\sigma_{t}\Bigr{)}\geq\zeta\}$ from which we conclude that
$\lim_{\delta\rightarrow 0}\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-\lambda|\leq\delta)\leq-\infty\,.$
We do not need to prove (2.3) at $\lambda=\ell_{\nu,t}$ either. Indeed, if $\ell_{\nu,t}=\ell^{\Lambda}_{\nu,t}$, then Theorem 1.1 shows that for every $\delta>0$,
$\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-\ell_{\nu,t}^{\Lambda}|\leq\delta)=\liminf_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-\ell_{\nu,t}^{\Lambda}|\leq\delta)=0,$
which corresponds to the announced statement, since by Lemma B.1, $I_{\nu,t}^{\Lambda}(\ell_{\nu,t}^{\Lambda})=0$.
If ${\ell_{\nu,t}}>\ell^{\Lambda}_{\nu,t}$, the monotonicity of the large deviation rate function proven in Lemma B.1 implies that for every $\delta>0$ small enough, every $\gamma<\delta/4$,
$\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-\ell_{\nu,t}|\leq\delta)\leq\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-\ell_{\nu,t}-\delta|\leq\gamma)$
so that (2.3) at $\lambda=\ell_{\nu,t}-\delta$ gives, after taking the limit $\gamma\rightarrow 0$,
$\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-\ell_{\nu,t}|\leq\delta)\leq-I_{\nu,t}^{\Lambda}(\ell_{\nu,t}-\delta)$
and the left-continuity of $I_{\nu,t}^{\Lambda}$ at $\ell_{\nu,t}$ implies
$\lim_{\delta\rightarrow 0}\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-\ell_{\nu,t}|\leq\delta)\leq-I_{\nu,t}^{\Lambda}(\ell_{\nu,t})\,.$ (2.5)
Similarly for the lower bound, for every $\delta>0$ and $\gamma\in(0,\delta/4)$, we have
$\liminf_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-\ell_{\nu,t}|\leq\delta)\geq\liminf_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-\ell_{\nu,t}-\delta/2|\leq\gamma),$
which implies by (2.3) that
$\liminf_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-\ell_{\nu,t}|\leq\delta)\geq-I_{\nu,t}^{\Lambda}(\ell_{\nu,t}-\delta/2).$
Equation (2.3) at ${\lambda=\ell_{\nu,t}<\ell^{\Lambda}_{\nu,t}}$, then follows from the left-continuity of $I_{\nu,t}^{\Lambda}$ at $\ell_{\nu,t}$ (and (2.5)).
Hence, the rest of this section is dedicated to the proof of (2.3). As announced in the heuristics, the proof of (2.3) starts by factorizing the terms depending on $\lambda_{1}$.
2.1. Projection onto the first eigenvector
Let $v_{1}$ be an eigenvector of $X_{N}$ of norm $1$ associated to $\lambda_{1}$. One can first project onto $v_{1}^{\perp}$ using the following lemma:
Lemma 2.2.
Let $M,X\in\mathcal{S}_{N}(\mathbb{R})$. Let $\lambda_{1}$ be an eigenvalue of $X$ and $v_{1}\in S^{N-1}$ be an associated eigenvector. One may write
$\mathrm{Tr}((X-M)^{2})=\lambda_{1}^{2}-2\lambda_{1}\langle v_{1},Mv_{1}\rangle+2\langle v_{1},M^{2}v_{1}\rangle-\langle v_{1},Mv_{1}\rangle^{2}+\mathrm{Tr}((p_{v_{1}}^{\perp}(X-M)p_{v_{1}}^{\perp})^{2}),$ (2.6)
where $p_{v_{1}}^{\perp}$ is the project onto the orthocomplement of $v_{1}$.
Proof.
Let $v_{1},\ldots,v_{N}$ be a family of eigenvectors of $X$. By writing the matrix $X-M$ in the basis $v_{1},\ldots,v_{N}$, one can check that
$\mathrm{Tr}((X-M)^{2})=\langle v_{1},(X-M)v_{1}\rangle^{2}+2\sum_{j>1}\langle v_{1},(X-M)v_{j}\rangle^{2}+\mathrm{Tr}(p_{v_{1}}^{\perp}(X-M)p_{v_{1}}^{\perp})^{2}.$ (2.7)
One may then write
$(X-M)v_{1}=\sum_{j=1}^{N}\langle v_{j},(X-M)v_{1}\rangle v_{j}.$
Therefore
$\langle v_{1},(X-M)^{2}v_{1}\rangle=\sum_{j=1}^{N}\langle v_{j},(X-M)v_{1}\rangle^{2}$
which allows one to rewrite (2.7) as
$\mathrm{Tr}((X-M)^{2})=2\langle v_{1},(X-M)^{2}v_{1}\rangle-\langle v_{1},(X-M)v_{1}\rangle^{2}+\mathrm{Tr}(p_{v_{1}}^{\perp}(X-M)p_{v_{1}}^{\perp})^{2}.$
One can finally check using $Xv_{1}=\lambda_{1}v_{1}$ that
$2\langle v_{1},(X-M)^{2}v_{1}\rangle-\langle v_{1},(X-M)v_{1}\rangle^{2}=\lambda_{1}^{2}-2\lambda_{1}\langle v_{1},Mv_{1}\rangle+2\langle M^{2}v_{1},v_{1}\rangle-\langle v_{1},Mv_{1}\rangle^{2}.$
Let $w_{0}$ be a unitary eigenvector of $B_{N}$ associated to $\Lambda$. For each $k=1,\ldots,p$, let us denote by $P_{E_{\eta_{k}}}$ the orthogonal projection onto the eigenspace $E_{\eta_{k}}$ corresponding to the eigenvalue $\eta_{k}$. The quantities of Lemma 2.2 can then be written as follows:
$\langle v_{1},B_{N}v_{1}\rangle$ $=\Lambda\langle v_{1},w_{0}\rangle^{2}+\sum_{k=1}^{p}\eta_{k}\|P_{E_{\eta_{k}}}v_{1}\|^{2}$ (2.8)
$\langle v_{1},B_{N}^{2}v_{1}\rangle$ $=\Lambda^{2}\langle v_{1},w_{0}\rangle^{2}+\sum_{k=1}^{p}\eta_{k}^{2}\|P_{E_{\eta_{k}}}v_{1}\|^{2}.$ (2.9)
Define $Y^{N}=(Y_{0}^{N},\ldots,Y_{k}^{N})$ by
$Y_{0}^{N}=\langle v_{1},w_{0}\rangle^{2},\quad Y_{k}^{N}=\|P_{E_{\eta_{k}}}v_{1}\|^{2},\quad 1\leq k\leq p.$ (2.10)
The variable $Y^{N}$ takes values into the simplex $\Delta^{p+1}$.
2.2. LDP for Dirichlet variables
One can observe that the random variable $Y^{N}$ follows a Dirichlet law with parameter $(1,N_{1},\ldots,N_{p})$ which are such that $N_{i}/N$ goes to $\alpha_{i}>0$ for each $i=1,\ldots,p$. Using the explicit density of the Dirichlet distribution, it easy to prove [GH22] by Laplace’s method that $Y^{N}$ satisfies the following LDP:
Lemma 2.3.
Let $Y^{N}$ be as in (2.10). The sequence of random variables $(Y^{N})$ satisfies a large deviation principle at speed $N$ and with rate function given by
$I_{\nu}:Y=(y_{0},\ldots,y_{p})\in\Delta^{p+1}\mapsto-\frac{1}{2}\sum_{k=1}^{p}\alpha_{k}\log(y_{k}/\alpha_{k}).$ (2.11)
2.3. Outliers of $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$
As outlined in Subsection 1.4, the proof involves computing the probability that the smallest eigenvalue of $G_{N-1}+p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$ is larger than $x$, where $G_{N-1}$ is a GOE matrix of size $(N-1)\times(N-1)$ on the orthocomplement of $v_{1}$, for given $x\leq\ell_{\nu,t}$ and $Y\in\Delta^{p+1}$. Thus, we first study the outliers of the matrix $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$ when $v_{1}$ is such that $Y_{N}$ is equal to some $Y$, and then invoke known results on the BBP transition.
Lemma 2.4.
Let $B_{N}$ be a $N\times N$ diagonal matrix with eigenvalues $\eta_{0}:=\Lambda<\eta_{1}<\cdots<\eta_{p}$ with multiplicity $N_{0}=1,N_{1},\cdots,N_{p}$ such that that for each $i\in\{1,\ldots,p\}$, $N_{i}\geq 2$. For $Y\in\Delta^{p+1}$ so that $y_{0}\in(0,1)$, we set $\Phi(\Lambda,Y)\in[\Lambda,\eta_{1}]$ to be the smallest solution to the equation
$\frac{y_{0}}{\Lambda-\Lambda_{1}}+\sum_{k=1}^{p}\frac{y_{k}}{\eta_{k}-\Lambda_{1}}=0.$ (2.12)
When $y_{0}=0$, we set $\Phi(\Lambda,Y)=\Lambda$ and when $y_{0}=1$ we set $\Phi(\Lambda,Y)=\eta_{1}$. Then $\Phi(\Lambda,.)$ is continuous. Moreover:
  1. (1)
    Let $v_{1}\in\mathbb{R}^{N}$ be a unit vector, $v_{1}\notin\cup_{k=0}^{p}E_{\eta_{k}}$. For each $k=0,\ldots,p$, denote
    $y_{k}:=\|v_{1}\|_{E_{\eta_{k}}}^{2}.$ (2.13)
    Then the matrix $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$ has eigenvalues $\Lambda_{1}<\cdots<\Lambda_{p}$ of multiplicities $1$ and $\eta_{1},\ldots\eta_{p}$ of multiplicities $N_{1}-1,\ldots,N_{p}-1$. Moreover for each $k=1,\ldots,p$, $\Lambda_{k}\in(\eta_{k-1},\eta_{k})$. Finally, $\Lambda_{1}=\Phi(\Lambda,Y)$.
  2. (2)
    Let $v_{1}$ be a unit vector such that $v_{1}\in E_{\eta_{k}}$ for some $k=0,\ldots,p$. Then the matrix $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$ has eigenvalues $\eta_{i}$ with multiplicity $N_{i}$ for each $i\in\{0,\ldots,p\}\setminus\{k\}$, and $\eta_{k}$ with multiplicity $N_{k}-1$. Moreover, the smallest eigenvalue $\Lambda_{1}$ of $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$ equals $\Phi(\Lambda,Y)$.
Proof.
Let us prove (1). Let $v_{1}$ be a unit vector, which is not an eigenvector of $B_{N}$. Let $B_{N}^{\prime}:=p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$. Let us show that $\eta_{k}$ has multiplicity $N_{k}-1$. The eigenspace of $B_{N}^{\prime}$ associated to $\eta_{k}$ is
$E_{\eta_{k}}^{\prime}=E_{\eta_{k}}\cap v_{1}^{\perp}.$
Since $v_{1}\notin E_{\eta_{k}}$, $E_{\eta_{k}}^{\prime}$ has dimension $N_{k}-1$. Moreover $\Lambda$ is not an eigenvalue of $B_{N}^{\prime}$. There are therefore $p$ new eigenvalues $\Lambda_{1}\leq\cdots\leq\Lambda_{p}$. Recall that if $\lambda_{1}\leq\cdots\leq\lambda_{N}$ are the eigenvalues of $B_{N}$ and $\lambda_{1}^{\prime}\leq\cdots\leq\lambda_{N}^{\prime}$ are the eigenvalues of $B_{N}^{\prime}$, then by the Courant-Fischer theorem, $\lambda$ and $\lambda^{\prime}$ are interlaced:
$\lambda_{1}\leq\lambda_{1}^{\prime}\leq\cdots\leq\lambda^{\prime}_{N-1}\leq\lambda_{N}.$
Therefore for each $i=1,\ldots,p$, we have $\Lambda_{i}\in(\eta_{i-1},\eta_{i}).$
Moreover, by the Courant-Fischer theorem, $\Lambda_{1}$ can be expressed as
$\Lambda_{1}=\inf_{v\in\mathbb{R}^{N}:\|v\|_{2}=1,\langle v,v_{1}\rangle=0}\langle v,B_{N}v\rangle.$
Let $v\in\mathbb{R}^{N}$ be a minimizer in the above problem. By variational calculus, there exist $C_{1},C\in\mathbb{R}$ such that
$B_{N}v=Cv+C_{1}v_{1}.$
Since $\langle v,v_{1}\rangle=0$, we have
$\langle v,B_{N}v\rangle=C=\Lambda_{1}.$
Since $\Lambda_{1}\notin\{\lambda_{1},\ldots,\lambda_{N}\}$, we deduce that for each $i=1,\ldots,N$,
$v_{i}=C_{1}\frac{(v_{1})_{i}}{\lambda_{i}-\Lambda_{1}}.$
Thus $\Lambda_{1}$ satisfies
$\langle v,v_{1}\rangle=\sum_{i=1}^{N}\frac{(v_{1})_{i}^{2}}{\lambda_{i}-\Lambda_{1}}=0,$
which can be written
$\frac{y_{0}}{\Lambda_{1}-\Lambda}+\sum_{k=1}^{p}\frac{y_{k}}{\Lambda_{1}-\eta_{k}}=0.$
Item (2) is clear. Moreover, the continuity of $\Phi(\Lambda,Y)$ is clear from (2.12). Indeed, if $\varepsilon>0$ is such that $Y\subset[\varepsilon,1-\varepsilon]\times[0,1]^{p}\cap\Delta^{p+1}$, we see from (2.12) that
$\varepsilon(\eta_{1}-\Lambda_{1})\leq\Lambda_{1}-\Lambda\leq\varepsilon^{-1}(\eta_{1}-\Lambda_{1})$
so that $\Lambda_{1}$ is at distance of order $\varepsilon$ of $\Lambda$ and $\eta_{1}$. Then, if $Y,Y^{\prime}\in[\varepsilon,1-\varepsilon]\times[0,1]^{p}\cap\Delta^{p+1}$, we easily deduce that
$|\Phi(\Lambda,Y)-\Phi(\Lambda,Y^{\prime})|\leq C(\varepsilon)\sum_{i=1}^{p}|y_{i}-y_{i}^{\prime}|\,.$
When $y_{0}$ goes to zero or one, it is clear that $\Phi(\Lambda,Y)$ was extended by continuity, which in turn completes the proof of the continuity of $\Phi(\Lambda,.)$. ∎
We next study the typical behavior of the smallest eigenvalue of $p_{v_{1}}^{\perp}(G_{N}+B_{N})p_{v_{1}}^{\perp}$ in $v_{1}^{\perp}$, when $v_{1}$ is as in Lemma 2.4. It is a straightforward application of Lemma 2.4 and Theorem 1.1 as this matrix is the sum of a GOE matrix and the matrix $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$ with spectrum asymptotically distributed according to $\nu$ and outlier $\Phi(Y,\Lambda)$ as in Lemma 2.4.
Lemma 2.5 (Smallest eigenvalue of the remainder).
Let $B_{N}$ be a $N\times N$ diagonal matrix with eigenvalue $\Lambda$ with multiplicity $1$ and $\eta_{1},\ldots,\eta_{p}$ of multiplicities $N_{i}$ such that $N_{i}/N$ goes to $\alpha_{i}>0$. Let $\nu:=\sum_{k=1}^{p}\alpha_{k}\delta_{\eta_{k}}$. Let $\lambda\leq\ell_{\nu,t}$. Let $v_{1}$ and $Y$ be as in Lemma 2.4. Then the smallest eigenvalue of $p_{v_{1}}^{\perp}(G_{N}+B_{N})p_{v_{1}}^{\perp}$ in $v_{1}^{\perp}$ converges towards $H_{\nu,t}(\Phi(Y,\Lambda))$ if $\Phi(Y,\Lambda)\leq\omega_{\nu,t}(\ell_{\nu,t})$ and to $\ell_{\nu,t}$ otherwise.
2.4. A functional inequality for the large deviation upper bound
We denote for $\alpha\in\Delta^{p}$, $\mathcal{B}_{\delta}^{N}(\alpha)$ the $\ell_{\infty}$-ball
$\mathcal{B}_{\delta}^{N}(\alpha)=\Bigr{\{}\mathsf{M}=(M_{1},\ldots,M_{p})\in(\mathbb{N}^{*})^{p}:\sup_{1\leq i\leq p}\Bigr{|}\frac{M_{i}}{N}-\alpha_{i}\Bigr{|}\leq\delta,\quad\sum_{i=1}^{p}M_{i}=N-1\Bigr{\}}.$
Let $B_{N}$ be a diagonal matrix with eigenvalues $\Lambda$ with multiplicity $1$, $\eta_{1},\ldots,\eta_{p}$ of multiplicities $M_{1},\ldots,M_{p}$. Let $\mathsf{M}=(M_{1},\ldots,M_{p})\in\mathcal{B}_{\delta}^{N}(\alpha)$. For each $\Lambda\leq\ell_{\nu}$, let us denote $\mathbb{P}_{N,\Lambda,\mathsf{M}}$ the law of $X_{N}=G_{N}+B_{N}$ and
$P_{N,\Lambda,\mathsf{M}}(x,\varepsilon):=\mathbb{P}_{N,\Lambda,\mathsf{M}}(|\lambda_{1}(X_{N})-x|\leq\varepsilon),$
Recall that for $\alpha\in\Delta^{p}$, we denote $\nu=\sum\alpha_{i}\delta_{\eta_{i}}$. For all $\Lambda\leq\ell_{\nu}$ and $\lambda\leq\ell_{\nu,t}$, let us define
$F_{\nu,t}^{+}(\Lambda,\lambda):=-\limsup_{\Lambda^{\prime}\to\Lambda}\limsup_{\varepsilon\to 0}\lim_{\delta\downarrow 0}\limsup_{N\to\infty}\sup_{\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)}\frac{1}{N}\log P_{N,\Lambda,\mathsf{M}}(\lambda,\varepsilon).$ (2.14)
Remark 2.1.
Let us remark that, by construction, $\Lambda\mapsto F_{\nu,t}^{+}(\Lambda,\lambda)$ is lower semicontinuous. Indeed, if $f:\mathbb{R}\to\mathbb{R}$ is a measurable function, the function
$g:x\in\mathbb{R}\mapsto\lim_{\varepsilon\to 0}\sup_{B(x,\varepsilon)}f$
is upper semicontinuous.
Recall the map $\Phi$ defined in Lemma 2.4. Let us define
$C_{t}:=\frac{1}{2}-\frac{1}{2}\ln t.$ (2.15)
We consider the map $L_{\nu,t}^{\Lambda}:(-\infty,\ell_{\nu,t}]\times\Delta^{p+1}\to\mathbb{R}$, given for all $\lambda\in(-\infty,\ell_{\nu,t}]$ and $Y\in\Delta^{p+1}$ by
$L_{\nu,t}^{\Lambda}(\lambda,Y):=-\frac{\lambda}{2t}\Bigl{(}\Lambda y_{0}+\sum_{k=1}^{p}\eta_{k}y_{k}\Bigr{)}+\frac{1}{2t}\Bigl{(}\sum_{k=1}^{p}\eta_{k}^{2}y_{k}+\Lambda^{2}y_{0}\Bigr{)}-\frac{1}{4t}\Bigl{(}\Lambda y_{0}+\sum_{k=1}^{p}\eta_{k}y_{k}\Bigr{)}^{2}.$ (2.16)
Recall the map $I_{\nu}:Y\in\Delta^{p+1}\mapsto-\frac{1}{2}\sum_{k=1}^{p}\alpha_{k}\log(y_{k}/\alpha_{k})$ from Lemma 2.3. We set
$J_{\nu,t}^{\Lambda}:(\lambda,Y)\in(-\infty,\ell_{\nu,t}]\times\Delta^{p+1}\mapsto L_{\nu,t}^{\Lambda}(\lambda,Y)+I_{\nu}(Y).$ (2.17)
We also let
$K_{\nu,t}:(\Lambda,\lambda,Y)\in(-\infty,\ell_{\nu}]\times(-\infty,\ell_{\nu,t}]\times\Delta^{p+1}\mapsto J_{\nu,t}^{\Lambda}(\lambda,Y)-C_{t}-\int\log|\lambda-x|\mathrm{d}(\nu\boxplus\sigma_{t})(x)+\frac{\lambda^{2}}{4t}\,.$ (2.18)
We denote $f(\lambda^{+})=\lim_{\varepsilon\downarrow 0}f(\lambda+\varepsilon)$. The goal of this section is to prove the following:
Lemma 2.6 (Functional upper bound).
Let $\Lambda\leq\ell_{\nu}$. For every $\lambda<\ell_{\nu,t}$, we have
$F_{\nu,t}^{+}(\Lambda,\lambda)\geq\inf_{Y\in\Delta^{p+1}}\Bigr{\{}K_{\nu,t}(\Lambda,\lambda,Y)+F_{\nu,t}^{+}(\Phi(\Lambda,Y),\lambda^{-})1_{\omega_{\nu,t}(\lambda)>\Phi(\Lambda,Y)}\Bigr{\}}$ (2.19)
Moreover, $\lambda\mapsto F_{\nu,t}^{+}(\Lambda,\lambda)$ decreases on $(-\infty,\ell^{\Lambda}_{\nu,t})$ and increases on $(\ell^{\Lambda}_{\nu,t},\ell_{\nu,t})$, while it vanishes at $\ell^{\Lambda}_{\nu,t}$. It is bounded from below by $c(\lambda-\ell^{\Lambda}_{\nu,t})^{2}$ for some $c>0$.
Proof.
The properties of $F^{+}_{\nu,t}$ listed at the end of the lemma are proven in Lemma B.2 and Lemma B.1, and we concentrate below on the proof of the functional inequality (2.19). Let $\varepsilon>0$. Let us perform the change of variables $X_{N}\mapsto(\lambda_{1},\ldots,\lambda_{N},v_{1},\ldots,v_{N})$ where $\lambda_{1}<\ldots<\lambda_{N}$ are the eigenvalues of $X_{N}$ and $v_{1},\ldots,v_{N}$ the associated eigenvectors. One has
$P_{N,\Lambda,\mathsf{M}}(x,\varepsilon)=\frac{1}{Z_{N}^{t}}\int_{|\lambda_{1}-x|\leq\varepsilon}\prod_{i<j}|\lambda_{i}-\lambda_{j}|1_{\lambda_{1}<\ldots<\lambda_{N}}e^{-\frac{N}{4t}\mathrm{Tr}((OD(\lambda)O^{T}-B_{N})^{2})}\mathrm{d}O\mathrm{d}\lambda_{1}\ldots\mathrm{d}\lambda_{N},$
where $\mathrm{d}O$ is the Haar measure on $O_{N}(\mathbb{R})$, $D(\lambda)$ the diagonal matrix with entries $(\lambda_{i})_{1\leq i\leq N}$ and the constant $Z_{N}^{t}$ the value of the integral when the constraint on $\lambda_{1}$ is removed. Applying Lemma 2.2 to $M=B_{N}$ and using (2.9) gives, by the definition (2.16) of $L_{\nu,t}^{\Lambda}$,
$P_{N,\Lambda,\mathsf{M}}(x,\varepsilon)=\frac{1}{Z^{t}_{N}}\int_{S^{N-1}}\mathrm{d}v_{1}\int\mathrm{d}\mathbb{P}_{v_{1}}(v_{2},\ldots,v_{N})\\ \int_{|\lambda_{1}-x|\leq\varepsilon\atop\lambda_{1}<\ldots<\lambda_{N}}\prod_{i<j}|\lambda_{i}-\lambda_{j}|e^{-NL_{\nu,t}^{\Lambda}(\lambda_{1},\langle v_{1},w_{0}\rangle^{2},\ldots,\|P_{E_{\eta_{p}}}v_{1}\|^{2})-\frac{N}{4t}\lambda_{1}^{2}-\frac{N}{4t}\mathrm{Tr}(p_{v_{1}}^{\perp}(OD(0,\lambda)O^{T}-B_{N})p_{v_{1}}^{\perp})^{2}}\mathrm{d}\lambda_{1}\ldots\mathrm{d}\lambda_{N},$
where $\mathbb{P}_{v_{1}}$ is the Haar measure on the orthocomplement of $v_{1}$ and $D(0,\lambda)$ is the diagonal matrix with entries $(0,\lambda_{2},\ldots,\lambda_{N})$. If $\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)$, $d(\nu_{\alpha},\nu_{\mathsf{M}/N})\leq\delta$ and therefore $d(\nu_{\alpha}\boxplus\sigma_{t},\nu_{\mathsf{M}/N}\boxplus\sigma_{t})\leq\delta$. Hence, taking $\zeta=2\delta$ in (2.4), we have, uniformly on $\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)$,
$\mathbb{P}_{N,\Lambda,\mathsf{M}}\Bigr{(}d\Bigr{(}\frac{1}{N}\sum\delta_{\lambda_{i}},\nu_{\alpha}\boxplus\sigma_{t}\Bigr{)}\geq\delta\Bigr{)}\leq e^{-\frac{1}{2}C_{\delta}N^{2}}\,.$ (2.20)
Hereafter $\alpha$ is fixed and we denote in short $\nu$ for $\nu_{\alpha}$. Moreover, since the sequence $(B_{N})_{N\in\mathbb{N}}$ is bounded uniformly for the operator norm, there exists $L_{0}$ finite ($L_{0}=(2+\|B_{N}\|_{\infty})^{2}$) such that for every $L>L_{0}$, by [AGZ10, (2.6.21)]
$\sup_{\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)}\mathbb{P}_{N,\Lambda,\mathsf{M}}\Bigr{(}\frac{1}{N}\sum\lambda_{i}^{2}\geq L\Bigr{)}\leq e^{-(L-L_{0})N^{2}}\,.$ (2.21)
For every $\kappa>0$, set
$f_{\kappa}(x):=\min\{\log(\max(|x|,\kappa)),\kappa^{-1}\}+\frac{\kappa^{2}}{\log\kappa^{-1}}(L_{0}+1).$
It is easy to check that on $\{d(\frac{1}{N}\sum\delta_{\lambda_{i}},\nu\boxplus\sigma_{t})\leq\delta\}\cap\{\frac{1}{N}\sum\lambda_{i}^{2}\leq L_{0}+1\}$,
$\prod_{j:j>1}|\lambda_{1}-\lambda_{j}|\leq\exp\Bigr{\{}\sum_{i=1}^{N}f_{\kappa}(\lambda_{1}-\lambda_{i})\Bigr{\}}\leq\exp\Bigr{\{}N\int f_{\kappa}(\lambda_{1}-x)\mathrm{d}(\sigma_{t}\boxplus\nu)(x)+NC_{\kappa}\delta\Bigr{\}},$
where we finally used that $f_{\kappa}$ is bounded continuous to find a finite constant $C_{\kappa}$ such that the last inequality holds. Finally, if moreover $\lambda_{1}$ is close to $x<\ell_{\nu,t}$, and therefore at positive distance from the support of $\sigma_{t}\boxplus\nu$, because $\sigma_{t}\boxplus\nu$ is compactly supported,
$\int f_{\kappa}(\lambda_{1}-x)\mathrm{d}(\sigma_{t}\boxplus\nu)(x)=\int\log|\lambda_{1}-x|\mathrm{d}(\sigma_{t}\boxplus\nu)(x)+O(\varepsilon)+O(\kappa)\,.$
Let $\varepsilon^{\prime}>0$. Since $I_{\nu}$ is lower semicontinuous, one can cover $\Delta^{p+1}$ by a finite union of measurable sets $\cup_{1\leq k\leq M}E_{Y_{k},\kappa_{k},\varepsilon^{\prime}}$ where for $Y\in\Delta^{p+1}$, and $\kappa,\varepsilon^{\prime}>0$,
[…]
and $(Y_{k})_{1\leq k\leq M}$ is a family of points of $\Delta^{p+1}$. We denote in short $E_{Y_{k}}$ for $E_{Y_{k},\kappa_{k},\varepsilon^{\prime}}$. We deduce by continuity of $L_{\nu,\Lambda}$ together with (2.20), (2.21) that
$P_{N,\Lambda,\mathsf{M}}(x,\varepsilon)\leq\frac{2\varepsilon}{Z_{N}^{t}}\sum_{k=1}^{M}e^{N(\int\log|x-y|\mathrm{d}(\sigma_{t}\boxplus\nu)(x)-L_{\nu,t}^{\Lambda}(x,Y_{k})+o_{\kappa}(\delta,\epsilon,\epsilon^{\prime})-\frac{1}{4t}x^{2})}\int_{E_{Y_{k}}}\mathrm{d}v_{1}\int\mathrm{d}\mathbb{P}_{v_{1}}(v_{2},\ldots,v_{N})\qquad\qquad$
$\qquad\int_{\mathcal{B}(\delta,\varepsilon,x)}\prod_{2\leq i<j}|\lambda_{i}-\lambda_{j}|e^{-\frac{N}{4t}\mathrm{Tr}((p_{v_{1}}^{\perp}(OD(0,\lambda)O^{T}-B)p_{v_{1}}^{\perp})^{2})}\mathrm{d}\lambda_{2}\ldots\mathrm{d}\lambda_{N}+O(e^{-\min\{C_{\delta},1\}N^{2}})\,,$ (2.23)
where $o_{\kappa}(\delta,\epsilon,\epsilon^{\prime})=O(\varepsilon)+O(\varepsilon^{\prime})+C_{\kappa}(\delta)+O(\kappa)$ and
$\mathcal{B}(\delta,\varepsilon,x):=\{\lambda_{1}<\cdots<\lambda_{N}\}\cap\{|\lambda_{1}-x|\leq\varepsilon\}\cap\Bigr{\{}d\Bigr{(}\frac{1}{N}\sum\delta_{\lambda_{i}},\nu\boxplus\sigma_{t}\Bigr{)}\leq\delta\Bigr{\}}\cap\Bigr{\{}\frac{1}{N}\sum\lambda_{i}^{2}\leq L_{0}+1\Bigr{\}}\,.$
Fix $Y\in\Delta^{p+1}$ and $v_{1}\in E_{Y,\kappa,\varepsilon}$. We can write
$\int_{\mathcal{B}(\delta,\varepsilon,x)}\prod_{2\leq i<j}|\lambda_{i}-\lambda_{j}|e^{-\frac{N}{4t}\mathrm{Tr}((p_{v_{1}}^{\perp}(OD(0,\lambda)O^{T}-B)p_{v_{1}}^{\perp})^{2})}\mathrm{d}\lambda_{2}\cdots\mathrm{d}\lambda_{N}\\ \leq\int_{x-\varepsilon\leq\lambda_{2}\leq\cdots\leq\lambda_{N}}|\lambda_{i}-\lambda_{j}|e^{-\frac{N}{4t}\mathrm{Tr}((p_{v_{1}}^{\perp}(OD(0,\lambda)O^{T}-B_{N})p_{v_{1}}^{\perp})^{2})}\mathrm{d}\lambda_{2}\cdots\mathrm{d}\lambda_{N}.$
Let us remark that $OD(0,\lambda)O^{T}$ follows the law of a GOE of dimension $N-1$ and variance $t(N-1)/N$ so that
$\int\mathrm{d}\mathbb{P}_{v_{1}}(v_{2},\ldots,v_{N})\int_{\lambda_{2}\leq\cdots\leq\lambda_{N}}\prod_{2\leq i<j}|\lambda_{i}-\lambda_{j}|e^{-\frac{N}{4t}\mathrm{Tr}((p_{v_{1}}^{\perp}(OD(0,\lambda)O^{T}-B)p_{v_{1}}^{\perp})^{2})}\mathrm{d}\lambda_{2}\ldots\mathrm{d}\lambda_{N}=Z_{N-1}^{t\frac{N-1}{N}}\,.$
Moreover, $Z_{N}^{t}$ can be exactly computed by Selberg integral, see e.g [AGZ10, (2.5.11)], resulting in
$Z_{N}^{t}=N!\Bigr{(}\frac{2t}{N}\Bigr{)}^{\frac{N(N+1)}{4}}(2\pi)^{N}\prod_{j=0}^{N-1}\frac{\Gamma((j+1)/2)}{\Gamma(1/2)},$
from which the limit of $\frac{1}{N}\log\frac{Z_{N-1}^{t\frac{N-1}{N}}}{Z_{N}^{t}}$ is easily inferred and gives a contribution $C_{t}$ with $C_{t}$ as in (2.15).
By Lemma 2.4, the matrix $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$ has eigenvalues $\eta_{1},\ldots,\eta_{p}$ of multiplicities between $M_{k}-1$ and $M_{k}$ with $p$ possible outliers $\Lambda_{1}\leq\cdots\leq\Lambda_{p}$ with for each $k=2,\ldots,p$, $\Lambda_{k}\in[\eta_{k-1},\eta_{k}]$. Moreover, by the continuity of $\Lambda_{1}$ proven in Lemma 2.4, $\Lambda_{1}$ is at distance at most of order $\varepsilon$ from $\Phi(\Lambda,Y)$ if $v_{1}\in E_{Y,\kappa,\varepsilon^{\prime}}$ and $\kappa$ is small enough. We can bound $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$ from above by a matrix $B_{N}^{\prime}$ in $v_{1}^{\perp}$ with the same eigenspaces, the outlier $\Lambda_{1}$ and eigenvalues $\eta_{i}$ with multiplicity $M_{i}^{\prime}=M_{i}$ except for $\eta_{1}$ which has multiplicity $M_{1}^{\prime}=M_{1}-1$. We therefore get
$\frac{1}{Z_{N-1}^{t\frac{N-1}{N}}}\int_{x-\varepsilon\leq\lambda_{2}\leq\cdots\leq\lambda_{N}}|\lambda_{i}-\lambda_{j}|e^{-\frac{N}{4t}\mathrm{Tr}((p_{v_{1}}^{\perp}(OD(0,\lambda)O^{T}-B_{N})p_{v_{1}}^{\perp})^{2})}\mathrm{d}\lambda_{2}\ldots\mathrm{d}\lambda_{N}\qquad$
$\qquad\leq\mathbb{P}_{N-1,\Lambda_{1},\mathsf{M}^{\prime}}(\lambda_{1}(G_{N-1}+B_{N}^{\prime})\geq x-2\varepsilon),$
where we incorporated the error due to the small change in the variance (which can be removed after rescaling and boundedness of $B_{N}$) in an additional error of order $\varepsilon$ for $N$ large enough. Note that Theorem 1.1 and Lemma 2.4 imply that the smallest eigenvalue of $G_{N-1}+B_{N}^{\prime}$ converges towards $H_{\nu,t}(\Lambda_{1})\simeq H_{\nu,t}(\Phi(\Lambda,Y))$ if $\Phi(\Lambda,Y)<\omega_{\nu,t}(\ell_{\nu,t})$ and towards $\ell_{\nu,t}$ otherwise. We therefore have three cases:
$\bullet$ If $\Phi(\Lambda,Y)\geq\omega_{\nu,t}(\ell_{\nu,t})$, then the smallest eigenvalue of $G_{N-1}+B_{N}^{\prime}$ converges towards $\ell_{\nu,t}$ and $\mathbb{P}_{N-1,\Lambda_{1},\mathsf{M}^{\prime}}(\lambda_{1}(G_{N-1}+B_{N}^{\prime})\geq x-2\varepsilon)$ goes to one as soon as $x<\ell_{\nu,t}$ and $\varepsilon$ is small enough.
$\bullet$ If $x\leq H_{\nu,t}(\Phi(\Lambda,Y))$ and $\Phi(\Lambda,Y)\leq\omega_{\nu,t}(\ell_{\nu,t})$, then the smallest eigenvalue of $G_{N-1}+B_{N}^{\prime}$ goes to $H_{\nu,t}(\Phi(\Lambda,Y))$ when $N$ goes to infinity and is therefore greater than $x-2\varepsilon$. Hence, we see that $\mathbb{P}_{N-1,\Lambda_{1},\mathsf{M}^{\prime}}(\lambda_{1}(G_{N-1}+B_{N}^{\prime})\geq x-2\varepsilon)$ goes to one as $N$ goes to infinity.
$\bullet$ If $x>H_{\nu,t}(\Phi(\Lambda,Y))$ and $\Phi(\Lambda,Y)\leq\omega_{\nu,t}(\ell_{\nu,t})$, then $\mathbb{P}_{N-1,\Lambda_{1},\mathsf{M}^{\prime}}(\lambda_{1}(G_{N-1}+B_{N}^{\prime})\geq x-2\varepsilon)$ goes to zero exponentially fast by large deviations. We can replace $B_{N}^{\prime}$ by another matrix $B_{N}^{\prime\prime}$ with the same eigenvalues except that $\Lambda_{1}$ is replaced by $\Phi(\Lambda,Y)$ up to an additional error of order $\varepsilon$. Observe that $\lambda_{1}(G_{N-1}+B_{N}^{\prime\prime})\leq\ell_{\nu,t}$ with overwhelming probability by (2.20). We can cut the compact set $[x-3\varepsilon,\ell_{\nu,t}]$ into a finite number $K$ of intervals of size $\gamma$ and replace $\Lambda_{1}$ by $\Phi(\Lambda,Y)$ up to a small error (an error of size $2\varepsilon$) to deduce that
$\mathbb{P}_{N-1,\Lambda_{1},\mathsf{M}^{\prime}}(\lambda_{1}(G_{N-1}+B_{N}^{\prime})\geq x-2\varepsilon)\leq\sum_{i=1}^{K}P_{N-1,\Phi(\Lambda,Y),\mathsf{M}^{\prime}}(x-3\varepsilon+i\gamma,\gamma)\,.$
According to the proof of Lemma B.1, there exists $c>0$ such that
$P_{N-1,\Phi(\Lambda,Y),\mathsf{M}^{\prime}}(x-3\varepsilon+i\gamma,\gamma)\leq 2N^{2}P_{N-1,\Phi(\Lambda,Y),\mathsf{M}^{\prime}}(x-3\varepsilon,\gamma)+2N^{2}e^{-cN^{2}}$
so that we obtain
$\mathbb{P}_{N-1,\Lambda_{1},\mathsf{M}^{\prime}}(\lambda_{1}(G_{N-1}+B_{N}^{\prime})\geq x-2\varepsilon)\leq 2KN^{2}P_{N-1,\Phi(\Lambda,Y),\mathsf{M}^{\prime}}(x-3\varepsilon,\gamma)+2N^{2}Ke^{-cN^{2}}\,.$
Hence, for $x>H_{\nu,t}(\Phi(\Lambda,Y))$ and $\Phi(\Lambda,Y)\leq\omega_{\nu,t}(\ell_{\nu,t})$, after taking the large $N$ limit and $\gamma$ to zero, we find
$\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}_{N-1,\Lambda_{1},\mathsf{M}^{\prime}}(\lambda_{1}(G_{N-1}+B_{N}^{\prime})\geq x-2\varepsilon)\leq-F^{+}_{\nu,t}(\Phi(\Lambda,Y),x-3\varepsilon)\,.$
To conclude we need to first take the supremum over $\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)$, so $\mathsf{M}^{\prime}\in\mathcal{B}^{N}_{\delta(\alpha)}$ before the large $N$ limit. We have to be careful because $\nu$ depends implicitly on $\mathsf{M}/N$ and $Y_{N}$ is not equal to $Y$ but is close to it. But the condition $\{\Phi(\Lambda,Y)\leq\omega_{\nu,t}(\ell_{\nu,t})\}\cap\{x\geq H_{\nu,t}(\Phi(\Lambda,Y))\}=\{\omega_{\nu,t}(x)\geq\Phi(\Lambda,Y)\}$ is continuous in $\nu$ (as $\omega_{\nu,t}(x)$ is) and $Y$. Moreover, $F^{+}_{\nu,t}(\Phi(\Lambda,Y),x-3\varepsilon)$ is lower semicontinuous in $Y$. The argument above is therefore valid uniformly on $\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)$ as long as we are in the interior of the set $\{\omega_{\nu,t}(x)\geq\Phi(\Lambda,Y)\}$ for $\delta$ small enough. We therefore conclude that for $\kappa$ small enough,
$\lim_{\delta\downarrow 0}\limsup_{N\to\infty}\sup_{v_{1}\in E_{Y_{k},\varepsilon^{\prime}}\atop M\in\mathcal{B}_{\delta}^{N}(\alpha)}\frac{1}{N}\log\Bigr{(}\frac{1}{Z_{N-1}^{t\frac{N-1}{N}}}\int_{x-\varepsilon\leq\lambda_{2}\leq\cdots\leq\lambda_{N}}|\lambda_{i}-\lambda_{j}|e^{-\frac{N}{4t}\mathrm{Tr}((p_{v_{1}}^{\perp}(OD(0,\lambda)O^{T}-B_{N})p_{v_{1}}^{\perp})^{2})}\mathrm{d}\lambda_{2}\ldots\mathrm{d}\lambda_{N}\Bigr{)}$
$\qquad\leq-F_{\nu,t}^{+}(\Phi(\Lambda,Y_{k}),x-3\varepsilon)1_{\omega_{\nu,t}(x)>\Phi(\Lambda,Y_{k})}+o(\kappa,\varepsilon).$
By Lemma 2.3, for every $k$,
$\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(E_{Y_{k}})\leq-I_{\nu}(Y_{k})+\varepsilon^{\prime}.$
We deduce from (2.23) that
$\lim_{\delta\downarrow 0}\limsup_{N\to\infty}\sup_{\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)}\frac{1}{N}\log P_{N,\Lambda,\mathsf{M}}(x,\varepsilon)\\ \leq\sup_{k}\Bigr{\{}-L_{\nu,t}^{\Lambda}(\lambda,Y_{k})-I_{\nu}(Y_{k})-F_{\nu,t}^{+}(\Phi(\Lambda,Y_{k}),x-3\varepsilon)1_{\omega_{\nu,t}(x)>\Phi(\Lambda,Y_{k})}\Bigr{\}}\\ +\int\log|\lambda-x|\mathrm{d}(\nu\boxplus\sigma_{t})(x)-\frac{\lambda^{2}}{4t}-C_{t}+o(\varepsilon^{\prime},\kappa,\varepsilon).\$
Letting $\varepsilon$ go to zero, we obtain
$\lim_{\varepsilon\downarrow 0}\limsup_{N\to\infty}\sup_{\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)}\frac{1}{N}\log P_{N,\Lambda,\mathsf{M}}(x,\varepsilon)\\ \leq\sup_{k}\Bigr{\{}-K_{\nu,t}(\Lambda,x,Y_{k})-F_{\nu,t}^{+}(\Phi(\Lambda,Y_{k}),x^{-})1_{\omega_{\nu,t}(x)\geq\Phi(\Lambda,Y_{k})}\Bigr{\}}+o(\varepsilon^{\prime},\kappa)\\ \leq\sup_{Y\in\Delta}\Bigr{\{}-K_{\nu,t}(\Lambda,x,Y)-F_{\nu,t}^{+}(\Phi(\Lambda,Y),x^{-})1_{\omega_{\nu,t}(x)>\Phi(\Lambda,Y)}\Bigr{\}}+o(\varepsilon^{\prime},\kappa).$
We can finally let $\varepsilon^{\prime}$ going to zero. Now, we take the supremum over $\Lambda\in[\Lambda^{\prime}-\varepsilon,\Lambda^{\prime}+\varepsilon]$ on both sides. In the right-hand side it can be interchanged with the supremum over $Y$. The map $J_{\nu,t}^{\Lambda}(\lambda,Y)=L_{\nu,t}^{\Lambda}(\lambda,Y)+I_{\nu}(Y)$ is uniformly continuous in $\Lambda$. Moreover
$\lim_{\varepsilon\downarrow 0}\inf_{\Lambda\in[\Lambda^{\prime}-\varepsilon,\Lambda^{\prime}+\varepsilon]}F_{\nu,t}^{+}(\Phi(\Lambda,Y),x^{-})1_{\omega_{\nu,t}(x)>\Phi(\Lambda,Y)}\geq F_{\nu,t}^{+}(\Phi(\Lambda^{\prime},Y),x^{-})1_{\omega_{\nu,t}(x)>\Phi(\Lambda^{\prime},Y)},$
by lower semi-continuity of $F_{\nu,t}^{+}(.,x^{-})$, and the continuity of $\Lambda\mapsto\Phi(\Lambda,Y)$. Note that the boundary point when $\omega_{\nu,t}(x)=\Phi(\Lambda,Y)$ is also such that $\omega_{\nu,t}(x^{-})=\Phi(\Lambda,Y)$ by continuity of $\omega_{\nu,t}$ and then since $F_{\nu}^{+}(\omega_{\nu,t}(x),x)$ vanishes at every $x$, so that the above quantity vanishes at this point. We therefore conclude that (2.19) holds.
2.5. A functional inequality for the large deviation lower bound
For every $\Lambda\leq\ell_{\nu}$ and $\lambda\leq\ell_{\nu,t}$, define
$F_{\nu}^{-}(\Lambda,\lambda):=-\liminf_{{\Lambda^{\prime}\to\Lambda}}\liminf_{\varepsilon\to 0}\lim_{\delta\downarrow 0}\liminf_{N\to\infty}\inf_{\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)}\frac{1}{N}\log P_{N,\Lambda^{\prime},\mathsf{M}}(\lambda,\varepsilon).$ (2.24)
Denoting $f(x^{+})=\lim_{\varepsilon\downarrow 0}f(x+\varepsilon)$, our goal is to prove the following:
Lemma 2.7 (Functional lower bound).
Let $\Lambda\leq\ell_{\nu}$. For every $\lambda<\ell_{\nu,t}$, we have
$F_{\nu,t}^{-}(\Lambda,\lambda)\leq\inf_{Y\in\Delta^{p+1}}\Bigr{\{}K_{\nu,t}(\Lambda,\lambda,Y)+F_{\nu,t}^{-}(\Phi(\Lambda,Y),\lambda^{+})1_{\omega_{\nu,t}(\lambda)\geq\Phi(\Lambda,Y)}\Bigr{\}}.$ (2.25)
Moreover, $0\leq F_{\nu,t}^{+}\leq F_{\nu,t}^{-}$. Furthermore, $\lambda\mapsto F_{\nu,t}^{-}(\Lambda,\lambda)$ decreases on $(-\infty,{\ell^{\Lambda}_{\nu,t})}$ and increases on ${(\ell^{\Lambda}_{\nu,t},\ell_{\nu,t})}$, while it vanishes at ${\ell^{\Lambda}_{\nu,t}}$. It is bounded from below by $c(\lambda-\ell^{\Lambda}_{\nu,t})^{2}$ for some $c>0$.
Proof.
The only remaining task is to prove (2.25), as the rest of the lemma is established in Lemma B.2. Let $x\in(-\infty,\ell_{\nu,t})$ and $\varepsilon>0$.
Again $B_{N}$ has eigenvalues $\eta_{i}$ with multiplicity $M_{i}$ and $\Lambda^{\prime}$ with multiplicity $1$. Let $Y\in\Delta^{p+1}$ and $\kappa,\varepsilon^{\prime}>0$. Letting $E_{Y,\kappa,\varepsilon^{\prime}}$ be as in (2.22), one can bound $P_{N,\Lambda^{\prime},\mathsf{M}}(x,\varepsilon)$ from below by
$P_{N,\Lambda^{\prime},\mathsf{M}}(x,\varepsilon)\geq\frac{1}{Z_{N}^{t}}\int_{E_{Y,\kappa,\varepsilon^{\prime}}}\mathrm{d}v_{1}\int\mathrm{d}\mathbb{P}_{v_{1}}(v_{2},\ldots,v_{N})\\ \int_{\mathcal{B}(\delta,\varepsilon,x)}e^{-NL_{\nu,t}^{\Lambda^{\prime}}(\lambda_{1},\langle v_{a},w_{0}\rangle^{2},...)-\frac{N}{4t}\lambda_{1}^{2}}\prod_{1\leq i<j}|\lambda_{i}-\lambda_{j}|e^{-\frac{N}{4t}\mathrm{Tr}((p_{v_{1}}^{\perp}(OD(0,\lambda)O^{T}-B)p_{v_{1}}^{\perp})^{2})}\mathrm{d}\lambda_{1}\cdots\mathrm{d}\lambda_{N},$ (2.26)
where $B(\delta,\varepsilon,x):=\{|\lambda_{1}-x|\leq\varepsilon\}\cap C(\delta,\varepsilon,x)$ with, if $\nu=\sum\alpha_{i}\delta_{\eta_{i}}$,
$C(\delta,\varepsilon,x):=\{x+2\varepsilon<\lambda_{2}<\cdots<\lambda_{N}\}\cap\Bigr{\{}d\Bigr{(}\frac{1}{N}\sum_{i=1}^{N}\delta_{\lambda_{i}},\nu\boxplus\sigma_{t}\Bigr{)}\leq 2\delta\Bigr{\}}\cap\Bigr{\{}\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}^{2}\leq L_{0}+1\Bigr{\}}.$
Using the continuity of $L_{\nu,t}^{\Lambda^{\prime}}$ and the fact that $\lambda_{1}$ is at distance at least $\varepsilon$ from the rest of the spectrum (if $x+2\varepsilon<\ell_{\nu,t}$) so that
$\frac{1}{N}\sum_{i=2}^{N}\log|\lambda_{1}-\lambda_{i}|=\int\log|\lambda_{1}-x|\mathrm{d}(\nu\boxplus\sigma_{t})(x)+O(\delta/\varepsilon)\quad\text{on $B(\delta,\varepsilon,x)$},$
we get
$P_{N,\Lambda,\mathsf{M}}(x,\varepsilon)\geq\frac{2\varepsilon Z^{t(1-N^{-1})}_{N-1}}{Z^{t}_{N}}e^{N(\int\log|x-\lambda|\mathrm{d}(\sigma_{t}\boxplus\nu)(\lambda)-L_{\nu,t}^{\Lambda}(x,Y)+o(\delta,\epsilon)-\frac{1}{4t}x^{2}+o(\varepsilon,\delta))}\\ \times\int_{E_{Y,\kappa,\varepsilon^{\prime}}}\mathrm{d}v_{1}\mathbb{P}(\lambda(G_{N-1}+p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp})\in C(\delta,\varepsilon,x)).$
Moreover, recall that $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$, as a $(N-1)\times(N-1)$ matrix in $v_{1}^{\perp}$, has a unique outlier in $[\Lambda,\ell_{\nu}]$, which is close to $\Phi(\Lambda^{\prime},Y)$, and eigenvalues $\eta_{i}$ with multiplicities in $\{N_{i},N_{i}-1\}$. Let $B_{N}^{\prime}$ be the matrix with outlier $\Phi(\Lambda^{\prime},Y)$ and eigenvalues $\eta_{i}$ with multiplicity $M_{i}^{\prime}=M_{i}$ except for $\eta_{p}$ with multiplicity $M_{p}^{\prime}=M_{p}-1$ (with the same eigenspaces as $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$). Then $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}\geq B_{N}^{\prime}$ so that
$\mathbb{P}(\lambda_{1}(G_{N-1}+p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp})\in C(\delta,\varepsilon,x))$$\geq$ $\mathbb{P}(\lambda_{1}(G_{N-1}+p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp})\geq x+2\varepsilon)$
$-\mathbb{P}(d\Bigr{(}\frac{1}{N}\sum_{i=1}^{N}\delta_{\lambda_{i}},\nu\boxplus\sigma_{t}\Bigr{)}\geq\delta)-\mathbb{P}(\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}^{2}\geq L_{0}+1)$
$\geq$ $\mathbb{P}(\lambda_{1}(G_{N-1}+B_{N}^{\prime})\geq x+2\varepsilon)-e^{-cN^{2}},$
where we finally used (2.20) and (2.21) and took some $\delta\leq\varepsilon$. As for the proof of the upper bound, the right-hand side goes to $1$ for $\varepsilon$ small enough except when $\omega_{\nu,t}(x)>\Phi(\Lambda,Y)$. Notice that $B_{N}^{\prime}$ satisfy the same assumptions as $B_{N}$ except that its outlier is now going to $\Phi(\Lambda^{\prime},Y)$ (when $\kappa$ goes to zero) so that, for every $\Lambda<\ell_{\nu}$, for every $Y\in\Delta^{p+1}$,
$\lim_{\delta\to 0}\liminf_{N\to\infty}\inf_{\mathsf{M}\in B_{\delta}^{N}(\alpha)}\frac{1}{N}\log P_{N,\Lambda^{\prime},\mathsf{M}}(\lambda,\varepsilon)\geq-K_{\nu,t}(\Lambda^{\prime},\lambda,Y)-F_{\nu,t}^{-}(\Phi(\Lambda^{\prime},Y),x+3\varepsilon)+o(\varepsilon,\kappa,\varepsilon^{\prime}),$
where we took $\delta$ to zero. We can then let $\varepsilon$ going to zero and take the supremum over all possible $Y$. We can finally take the infimum over $\Lambda\in[\Lambda^{\prime}-\varepsilon,\Lambda^{\prime}+\varepsilon]$ and use the upper semi-continuity of $F_{\nu,t}^{-}$. This completes the proof of (2.25) as soon as $x+3\varepsilon<\ell_{\nu,t}$, which is true for $\varepsilon$ small enough as long as $x<\ell_{\nu,t}$. ∎
2.6. Solving the functional equation in the region $\omega_{\nu,t}(\lambda)\leq\Lambda$
Recall that for $\Lambda\leq\ell_{\nu}$ and $\lambda\leq\ell_{\nu,t}$,
$\gamma(\Lambda,\lambda):=\begin{cases}\Lambda&\text{if $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$, or $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda\leq\rho_{\nu,t}^{\Lambda}$}\\ \omega^{*}_{\nu,t}(\lambda)&\text{if $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$and $\lambda\geq\rho_{\nu,t}^{\Lambda}$},\end{cases}$ (2.27)
where $\omega^{*}_{\nu,t}(\lambda)$ stands for the unique solution $\gamma$ in $[\omega_{\nu,t}(\ell_{\nu,t}),\inf\eta_{k})$ of the equation $H_{\nu,t}(\gamma)=\lambda$. Also define
$I_{\nu,t}^{\Lambda}(\lambda,\gamma):=\frac{1}{2}(S_{\nu}(\omega_{\nu,t}(\lambda))-S_{\nu}(\gamma))+\frac{1}{4t}((\lambda-\gamma)^{2}-(\lambda-\omega_{\nu,t}(\lambda))^{2}).$ (2.28)
and set ${I_{\nu,t}^{\Lambda}(\lambda):=I_{\nu,t}^{\Lambda}(\lambda,\gamma(\Lambda,\lambda))}$. We first prove that $F_{\nu,t}^{+}$ and $F_{\nu,t}^{-}$ equal $I_{\nu,t}^{\Lambda}$ in the simplest case where the functional equations (2.19) and (2.25) are such that the infimum is taken when the term depending on $F_{\nu,t}^{+}$ or $F_{\nu,t}^{-}$ vanishes.
Lemma 2.8.
Let $\Lambda\leq\ell_{\nu}$ and $\lambda<\ell_{\nu,t}$. We have for every $\lambda<\ell_{\nu,t}$ and $\Lambda\leq\ell_{\nu}$,
$F_{\nu,t}^{+}(\Lambda,\lambda)\leq F_{\nu,t}^{-}(\Lambda,\lambda)\leq I_{\nu,t}^{\Lambda}(\lambda).$ (2.29)
Moreover, if $\Lambda\geq\omega_{\nu,t}(\lambda)$ or $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda\leq\rho_{\nu,t}^{\Lambda}$, then
$F_{\nu,t}^{-}(\Lambda,\lambda)=F_{\nu,t}^{+}(\Lambda,\lambda)=I_{\nu,t}^{\Lambda}(\lambda)=\inf K_{\nu,t}(\Lambda,\lambda,\cdot).$ (2.30)
A key point to prove Lemma 2.8 is to study the minimizers of $K_{\nu,t}(\Lambda,\lambda,.)$, namely the minimizers of the function $J_{\nu,t}^{\Lambda}(\lambda,.)$ defined in (2.17).
Lemma 2.9.
Fix $\lambda\in(-\infty,\ell_{\nu,t}]$.
  1. (1)
    For every $\Lambda\leq\ell_{\nu}$ and with either $\gamma=\omega_{\nu,t}^{*}(\lambda)$ or $\gamma=\Lambda$, let $Y(\gamma):=(y_{0}(\gamma),y_{1}(\gamma),\ldots,y_{p}(\gamma))$ be given by
    $y_{i}(\gamma):=\frac{t}{(\eta_{i}-\omega_{\nu,t}(\lambda))(\eta_{i}-\gamma)},\quad{\text{for each $1\leq i\leq p$}},$ (2.31)
    and $y_{0}(\gamma)=1-\sum_{i=1}^{p}y_{i}(\gamma)$. The local minimizers of $J^{\Lambda}_{\nu,t}(\lambda,.)$ are included in $\{Y(\omega_{\nu,t}^{*}(\lambda)),Y(\Lambda)\}$ whenever $y_{0}(\Lambda)\geq 0$ or equal to $\{Y(\omega_{\nu,t}^{*}(\lambda))\}$ whenever $y_{0}(\Lambda)<0$. Moreover, for $\gamma=\omega_{\nu,t}^{*}(\lambda)$ and $\gamma=\Lambda$, we have
    $K_{\nu,t}(\Lambda,\lambda,Y(\gamma))=I_{\nu,t}^{\Lambda}(\lambda,\gamma).$ (2.32)
  2. (2)
    The function $J_{\nu,t}^{\Lambda}(\lambda,\cdot)$ admits a unique minimizer $(y_{0},Y)\in\Delta^{p+1}$ in the following cases:
    • If $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda\leq\rho_{\nu,t}^{\Lambda}$, or $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda\leq\rho_{\nu,t}^{\Lambda}$, or $\Lambda<\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda=\ell_{\nu,t}$, then the minimizer is given by $Y(\Lambda)$ and
      $y_{0}(\Lambda)=\begin{cases}1+tG_{\nu}^{\prime}(\Lambda)&\text{if $\lambda=\rho_{\nu,t}^{\Lambda}$}\\ \frac{\lambda-\rho_{\nu,t}^{\Lambda}}{\omega_{\nu,t}(\lambda)-\Lambda}&\text{if $\lambda\neq\rho_{\nu,t}^{\Lambda}$}.\end{cases}$
    • If $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda\in(\rho_{\nu,t}^{\Lambda},\ell_{\nu,t}]$, then the unique minimizer is $Y({\omega_{\nu,t}^{*}(\lambda)})$ with $y_{0}(\omega_{\nu,t}^{*}(\lambda))$ vanishing.
Proof.
Step 1: existence of a minimizer. By Cauchy-Schwarz inequality, since $t\geq 0$,
$J_{\nu,t}^{\Lambda}(\lambda,Y)\geq-\frac{\lambda}{2t}\Bigr{(}\Lambda y_{0}+\sum_{k=1}^{p}\eta_{k}y_{k}\Bigr{)}+\frac{1}{4t}\Bigr{(}\sum_{k=1}^{p}\eta_{k}^{2}y_{k}+\Lambda^{2}y_{0}\Bigr{)}-\frac{1}{2}\sum_{k=1}^{p}\alpha_{k}\log(y_{k}/\alpha_{k})$
Thus, since the last term is non-negative by Jensen’s inequality, one can see by completing the square that
${J_{\nu,t}^{\Lambda}(\lambda,Y)}\geq\frac{1}{4t}\Bigr{(}\sum_{k=1}^{p}(\eta_{k}-\lambda)^{2}y_{k}+(\Lambda-\lambda)^{2}y_{0}\Bigr{)}-\frac{1}{4t}\lambda^{2}\geq-\frac{1}{4t}\lambda^{2},$
is uniformly bounded from below on $\Delta^{p+1}$. It is well-known that $Y\in(1-y_{0})\Delta^{p}\mapsto I_{\nu}(Y)$ is lower semicontinuous. Moreover,
$Y\in(1-y_{0})\Delta^{p}\mapsto\sum_{k=1}^{p}\eta_{k}y_{k}\quad\text{ and}\quad Y\in({1-y_{0}})\Delta^{p}\mapsto\sum_{k=1}^{p}\eta_{k}^{2}y_{k}$
are continuous. Thus, $J_{\nu,t}^{\Lambda}(\lambda,\cdot)$ is lower semicontinuous and since it is bounded from below, it achieves its minimal value. The function $J_{\nu,t}^{\Lambda}(\lambda,\cdot)$ can be differentiated and its critical points can be described.
Step 2: interior critical points. Let $Y\in\Delta^{p+1}$ be a critical point of $J_{\nu,t}^{\Lambda}(\lambda,\cdot)$ such that $y_{0}>0$. By variational calculus, there exist constants $R_{1},R_{2}$ such that for every $k\in\{1,\ldots,p\}$
$-\lambda\eta_{k}+\eta_{k}^{2}-R_{1}\eta_{k}-\frac{t}{y_{k}}=R_{2}$
and
$-\lambda\Lambda+\Lambda^{2}-R_{1}\Lambda=R_{2}.$
Therefore, for each $k=1,\ldots,p$,
$y_{k}=\frac{t}{\eta_{k}^{2}-\lambda\eta_{k}-R_{1}\eta_{k}-R_{2}},$
where $R_{1}$ and $R_{2}$ are such that
$\begin{cases}y_{0}+\sum_{i=1}^{p}y_{i}=1\\ \Lambda y_{0}+\sum_{i=1}^{k}\eta_{k}y_{k}=R_{1}\\ -\lambda\Lambda+\Lambda^{2}-R_{1}\Lambda=R_{2}.\end{cases}$ (2.33)
One may write $Y$ in the form
$y_{k}=\frac{t}{(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}{\quad\text{for each$k=1,\ldots,p$}}$ (2.34)
with constants $\gamma_{1},\gamma_{2}\in\mathbb{C}$ satisfying $\gamma_{1}+\gamma_{2}=\lambda+R_{1}$ and $\gamma_{1}\gamma_{2}=-R_{2}$. The last equation in (2.33) therefore gives
$(\Lambda-\gamma_{1})(\Lambda-\gamma_{2})=0.$
It follows that $\gamma_{1}=\Lambda$ or $\gamma_{2}=\Lambda$ and $\gamma_{1},\gamma_{2}\in\mathbb{R}$. Thus, for each $k=1,\ldots,p$,
$y_{k}=\frac{t}{(\eta_{k}-\Lambda)(\eta_{k}-\alpha)},$
for some $\alpha\in\mathbb{R}$. Note that $y_{k}$ is nonnegative for every $k$ such that $\alpha$ belongs to $(-\infty,\ell_{\nu})$ since $\Lambda\leq\ell_{\nu}$. If $\alpha\neq\Lambda$, the first equation of (2.33) gives
$\sum_{k=1}^{p}y_{k}=\frac{t}{\Lambda-\alpha}(G_{\nu}(\alpha)-G_{\nu}(\Lambda))=1-y_{0},$
otherwise if $\alpha=\Lambda$, we have
$1+tG_{\nu}^{\prime}(\Lambda)=y_{0}.$
Furthermore the second equation of (2.33) gives
$H_{\nu,t}(\alpha)=\lambda,\quad{\alpha<\ell_{\nu}}.$ (2.35)
By Remark A.2, whenever $\lambda<\ell_{\nu,t}$, the above equation has two distinct solutions $\gamma_{1}<\gamma_{2}$ in $(-\infty,\ell_{\nu})$ and moreover $\gamma_{1}=\omega_{\nu,t}(\lambda)$ and $\gamma_{2}=\omega^{*}_{\nu,t}(\lambda)$. If $\lambda=\ell_{\nu,t}$, then Equation (2.35) has a unique solution $\alpha=\omega_{\nu,t}(\ell_{\nu,t})$.
In the rest of the proof, we let $\gamma_{1}\leq\gamma_{2}$ be the two solutions of (2.35). For $\gamma\in\{\gamma_{1},\gamma_{2}\}$, let ${y(\gamma)}$ be given by
${y_{k}(\gamma)}:=\frac{t}{(\eta_{k}-\gamma)(\eta_{k}-\Lambda)}{\quad\text{for each $k=1,\ldots,p$}}.$ (2.36)
Let also
$y_{0}(\gamma):=\begin{cases}1+tG_{\mu}^{\prime}(\Lambda)&\text{if $\gamma=\Lambda$}\\ \frac{H_{\nu,t}(\gamma)-H_{\nu,t}(\Lambda)}{\gamma-\Lambda}&\text{if $\gamma\neq\Lambda$}.\end{cases}$ (2.37)
Observe that if $\Lambda>\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda>\rho_{\nu,t}^{\Lambda}$, then for $\gamma\in\{\gamma_{1},\gamma_{2}\}$, $y_{0}(\gamma)<0$, showing that $Y({\gamma})$ is not a critical point in $\Delta^{p+1}$. Let us now assume that $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$ or $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda\leq\rho_{\nu,t}^{\Lambda}$. We check whether $y_{0}({\gamma})$ belongs to $[0,1]$. Note that we always have $y_{0}({\gamma})<1$. Moreover, in the case $\gamma=\Lambda$,
$y_{0}=1+tG^{\prime}_{\nu}(\Lambda),$
which is strictly positive if and only if $\Lambda<\omega_{\nu,t}(\ell_{\nu,t})$. Therefore if $\lambda=\rho_{\nu,t}^{\Lambda}=H_{\nu,t}(\Lambda)$ (giving $\gamma_{1}=\gamma_{2}=\Lambda$) and $\Lambda<\omega_{\nu,t}(\ell_{\nu,t})$, then $y_{0}\in(0,1)$. If ${\Lambda\notin\{\gamma_{1},\gamma_{2}\}}$, then
$y_{0}({\gamma})=\frac{\lambda-\rho_{\nu,t}^{\Lambda}}{\gamma-\Lambda}$
and $y_{0}({\gamma})$ belongs to $(0,1)$ only if
  • $\Lambda<\omega_{\nu,t}(\ell_{\nu,t})$ (since $H_{\nu,t}$ is increasing on $(-\infty,\omega_{\nu,t}(\ell_{\nu,t}))$),
  • or $\Lambda>\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda<\rho_{\nu,t}^{\Lambda}$.
Hence, in this range of parameters $Y({\gamma_{i}})_{i=1,2}$ are critical points in $\Delta^{p+1}$, but none of them belong to $\Delta^{p+1}$ otherwise. In the following we compute the value of $J_{\nu,t}^{\Lambda}(\lambda,\cdot)$ for each of these two solutions in order to select the minimizer. Let $\gamma\in\{\gamma_{1},\gamma_{2}\}$. One may compute
$X:=\Lambda y_{0}+\sum_{i=1}^{p}\eta_{i}y_{i}=\Lambda-tG_{\mu}(\gamma)=\Lambda+\gamma-\lambda,$ (2.38)
where we have used the fact that $H_{\nu,t}(\gamma)=\lambda$ in the last equality. Moreover,
$\begin{split}\Lambda^{2}y_{0}+\sum_{k=1}^{p}\eta_{k}^{2}y_{k}&=\Lambda^{2}y_{0}+t+(\gamma+\Lambda)\sum\eta_{k}y_{k}-\gamma\Lambda(1-y_{0})\\ &=\Lambda^{2}y_{0}+t+(\gamma+\Lambda)X-\gamma\Lambda(1-y_{0})-(\gamma+\Lambda)\Lambda y_{0}\\ &=t+(\gamma+\Lambda)X-\gamma\Lambda.\end{split}$
Thus
$\begin{split}J_{\nu,t}^{\Lambda}(\lambda,Y(\gamma))&=\frac{1}{4t}(-2\lambda-X+2(\gamma+\Lambda))X+\frac{1}{2}-\frac{\gamma\Lambda}{2t}-\frac{1}{2}(S_{\nu}(\Lambda)+S_{\nu}(\gamma))-\frac{1}{2}\ln t\\ &=\frac{1}{4t}(\gamma+\Lambda-\lambda)^{2}-2\gamma\Lambda)+\frac{1}{2}-\frac{1}{2}(S_{\nu}(\Lambda)+S_{\nu}(\gamma))-\frac{1}{2}\ln t\\ &=\frac{1}{4t}(\gamma^{2}+\Lambda^{2}+\lambda^{2}-2\gamma\lambda-2\Lambda\lambda)+\frac{1}{2}-\frac{1}{2}(S_{\nu}(\Lambda)+S_{\nu}(\gamma))-\frac{1}{2}\ln t.\end{split}$ (2.39)
Define the map
$\phi:\gamma\in(-\infty,\ell_{\nu})\mapsto\frac{1}{4t}(\gamma^{2}+\Lambda^{2}+\lambda^{2}-2\gamma\lambda-2\Lambda\lambda)+\frac{1}{2}-\frac{1}{2}(S_{\nu}(\Lambda)+S_{\nu}(\gamma))-\frac{1}{2}\ln t.$
One can compute that for all $\gamma\in(-\infty,\ell_{\nu})$,
$\phi^{\prime}(\gamma)=\frac{1}{2t}(H_{\nu,t}(\gamma)-\lambda).$
Therefore $\phi$ is strictly increasing on $[\gamma_{1},\gamma_{2}]$ and therefore since we assumed $\lambda\neq\ell_{\nu,t}$,
${J_{\nu,t}^{\Lambda}(\lambda,Y({\gamma_{1}}))=\phi(\gamma_{1})<\phi(\gamma_{2})=J_{\nu,t}^{\Lambda}(\lambda,Y({\gamma_{2}}))}.$
We deduce that when $\lambda\neq\ell_{\nu,t}$, $Y({\gamma_{2}})$ is not a minimizer of $J_{\nu,t}^{\Lambda}(\lambda,\cdot)$.
Step 3: study of the critical points on the boundary. Let $Y\in\Delta^{p}$ be a critical point of $J_{\nu,t}^{\Lambda}(\cdot,0,Y)$. By variational calculus, there exist $R_{1},R_{2}$ such that
$-\lambda\eta_{k}+\eta_{k}^{2}-R_{1}\eta_{k}-\frac{t}{y_{k}}=R_{2},{\quad\text{for each $1\leq k\leq p$}},$
where $R_{1}=\sum\eta_{k}y_{k}$. Therefore
$y_{k}=\frac{t}{(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})},{\quad\text{for each $1\leq k\leq p$}}$ (2.40)
where $\gamma_{1},\gamma_{2}\in\mathbb{C}$ satisfy $\gamma_{1}+\gamma_{2}=\lambda+R_{1}$ and $\gamma_{1}\gamma_{2}=-R_{2}.$ The equation $\sum y_{k}=1$ gives
$H_{\nu,t}(\gamma_{1})=H_{\nu,t}(\gamma_{2}).$ (2.41)
Moreover
$\sum_{k=1}^{p}\eta_{k}y_{k}=-tG_{\nu}(\gamma_{2})+\gamma_{1}=R_{1}=\gamma_{1}+\gamma_{2}-\lambda.$
Inserting (2.41), this shows that
$H_{\nu,t}(\gamma_{1})=H_{\nu,t}(\gamma_{2})=\lambda.$
Therefore $\gamma_{1}=\omega_{\nu,t}(\lambda)$ and $\gamma_{2}=\omega_{\nu,t}^{*}(\lambda)$. Let us now compute $J_{\nu,t}^{\Lambda}(\lambda,0,Y)$. Let $X:=\sum\eta_{k}y_{k}=\gamma_{1}+\gamma_{2}-\lambda$. We have
${\sum_{k=1}^{p}}\eta_{k}^{2}y_{k}=t+(\gamma_{1}+\gamma_{2})X-\gamma_{1}\gamma_{2}.$
Therefore we get
$\begin{split}J_{\nu,t}^{\Lambda}(\lambda,0,Y)&=\frac{1}{4t}(-2\lambda-X+2(\gamma_{1}+\gamma_{2}))X+\frac{1}{2}-\frac{\gamma_{1}\gamma_{2}}{2t}-\frac{1}{2}(S_{\nu}(\gamma_{1})+S_{\nu}(\gamma_{2}))-\frac{1}{2}\ln t\\ &=\frac{1}{4t}(\gamma_{1}^{2}+\gamma_{2}^{2}+\lambda^{2}-2\gamma_{1}\lambda-2\gamma_{2}\lambda)+\frac{1}{2}-\frac{1}{2}(S_{\nu}(\gamma_{1})+S_{\nu}(\gamma_{2}))-\frac{1}{2}\ln t.\end{split}$ (2.42)
Step 4: Identification of the minimizers.
$\bullet$ If $\Lambda>\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda>\rho_{\nu,t}^{\Lambda}$, then by Step 2, there is no critical point with $y_{0}>0$ and the only minimizer is therefore given by $Y({\omega^{*}_{\nu,t}(\lambda)})$ where ${Y(\gamma)}$ is as defined in (2.31).
$\bullet$ If $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda\leq\rho_{\nu,t}^{\Lambda}$, or $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda\leq\rho_{\nu,t}^{\Lambda}$, then the minimizer is either $Y(\gamma_{1})$ defined in (2.36) or $(0,Y)$ defined in (2.40). But
$J_{\nu,t}^{\Lambda}(\lambda,0,Y)=\frac{1}{4t}(\gamma_{1}^{2}+\gamma_{2}^{2}+\lambda^{2}-2\gamma_{1}\lambda-2\gamma_{2}\lambda)+\frac{1}{2}-\frac{1}{2}(S_{\nu}(\gamma_{1})+S_{\nu}(\gamma_{2}))-\frac{1}{2}\ln t,$
while
$J_{\nu,t}^{\Lambda}(\lambda,Y(\gamma_{1}))=\frac{1}{4t}(\gamma_{1}^{2}+\Lambda^{2}+\lambda^{2}-2\gamma_{1}\lambda-2\Lambda\lambda)+\frac{1}{2}-\frac{1}{2}(S_{\nu}(\gamma_{1})+S_{\nu}(\Lambda))-\frac{1}{2}\ln t{.}$
Define
$\psi:\gamma\in(-\infty,\ell_{\nu}]\mapsto\frac{1}{4t}(\gamma_{1}^{2}+\gamma^{2}+\lambda^{2}-2\gamma_{1}\lambda-2\gamma\lambda)+\frac{1}{2}-\frac{1}{2}(S_{\nu}(\gamma_{1})+S_{\nu}(\gamma))-\frac{1}{2}\ln t$
Note that for all $\gamma\in(-\infty,\ell_{\nu})$,
$\psi^{\prime}(\gamma)=\frac{1}{2t}(H_{\nu,t}(\gamma)-\lambda).$
Hence $\psi$ is strictly increasing on $[\gamma_{1},\gamma_{2}]$. Recall that $\Lambda\in[\gamma_{1},\gamma_{2}]$. Therefore if $\Lambda<\gamma_{2}$, then
$J_{\nu,t}^{\Lambda}(\lambda,0,Y)=\psi(\gamma_{2})>\psi(\Lambda)=J_{\nu,t}^{\Lambda}(\lambda,Y(\gamma_{1}))$
which implies that $Y({\gamma_{1}})$ is the unique minimizer of $J_{\nu,t}^{\Lambda}(\lambda,\cdot)$. If $\Lambda=\gamma_{2}$, then $Y(\gamma_{1})$ is again the unique minimizer of $J_{\nu,t}^{\Lambda}(\lambda,\cdot)$. We can also conclude in the case where $\Lambda<\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda=\ell_{\nu,t}$. In this case, one can notice that $\gamma_{2}=\gamma_{1}=\omega_{\nu,t}(\ell_{\nu,t})$. Moreover $\psi$ is strictly increasing on $(-\infty,\gamma_{1})$. Therefore
$J_{\nu,t}^{\Lambda}(\lambda,0,Y)=\psi(\gamma_{1})>J_{\nu,t}^{\Lambda}(\lambda,Y(\Lambda))=\psi(\Lambda).$
Therefore $Y(\Lambda)$ is the unique minimizer of $J_{\nu,t}^{\Lambda}(\lambda,\cdot)$.
Step 5: Proof of (2.32). Let $Y(\gamma)$ be as in (2.31). Recall that by (2.39) and (2.42), for $\gamma\in\{\omega_{\nu,t}^{*}(\lambda),\Lambda\}$,
$J_{\nu,t}^{\Lambda}(\lambda,Y(\gamma))=\frac{1}{4t}(\gamma_{1}^{2}+\gamma^{2}+\lambda^{2}-2\gamma_{1}\lambda-2\gamma\lambda)+\frac{1}{2}-\frac{1}{2}(S_{\nu}(\gamma_{1})+S_{\nu}(\gamma))-\frac{1}{2}\ln t.$
By Lemma A.4,
$-\int\log|\lambda-x|\mathrm{d}(\mu\boxplus\sigma_{t})(x)=S_{\mu\boxplus\sigma_{t}}(\lambda)=S_{\nu}(\gamma_{1})-\frac{1}{2t}(\gamma_{1}^{2}+\lambda^{2}-2\lambda\gamma_{1}).$
Therefore, combining this with (2.39) and recalling that $C_{t}=\frac{1}{2}-\frac{1}{2}\ln t$, we obtain
$K_{\nu,t}(\Lambda,\lambda,Y(\gamma))$ $=$ $\frac{1}{4t}(\gamma_{1}^{2}+\gamma^{2}+\lambda^{2}-2\gamma_{1}\lambda-2\gamma\lambda)-\frac{1}{2}(S_{\nu}(\gamma_{1})+S_{\nu}(\gamma))+S_{\nu}(\gamma_{1})$
$-\frac{1}{4t}(2\gamma_{1}^{2}+2\lambda^{2}-4\lambda\gamma_{1}-\lambda^{2})$
$=$ $\frac{1}{2}(S_{\nu}(\gamma_{1})-S_{\nu}(\gamma))+\frac{1}{4t}(\gamma^{2}-\gamma_{1}^{2}-2\gamma\lambda+2\gamma_{1}\lambda)$
$=$ $\frac{1}{2}(S_{\nu}(\gamma_{1})-S_{\nu}(\gamma))+\frac{1}{4t}((\gamma-\lambda)^{2}-(\gamma_{1}-\lambda)^{2})=I_{\nu,t}^{\Lambda}(\lambda,\gamma),$
which proves (2.32) by recalling (2.28). ∎
Before giving the proof of Lemma 2.8 we establish the following:
Lemma 2.10.
  1. (1)
    Let $\lambda\in(-\infty,\ell_{\nu,t}]$. If $\Lambda<\omega_{\nu,t}(\ell_{\nu,t})$, or $\Lambda>\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda<\rho_{\nu,t}^{\Lambda}$, let $Y(\Lambda)=(y_{0}(\Lambda),\ldots,y_{p}(\Lambda))\in\Delta^{p+1}$ be given by
    $y_{k}(\Lambda)=\frac{t}{(\eta_{k}-\omega_{\nu,t}(\lambda))(\eta_{k}-\Lambda)},\quad{\text{for each $k=1,\ldots,p$}}.$
    Then $y_{0}(\Lambda):=1-\sum_{k}y_{k}(\Lambda)$ is strictly positive. Moreover, $\Phi(\Lambda,Y(\Lambda))=\omega_{\nu,t}^{*}(\lambda)$.
  2. (2)
    Let $Y=(0,y_{1},\ldots,y_{p})\in\Delta^{p+1}$. Then, $\Phi(\Lambda,Y)=\Lambda$.
Proof.
By Lemma 2.4 and the proof of the previous lemma, it is enough to prove that $\Phi(\Lambda,Y(\Lambda))=\omega_{\nu,t}^{*}(\lambda)$ where we recall that $\Phi(\Lambda,Y)$ is defined as the smallest solution $\Lambda_{1}$ in (2.12). We denote in short $\alpha=\omega_{\nu,t}(\lambda)$. Since $y_{0}(\Lambda)\neq 0$, we may assume $\Lambda_{1}\neq\Lambda$. We claim that also $\Lambda_{1}\neq\alpha$ because
$\sum_{k=0}^{p}\frac{y_{k}}{\eta_{k}-\alpha}\neq 0.$ (2.43)
Indeed, using that for every real numbers $a\neq b\neq x$,
$\frac{1}{(x-a)^{2}(x-b)}=-\frac{1}{(a-b)^{2}}\left(\frac{1}{x-a}+\frac{b-a}{(x-a)^{2}}-\frac{1}{x-b}\right),$ (2.44)
and setting $a=\alpha$ and $b=\Lambda$, one gets
$\begin{split}\sum_{k=0}^{p}\frac{y_{k}}{\eta_{k}-\alpha}&=\frac{y_{0}}{\Lambda-\alpha}-\frac{t}{(\Lambda-\alpha)^{2}}(-G_{\mu}(\alpha)+G_{\mu}(\Lambda)-(\Lambda-\alpha)G_{\mu}^{\prime}(\alpha))+O(\varepsilon_{n})\\ &=-\frac{1}{(\Lambda-\alpha)^{2}}(H_{\nu,t}(\alpha)-H_{\nu,t}(\Lambda)-tG_{\mu}(\alpha)+tG_{\mu}(\Lambda)-t(\Lambda-\alpha)G_{\mu}^{\prime}(\alpha))+O(\varepsilon_{n})\\ &=\frac{1+tG_{\nu}^{\prime}(\alpha)}{\Lambda-\alpha}.\end{split}$
Since $\lambda<\ell_{\nu,t}$, we have $\alpha=\omega_{\nu,t}(\lambda)<\omega_{\nu,t}(\ell_{\nu,t})$ and therefore $01+tG_{\nu}^{\prime}(\alpha)>$ 0, which yields
$\sum_{k=0}^{p}\frac{y_{k}}{\eta_{k}-\alpha}\neq 0,$
which proves (2.43). We can thus take $\Lambda_{1}\notin\{\Lambda,\alpha\}$. Then, by partial fraction decomposition,
$\sum_{k=1}^{p}\frac{\alpha_{k}}{(\eta_{k}-\Lambda_{1})(\eta_{k}-\Lambda)(\eta_{k}-\alpha)}\\ =-\frac{1}{(\Lambda_{1}-\Lambda)(\Lambda_{1}-\alpha)}G_{\nu}(\Lambda_{1})-\frac{1}{(\Lambda-\alpha)(\Lambda-\Lambda_{1})}G_{\nu}(\Lambda)-\frac{1}{(\Lambda_{1}-\alpha)(\Lambda-\alpha)}G_{\nu}(\alpha),$
which implies
$\frac{y_{0}}{\Lambda-\Lambda_{1}}-t\frac{1}{(\Lambda-\Lambda_{1})(\alpha-\Lambda_{1})}G_{\nu}(\Lambda_{1})+t\frac{1}{(\alpha-\Lambda)(\Lambda-\Lambda_{1})}G_{\nu}(\Lambda)-t\frac{1}{(\alpha-\Lambda_{1})(\alpha-\Lambda)}G_{\nu}(\alpha)=0.$
Multiplying by $(\Lambda-\Lambda_{1})(\alpha-\Lambda_{1})(\alpha-\Lambda)$ gives
$y_{0}(\alpha-\Lambda_{1})(\alpha-\Lambda)-t(\alpha-\Lambda)G_{\nu}(\Lambda_{1})+t(\alpha-\Lambda_{1})G_{\nu}(\Lambda)-t(\Lambda-\Lambda_{1})G_{\nu}(\alpha)=0.$
Recalling that $y_{0}(\alpha-\Lambda)=\lambda-H_{\nu,t}(\Lambda)$, this can be simplified into
$(\lambda-H_{\nu,t}(\Lambda))(\alpha-\Lambda_{1})-t(\alpha-\Lambda)G_{\nu}(\Lambda_{1}^{n})+t(\alpha-\Lambda_{1})G_{\nu}(\Lambda)-t(\Lambda-\Lambda_{1})G_{\nu}(\alpha)=O(\varepsilon_{n})$
$\Longleftrightarrow$ $(\alpha-\Lambda)(\alpha-\Lambda_{1})+tG_{\nu}(\alpha)(\alpha-\Lambda)-tG_{\nu}(\Lambda_{1})(\alpha-\Lambda)=0$
$\Longleftrightarrow$ $(\alpha-\Lambda_{1})+tG_{\nu}(\alpha)-tG_{\nu}(\Lambda_{1})=0$
$\Longleftrightarrow$ $H_{\nu,t}(\Lambda_{1})=\lambda\,.$ (2.45)
Since $\Lambda_{1}\neq\alpha$, it follows that $\Lambda_{1}$ equals $\omega_{\nu,t}^{*}(\lambda)$.
We can now give the proof of Lemma 2.8. Let us prove the upper bound (2.29). Let $Y(\gamma)\in\Delta^{p+1}$ be given by (2.31) with $\gamma$ as in (2.27). By (2.32), we know that $K_{\nu,t}(\Lambda,\lambda,Y(\gamma))=I_{\nu,t}^{\Lambda}(\lambda).$ Using (2.25), and taking $Y=Y(\gamma)$ in the infimum yields
${F_{\nu,t}^{-}(\Lambda,\lambda)\leq K_{\nu,t}(\Lambda,\lambda,Y(\gamma))+F_{\nu,t}^{-}(\Phi(\Lambda,Y),\lambda^{-})1_{\omega_{\nu,t}(\lambda)\geq\Phi(\Lambda,Y(\gamma))}.}$
As shown in Lemma 2.10, $\Phi(\Lambda,Y(\gamma))\geq\omega_{\nu,t}(\ell_{\nu,t})\geq\omega_{\nu,t}(\lambda)$ so that the last term in the right-hand side vanishes. It follows that
$F_{\nu,t}^{-}(\Lambda,\lambda)\leq K_{\nu,t}(\Lambda,\lambda,Y(\gamma))=I_{\nu,t}^{\Lambda}(\lambda)$
where we finally used (2.32). Using $F_{\nu,t}^{+}\leq F_{\nu,t}^{-}$ we get the desired upper bounds. Let us now assume that $\Lambda\geq\omega_{\nu,t}(\lambda)$ or $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda\leq\rho_{\nu,t}^{\Lambda}$. Since $F_{\nu,t}^{+}\geq 0$, one has
$F_{\nu,t}^{+}(\Lambda,\lambda)\geq\inf K_{\nu,t}(\Lambda,\lambda,\dot{)}.$
By Lemma 2.9, in the range of parameters that we consider, $K_{\nu,t}(\Lambda,\lambda,.)$ and $J_{\nu,t}^{\Lambda}$ have a unique minimizer $Y(\gamma)$ with $\gamma=\gamma(\Lambda,\lambda)$. Hence
$F_{\nu,t}^{+}(\Lambda,\lambda)\geq K_{\nu,t}(\Lambda,\lambda,Y(\gamma))=I_{\nu,t}^{\Lambda}(\lambda),$
which gives $F_{\nu,t}^{+}(\Lambda,\lambda)=I_{\nu,t}^{\Lambda}(\lambda)$ using (2.29). Since $F_{\nu,t}^{-}(\Lambda,\lambda)\geq F_{\nu,t}^{+}(\Lambda,\lambda)$ we get that $F_{\nu,t}^{-}(\Lambda,\lambda)=I_{\nu,t}^{\Lambda}(\lambda)$.
2.7. Solving the functional inequalities
In the next lemma we prove that the equality (2.30) in fact holds for any $\Lambda\leq\ell_{\nu}$ and $\lambda\leq\ell_{\nu,t}$, hence completing the proof of the weak large deviation principle (2.3).
Lemma 2.11.
Let $\Lambda\leq\ell_{\nu}$ and $\lambda<\ell_{\nu,t}$. Let $I_{\nu,t}^{\Lambda}(\lambda)$ be as in (2.28). We have
$F_{\nu,t}^{-}(\Lambda,\lambda)=F_{\nu,t}^{+}(\Lambda,\lambda)=I_{\nu,t}^{\Lambda}(\lambda).$
To prove Lemma 2.11, we first demonstrate that $I_{\nu,t}^{\Lambda}$ satisfies the functional equation necessary for it to be equal to $F_{\nu,t}^{\pm}$, as suggested by (2.19) and (2.25).
Lemma 2.12.
Let $\Lambda\leq\ell_{\nu}$. Let $\Phi(\Lambda,Y)$ be as in (2.12). Let $I_{\nu,t}^{\Lambda}$ be as in (2.28). Let $\lambda<\ell_{\nu,t}$. Let $\omega_{\nu,t}^{*}(\lambda)$ be the unique solution $\gamma$ in $[\omega_{\nu,t}(\ell_{\nu,t}),\ell_{\nu})$ of the equation $H_{\nu,t}(\gamma)=\lambda$. Define
$\mathcal{B}(\Lambda,\lambda):=\inf_{Y\in\Delta^{p+1}}\Bigr{\{}K_{\nu,t}(\Lambda,\lambda,Y)+I_{\nu,t}^{\Phi(\Lambda,Y)}(\lambda)1_{\omega_{\nu,t}(\lambda)\geq\Phi(\Lambda,Y)}\Bigr{\}}.$
Then, the function $K_{\nu,t}(\Lambda,\lambda,\cdot)+I_{\nu,t}^{\Phi(\Lambda,\cdot)}(\lambda)1_{\omega_{\nu,t}(\lambda)\geq\Phi(\Lambda,\cdot)}$ admits a unique minimizer $Y^{*}\in\Delta^{p+1}$. Furthermore,
$\Phi(\Lambda,Y^{*})\geq\omega_{\nu,t}(\ell_{\nu,t})$
and
$\mathcal{B}(\Lambda,\lambda)=I_{\nu,t}^{\Lambda}(\lambda).$
Moreover,
  • If $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$, or $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda\leq\rho_{\nu,t}^{\Lambda}$, then $Y^{*}=Y(\Lambda)$ with $Y(\Lambda)$ as defined in (2.31),
  • If $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda\in(\rho_{\nu,t}^{\Lambda},\ell_{\nu,t}]$, then $y_{0}^{*}$ vanishes. Moreover, $Y^{*}=Y(\omega^{*}_{\nu,t}(\lambda))$ with $Y(\omega^{*}_{\nu,t}(\lambda))$ as defined in (2.31).
Proof.
Denote in short
$G:Y\in\Delta^{p+1}\mapsto J_{\nu,t}^{\Lambda}(\lambda,Y)-\int\log|\lambda-x|\mathrm{d}(\nu\boxplus\sigma_{t})(x)+\frac{\lambda^{2}}{4t}+I_{\nu,t}^{\Phi(\Lambda,Y)}(\lambda)1_{\Phi(\Lambda,Y)<\omega_{\nu,t}(\lambda)}-C_{t}.$
Since $I_{\nu,t}^{\Phi(\Lambda,Y)}\geq 0$, we deduce from the first step of the proof of Lemma 2.9 that $G$ is bounded from below (by $K_{\nu,t}$). We claim that it is lower semicontinuous in $Y\in\Delta^{p+1}$ when $\Delta^{p+1}$ is equipped with the weak topology. We already checked in the previous proof that $J_{\nu,t}^{\Lambda}(\lambda,.)$ is lower semicontinuous . Moreover $Y\mapsto\Phi(\Lambda,Y)$ is continuous by Lemma 2.4. Furthermore, when $\Phi(\Lambda,Y)\leq\omega_{\nu,t}(\lambda)$, we have
$I_{\nu,t}^{\Phi(\Lambda,Y)}(\lambda)=\frac{1}{2}(S_{\nu}(\omega_{\nu,t}(\lambda))-S_{\nu}(\Phi(\Lambda,Y)))+\frac{1}{4t}((\lambda-\Phi(\Lambda,Y))^{2}-(\lambda-\omega_{\nu,t}(\lambda))^{2}){.}$
Since $S_{\nu}$ is continuous on $(-\infty,\ell_{\nu})$ and ${\omega_{\nu,t}(\lambda)\leq\ell_{\nu}}$, the lower semi-continuity of
$Y\mapsto I_{\nu,t}^{\Phi(\Lambda,Y)}(\lambda)$
follows. The lower semi-continuity of $Y\mapsto I_{\nu,t}^{\Phi(\Lambda,Y)}(\lambda)1_{\Phi(\Lambda,Y)<\omega_{\nu,t}(\lambda)}$ then follows from the fact that $I_{\nu,t}^{\Phi(\Lambda,Y)}(\lambda)$ vanishes when $\Phi(\Lambda,Y)=\omega_{\nu,t}(\lambda)$. It follows that $G$ is bounded from below and lower semicontinuous , and hence achieves its minimal value. Next, we characterize the minimizers of $G$.
Step 1: study of the local minima such that $\Phi(\Lambda,\cdot)\geq\omega_{\nu,t}(\lambda)$.
On the set $\{Y\in\Delta^{p+1}:\Phi(\Lambda,\cdot)\geq\omega_{\nu,t}(\lambda)\}$, the last term in $G$ vanishes and therefore the minimizers of $G$ on this set are the minimizers $Y$ of $J_{\nu,t}^{\Lambda}(\lambda,\cdot)$ that satisfy $\Phi(\Lambda,Y)\geq\omega_{\nu,t}(\lambda)$.
The minimizers of $J_{\nu,t}^{\Lambda}(\lambda,\cdot)$ are in $\{Y(\Lambda),Y(\omega_{\nu,t}^{*}(\lambda))\}$ by Lemma 2.9. In view of Lemma 2.5, $\Phi(\Lambda,Y(\Lambda))=\omega_{\nu,t}^{*}(\lambda)\geq\omega_{\nu,t}(\lambda)$. Besides $\Phi(\Lambda,Y(\omega_{\nu,t}^{*}(\lambda)))=\Lambda$. Therefore the possible minimizers of $G$ on the set $\{Y\in\Delta^{p+1}:\Phi(\Lambda,\cdot)\geq\omega_{\nu,t}(\lambda)\}$ are $Y(\Lambda)$, and $Y(\omega_{\nu,t}^{*}(\lambda))$ whenever $\Lambda\geq\omega_{\nu,t}(\lambda)$. Additionally, (2.32) shows that the value of $G(Y({\gamma}))$ for $\gamma\in\{\omega_{\nu,t}^{*}(\lambda),\Lambda\}$ is $I_{\nu,t}^{\Lambda}(\lambda,\gamma)$.
Step 2: study of the critical points such that $\Phi(\Lambda,\cdot)<\omega_{\nu,t}(\lambda)$ and $y_{0}>0$
Let $Y\in\Delta^{p+1}$ be a critical point of $G$ such that $y_{0}>0$ and
$\Phi(\Lambda,Y)<\omega_{\nu,t}(\lambda).$ (2.46)
By variational calculus, there exist constants $R_{1},R_{2}$ and $C_{0}$ such that for ${k\in\{1,\ldots,p\}}$,
$-\lambda\eta_{k}+\eta_{k}^{2}-R_{1}\eta_{k}-\frac{t}{y_{k}}-C_{0}\frac{1}{\eta_{k}-\Lambda_{1}}=R_{2}$
and
$-\lambda\Lambda+\Lambda^{2}-R_{1}\Lambda-\frac{C_{0}}{\Lambda-\Lambda_{1}}=R_{2},$ (2.47)
with $R_{1}=\Lambda y_{0}+\int\eta_{\nu}\mathrm{d}\mu$, ${C_{0}=\partial_{\Lambda}I_{\nu,t}^{\Phi(\Lambda,Y)}(.,\lambda)|_{\Phi(\Lambda,Y)}}$ and $\Lambda_{1}=\Phi(\Lambda,Y)$. It follows that for each $k=1,\ldots,p$,
$y_{k}=\frac{t}{\eta_{k}^{2}-(\lambda+R_{1})\eta_{\nu}-R_{2}-\frac{C_{0}}{\eta_{k}-\Lambda_{1}}}=\frac{t(\eta_{k}-\Lambda_{1})}{(\eta_{k}^{2}-(\lambda+R_{1})\eta_{k}-R_{2})(\eta_{\nu}-\Lambda_{1})-C_{0}}.$
Let us write the denominator in the last display as
$P(\eta_{k}):=(\eta_{k}^{2}-(\lambda+R_{1})\eta_{k}-R_{2})(\eta_{k}-\Lambda_{1})-C_{0}=(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})(\eta_{k}-\gamma_{3}),$
with $\gamma_{3}\in\mathbb{R}$, $\gamma_{2},\gamma_{1}\in\mathbb{C}$. Identifying the coefficient in front of $\eta_{\nu}^{2}$, we observe that
$\Lambda+\gamma_{1}+\gamma_{2}=\lambda+R_{1}+\Lambda_{1}.$ (2.48)
In view of (2.47), $\Lambda$ is one of the roots of $P$, say $\gamma_{3}=\Lambda$.
$\bullet$ Assume that $\Lambda\neq\gamma_{1}\neq\gamma_{2}$. By definition of $\Phi(\Lambda,Y)=\Lambda_{1}$, see (2.12), we have
$\frac{y_{0}}{\Lambda-\Lambda_{1}}+\sum_{k=1}^{p}\frac{t}{(\eta_{k}-\Lambda)(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}=0.$ (2.49)
Using the partial fraction decomposition
$\frac{1}{(\eta_{k}-\Lambda)(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}$ $=$ $\frac{1}{(\Lambda-\gamma_{1})(\Lambda-\gamma_{2})}\frac{1}{\eta_{k}-\Lambda}+\frac{1}{(\gamma_{1}-\Lambda)(\gamma_{1}-\gamma_{2})}\frac{1}{\eta_{k}-\gamma_{1}}$
$+\frac{1}{(\gamma_{2}-\Lambda)(\gamma_{2}-\gamma_{1})}\frac{1}{\eta_{k}-\gamma_{2}},$
one may rewrite (2.49) as
$\frac{y_{0}}{\Lambda-\Lambda_{1}}-\frac{1}{(\Lambda-\gamma_{1})(\Lambda-\gamma_{2})}G_{\nu}(\Lambda)+\frac{1}{(\Lambda-\gamma_{1})(\gamma_{1}-\gamma_{2})}G_{\nu}(\gamma_{1})-\frac{1}{(\Lambda-\gamma_{2})(\gamma_{1}-\gamma_{2})}G_{\nu}(\gamma_{2})=0.$ (2.50)
The condition $\sum_{k=0}^{p}y_{k}=1$ reads
$y_{0}+t\sum_{k=1}^{p}\frac{\eta_{k}-\Lambda_{1}}{(\eta_{k}-\Lambda)(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}\mathrm{d}x=1,$
which gives
$t\sum_{k=1}^{p}\frac{1}{(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}+t(\Lambda-\Lambda_{1})\sum_{k=1}^{p}\frac{1}{(\eta_{k}-\Lambda)(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}=1-y_{0}.$
Hence, using (2.49), one gets
$\sum_{k=1}^{p}\frac{t}{(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}=1.$ (2.51)
Using
$\frac{1}{(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}=\frac{1}{\gamma_{1}-\gamma_{2}}\frac{1}{\eta_{k}-\gamma_{1}}-\frac{1}{\gamma_{1}-\gamma_{2}}\frac{1}{\eta_{k}-\gamma_{2}},$
we deduce that
$\frac{1}{\gamma_{1}-\gamma_{2}}(-tG_{\nu}(\gamma_{1})+tG_{\nu}(\gamma_{2}))=1,$
which gives
$H_{\nu,t}(\gamma_{1})=H_{\nu,t}(\gamma_{2}).$ (2.52)
In view of (2.48), we have
$\Lambda y_{0}+t\sum_{k=1}^{p}\frac{(\eta_{k}-\Lambda_{1})\eta_{k}}{(\eta_{k}-\Lambda)(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}=R_{1}=\Lambda+\gamma_{1}+\gamma_{2}-\lambda-\Lambda_{1},$
which implies
$\Lambda y_{0}+t\sum_{k=1}^{p}\frac{\eta_{k}-\Lambda_{1}}{(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}+\Lambda(1-y_{0})=\Lambda+\gamma_{1}+\gamma_{2}-\lambda-\Lambda_{1}.$
Simplifying further we obtain
$-tG_{\nu}(\gamma_{2})+(\gamma_{1}-\Lambda_{1})=\gamma_{1}+\gamma_{2}-\lambda-\Lambda_{1}.$
Therefore, inserting (2.52), we conclude that
$H_{\nu,t}(\gamma_{2})=H_{\nu,t}(\gamma_{1})=\lambda.$
Therefore $\gamma_{1}=\omega_{\nu,t}(\lambda)$ and $\gamma_{2}=\omega^{*}_{\nu,t}(\lambda)$. Moreover we can compute using (2.50) that
$\frac{y_{0}}{\Lambda-\Lambda_{1}}(\Lambda-\gamma_{1})(\Lambda-\gamma_{2})=\Lambda+tG_{\nu}(\Lambda)-\lambda$
which implies
$\Lambda_{1}=\Lambda-y_{0}\frac{(\Lambda-\gamma_{1})(\Lambda-\gamma_{2})}{H_{\nu,t}(\Lambda)-\lambda}\,.$ (2.53)
$\bullet$ Assume that $\Lambda=\gamma_{1}=\gamma_{2}$. Then proceeding as in (2.51), we get
$\sum_{k=1}^{p}\frac{t}{(\eta_{k}-\Lambda)^{2}}=1,$
implying that $\Lambda=\omega_{\nu,t}(\ell_{\nu,t})$. In particular, $\Phi(\Lambda,Y)\geq\omega_{\nu,t}(\lambda)$, which contradicts (2.46).
$\bullet$ Assume that $\Lambda=\gamma_{1}$ and $\gamma_{1}\neq\gamma_{2}$. Proceeding as above we easily get $H_{\nu,t}(\Lambda)=H_{\nu,t}(\gamma_{2})=\lambda$, from which it is inferred that $\Lambda=\omega_{\nu,t}(\lambda)$, implying that $\Phi(\Lambda,Y)\geq\omega_{\nu,t}(\lambda)$, thus contradicting (2.46).
$\bullet$ Assume that $\Lambda\neq\gamma_{1}$ and $\gamma_{1}=\gamma_{2}=\gamma$. Then we get
$\sum_{k=1}^{p}\frac{t}{(\eta_{k}-\gamma)^{2}}=1$
and $H_{\nu,t}(\gamma)=\lambda$, which implies that $\lambda=\ell_{\nu,t}$ and $\gamma=\omega_{\nu,t}(\ell_{\nu,t})$, contradicting the assumption that $\lambda<\ell_{\nu,t}$.
We conclude that if $Y$ is a critical point of $G$ satisfying (2.46), then it is of the form
$y_{k}=\frac{(\eta_{k}-\Lambda_{1})}{(\eta_{k}-\Lambda)(\eta_{k}-\omega_{\nu,t}(\lambda))(\eta_{k}-\omega_{\nu,t}^{*}(\lambda))},\quad\text{for each $1\leq k\leq p$},$ (2.54)
for some $\Lambda_{1}\in(\Lambda,\ell_{\nu})$ such that letting $y_{0}:=1-\sum_{k=1}^{p}y_{k}$, the relations (2.53), (2.49) hold. In the sequel we denote $\gamma_{1}:=\omega_{\nu,t}(\lambda)$ and $\gamma_{2}:=\omega_{\nu,t}^{*}(\lambda)$. We next show that this critical point will never be a minimizer of $G$.
Step 3: study of the minimizers.
A minimizer $Y$ of $G$ is either such that $\Phi(\Lambda,Y)<\omega_{\nu,t}(\lambda)$ or a critical point of ${J_{\nu,t}^{\Lambda}(\lambda,\cdot)}$ such that $\Phi(\Lambda,Y)\geq\omega_{\nu,t}(\lambda)$. Therefore it is either of the form (2.54) or of the form $Y(\gamma)$, for $\gamma\in\{\Lambda,\omega^{*}_{\nu,t}(\lambda)\}$ by Lemma 2.9.
$\bullet$ Assume that $\Lambda\geq\omega_{\nu,t}(\lambda)$. Then, since $\Phi(\Lambda,Y)\geq\Lambda\geq\omega_{\nu,t}(\lambda)$,
$G(Y)=J_{\nu,t}^{\Lambda}(\lambda,Y)-\int\log|\lambda-x|\mathrm{d}(\nu\boxplus\sigma_{t})(x)+\frac{\lambda^{2}}{4t}-C_{t}{.}$
The function $G(Y)$ is therefore uniquely minimized by the unique minimizer of $J_{\nu,t}^{\Lambda}(\lambda,\cdot)$ described in Lemma 2.9.
$\bullet$ Assume that $\Lambda<\omega_{\nu,t}(\lambda)$ and $\lambda=\ell_{\nu,t}$. By Lemma 2.9, the minimum of $J_{\nu,t}^{\Lambda}(\lambda,\cdot)$ is uniquely attained at $Y=Y(\Lambda)$. Moreover $\Phi(\Lambda,Y(\Lambda))=\omega^{*}_{\nu,t}(\ell_{\nu,t})=\omega_{\nu,t}(\ell_{\nu,t})$. Therefore $G(Y(\Lambda))=I_{\nu,t}^{\Lambda}(\ell_{\nu,t})$. Hence using $I_{\nu,t}^{\Lambda}\geq 0$, we see that if $Y$ is a minimizer of $G$, then $J^{\Lambda}_{\nu,t}(\ell_{\nu,t},Y)=J^{\Lambda}_{\nu,t}(\ell_{\nu,t},Y(\Lambda))$. Therefore $Y=Y(\Lambda)$.
$\bullet$ Assume that $\Lambda<\omega_{\nu,t}(\lambda)$ and $\lambda<\ell_{\nu,t}$. Then, for every $Y\in\Delta^{p}$, $\Phi(\Lambda,0,Y)=\Lambda$ and
$G(0,Y)=J_{\nu,t}^{\Lambda}(\lambda,0,Y)-\int\log|\lambda-x|\mathrm{d}(\nu\boxplus\sigma_{t})(x)+\frac{\lambda^{2}}{4t}+I_{\nu,t}^{\Lambda}(\lambda)1_{\Lambda<\omega_{\nu,t}(\lambda)}-C_{t}{.}$
Since the term depending on $I_{\nu,t}^{\Lambda}$ does not depend on $Y$ we conclude that $G(0,\cdot)$ is minimal at the minimum of $J_{\nu,t}^{\Lambda}(0,\cdot)$, namely at $(y_{1}(\omega_{\nu,t}^{*}(\lambda)),\ldots,y_{p}(\omega_{\nu,t}^{*}(\lambda)))$ and therefore
$\inf_{Y\in\Delta^{p}}G(0,Y)=I_{\nu,t}^{\omega_{\nu,t}^{*}(\lambda)}(\lambda)+I_{\nu,t}^{\Lambda}(\lambda).$
One can check that that for all $\lambda^{\prime}<\ell_{\nu,t}$,
$\frac{\partial}{\partial\lambda^{\prime}}I_{\nu,t}^{\omega_{\nu,t}^{*}(\lambda^{\prime})}(\lambda)=\frac{1}{2t}(\omega_{\nu,t}(\lambda^{\prime})-\omega_{\nu,t}^{*}(\lambda^{\prime}))<0.$
Since $I_{\nu,t}^{\omega_{\nu,t}^{*}(\ell_{\nu,t})}(\ell_{\nu,t})=0$ we deduce that for all $\lambda^{\prime}<\ell_{\nu,t}$,
$I_{\nu,t}^{\omega_{\nu,t}^{*}(\lambda^{\prime})}(\lambda^{\prime})>0$ (2.55)
and therefore
$\inf_{Y\in\Delta^{p}}G(0,Y)>I_{\nu,t}^{\Lambda}(\lambda)=G(Y(\Lambda))$
with $Y(\Lambda)\in\Delta^{p+1}$. It follows that minimizers of $G$ are not of the form $(0,Y)$ with $Y\in\Delta^{p+1}$.
Let us prove that $Y(\Lambda)$ uniquely minimizes $G$. Since $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$, $Y(\Lambda)\in\Delta^{p+1}$. By (2.32), we have $G(Y({\Lambda}))=I_{\nu,t}^{\Lambda}(\lambda).$ Therefore the minimizer of $G$ is either $Y(\Lambda)$ or $(y_{0},Y)$ of the form (2.54).
Let $(y_{0},Y)$ be of the form (2.54). Let us compute $G(y_{0},Y)$. Define $X:=\Lambda y_{0}+\sum_{k=1}^{p}\eta_{k}y_{k}$. Recall that $X=\Lambda+\gamma_{1}+\gamma_{2}-\lambda-\Lambda_{1}$. We have
$\sum_{k=1}^{p}\eta_{k}^{2}y_{k}=t-(\Lambda_{1}-\Lambda)tG_{\nu}(\Lambda)+(\gamma_{1}+\gamma_{2})X-\Lambda(\gamma_{1}+\gamma_{2})y_{0}-\gamma_{1}\gamma_{2}(1-y_{0}).$
Therefore
$\Lambda^{2}y_{0}+\sum_{k=1}^{p}\eta_{k}^{2}y_{k}=(\Lambda-\gamma_{1})(\Lambda-\gamma_{2})y_{0}-\gamma_{1}\gamma_{2}-(\Lambda_{1}-\Lambda)tG_{\nu}(\Lambda)+(\gamma_{1}+\gamma_{2})X.$
Inserting (2.53), we can therefore rewrite the above display as
$\Lambda^{2}y_{0}+\sum_{k=1}^{p}\eta_{k}^{2}y_{k}=-\gamma_{1}\gamma_{2}+(\gamma_{1}+\gamma_{2})X+(\Lambda-\Lambda_{1})(\Lambda-\lambda).$
It follows that
$\begin{split}G(y_{0},Y)&=\frac{1}{2}-\frac{1}{2}\ln t+\frac{1}{4t}(-2\lambda-X+2(\gamma_{1}+\gamma_{2}))X-\frac{\gamma_{1}\gamma_{2}}{2t}+\frac{1}{2t}(\Lambda-\Lambda_{1})(\Lambda-\lambda)\\ &-\frac{1}{2}(S_{\nu}(\Lambda)+S_{\nu}(\gamma_{1})+S_{\nu}(\gamma_{2})-S_{\nu}(\Lambda_{1}))+\frac{1}{2}(S_{\nu}(\gamma_{1})-S_{\nu}(\Lambda_{1}))+\frac{1}{4t}(\Lambda_{1}^{2}-2\lambda\Lambda_{1}-\gamma_{1}^{2}+2\lambda\gamma_{1})\\ &=-\frac{1}{2}(S_{\nu}(\gamma_{2})+S_{\nu}(\Lambda))+\frac{Y}{4t}+\frac{1}{2}-\frac{1}{2}\ln t,\end{split}$
with
$\begin{split}Y&:=(-2\lambda-X+2(\gamma_{1}+\gamma_{2}))X-2\gamma_{1}\gamma_{2}+2(\Lambda^{2}+\Lambda_{1}\lambda-\Lambda\lambda-\Lambda\lambda)+\Lambda_{1}^{2}-2\lambda\Lambda_{1}-\gamma_{1}^{2}+2\lambda\gamma_{1}\\ &=-\Lambda^{2}+\lambda^{2}+\gamma_{1}^{2}+\gamma_{2}^{2}-2\lambda\gamma_{1}-2\lambda\gamma_{2}+2\gamma_{1}\gamma_{2}+2\Lambda\Lambda_{1}-\Lambda_{1}^{2}-2\gamma_{1}\gamma_{2}\\ &+2(\Lambda^{2}+\Lambda_{1}\lambda-\Lambda\lambda-\Lambda\lambda)+\Lambda_{1}^{2}-2\lambda\Lambda_{1}-\gamma_{1}^{2}+2\lambda\gamma_{1}\\ &=\Lambda^{2}+\lambda^{2}+\gamma_{2}^{2}-2\gamma_{2}\lambda-2\lambda\Lambda.\end{split}$
We therefore find that
$G(y_{0},Y)=\frac{1}{4t}(\Lambda^{2}+\lambda^{2}+\gamma_{2}^{2}-2\gamma_{2}\lambda-2\lambda\Lambda)-\frac{1}{2}(S_{\nu}(\gamma_{2})+S_{\nu}(\Lambda))+\frac{1}{2}-\frac{1}{2}\ln t.$ (2.56)
We next show that this is greater than $G(Y({\Lambda}))$. Let us define
$\phi:\gamma\in(-\infty,\ell_{\nu})\mapsto\frac{1}{4t}(\Lambda^{2}+\lambda^{2}+\gamma^{2}-2\gamma\lambda-2\lambda\Lambda)-\frac{1}{2}(S_{\nu}(\gamma)+S_{\nu}(\Lambda))+\frac{1}{2}-\frac{1}{2}\ln t.$ (2.57)
For all $\gamma<\ell_{\nu}$, one has
$\phi^{\prime}(\gamma)=\frac{H_{\nu,t}(\gamma)-\lambda}{2t}.$
Therefore $\phi$ is strictly increasing on $[\gamma_{1},\gamma_{2}]$. It follows that if $\gamma_{1}<\gamma_{2}$,
$\phi(\gamma_{2})=G(y_{0},Y)>G(Y(\Lambda))=\phi(\gamma_{1})$
and $Y(\Lambda)$ is the unique minimizer of $G$. If $\gamma_{1}=\gamma_{2}$, then $Y(\Lambda)$ is still the unique minimizer of $G$.
Step 4: conclusion
The unique minimizer of $G$ is the same as that of $J_{\nu,t}^{\Lambda}$, except in the previously undecided case $\Lambda<\omega_{\nu,t}(\lambda)$, where it is now given by $Y({\Lambda})$. ∎
We finally prove Lemma 2.11 by iterating the functional inequality (2.19) and showing that after a finite number of steps it does not depend on $F^{+}_{\nu,t}$, so that $F^{+}_{\nu,t}$ is bounded from below by $I_{\nu,t}^{\Lambda}$.
Proof of Lemma 2.11.
By Lemma 2.8, we only need to consider the case where $\Lambda<\omega_{\nu,t}(\lambda)$ and $\lambda<\ell_{\nu,t}$. Let $\lambda<\ell_{\nu,t}$ and $\Lambda<\omega_{\nu,t}(\lambda)$. Recall
$F_{\nu,t}^{+}(\Lambda,\lambda)\geq\inf_{Y\in\Delta^{p+1}}\{K_{\nu,t}(\Lambda,\lambda,Y)+F_{\nu,t}^{+}(\Phi(\Lambda,Y),\lambda^{-})1_{\Phi(\Lambda,Y)<\omega_{\nu,t}(\lambda)}\}.$
Recall from Lemma B.2 that $\lambda\mapsto F_{\nu,t}^{+}(\Phi(\Lambda,Y),\lambda^{-})$ increases on $\{\lambda<\ell_{\nu,t}:\Phi(\Lambda,Y)<\omega_{\nu,t}(\lambda)\}$. Hence, for $\varepsilon>0$ small enough,
$F_{\nu,t}^{+}(\Lambda,\lambda)\geq\inf_{Y\in\Delta^{p+1}}\{K_{\nu,t}(\lambda,\Lambda,Y)+F_{\nu,t}^{+}(\Phi(\Lambda,Y),\lambda-\varepsilon)1_{\Phi(\Lambda,Y)<\omega_{\nu,t}(\lambda)}\},$ (2.58)
By iteration define $\Phi_{0}(\Lambda,Y)=\Lambda,\Phi_{1}(\Lambda,Y)=\Phi(\Lambda,Y)$ and for each $k\geq 1$,
$\Phi_{k}(\Lambda,Y_{1},\ldots,Y_{k}):=\Phi(\Phi_{k-1}(\Lambda,Y_{1},\ldots,Y_{k-1}),Y_{k})\quad{\text{for all $Y_{1},\ldots,Y_{k}\in\Delta^{p+1}$.}}$
Recall that for every $\Lambda,Y$, $\Phi(\Lambda,Y)\geq\Lambda$. Consequently for each $k\geq 1$ and all $Y_{1},\ldots,Y_{k}\in\Delta^{p+1}$,
$\bigcap_{i=1}^{k}\{\Phi_{i}(\Lambda,Y_{1},\ldots,Y_{i})<\omega_{\nu,t}(\lambda)\}=\{\Phi_{k}(\Lambda,Y_{1},\ldots,Y_{k})<\omega_{\nu,t}(\lambda)\}.$ (2.59)
Also recall that by Lemma 2.9, $\inf K_{\nu,t}(\Phi_{k-1}(\Lambda,Y_{1},\ldots,Y_{k-1}),\lambda,\cdot)=0$ when $\Phi_{k-1}(\Lambda,Y_{1},\ldots,Y_{k-1})=\omega_{\nu,t}(\lambda)$. Let $n\geq 0$. Applying (2.58) $n+1$ times for $\varepsilon$ small enough, gives
$F_{\nu,t}^{+}(\Lambda,\lambda)\geq\inf_{Y_{1},\ldots,Y_{n}}\Bigr{(}\sum_{k=1}^{n}K_{\nu,t}(\Phi_{k-1}(\Lambda,Y_{1},\ldots,Y_{k-1}),\lambda-k\varepsilon,Y_{k})1_{\cap_{\ell\leq k}\{\Phi_{\ell-1}(\Lambda,Y_{1},\ldots,Y_{\ell-1})<\omega_{\nu,t}(\lambda-(\ell-1)\varepsilon)\}}\\ +F_{\nu,t}^{+}(\Phi_{n}(\Lambda,Y_{1},\ldots,Y_{n}),\lambda-(n+1)\varepsilon)1_{\cap_{\ell\leq n}\{\Phi_{\ell}(\Lambda,Y_{1},\ldots,Y_{\ell})<\omega_{\nu,t}(\lambda-(\ell-1)\varepsilon)\}}\Bigr{)}.$ (2.60)
Recall that $\Lambda{\mapsto}F_{\nu}^{+}(\Lambda,\lambda)$ is lower semicontinuous and $H$ is continuous so that the above minimum is achieved on the compact set $(\Delta^{p+1})^{n}$. Let $(Y_{1}^{\varepsilon},\ldots,Y_{n}^{\varepsilon})$ be a minimizer. By compactness, one can extract a subsequence $(Y_{1}^{\varepsilon_{l}},\ldots,Y_{n}^{\varepsilon_{l}})$ converging as $\varepsilon$ goes to zero to some $(Y_{1},\ldots,Y_{n})\in\Delta^{p+1}$.
$\bullet$ Assume that $\Phi_{n}(\Lambda,Y_{1},\ldots,Y_{n})>\omega_{\nu,t}(\lambda)$. Then, by continuity of $\Phi$ and $\omega_{\nu,t}$, for $\varepsilon$ small enough, $\Phi_{n}(\Lambda,Y_{1}^{\varepsilon},\ldots,Y_{n}^{\varepsilon})>\omega_{\nu,t}(\lambda-(n-1)\varepsilon)$. Hence, in this case, by letting $\varepsilon$ going to zero and using (2.59), we find
$F_{\nu,t}^{+}(\Lambda,\lambda)$$\geq$ $\sum_{k=1}^{n}K_{\nu,t}(\Phi_{k-1}(\Lambda,Y_{1},\ldots,Y_{k-1}),\lambda,Y_{k})1_{\{\Phi_{k-1}(\Lambda,Y_{1},\ldots,Y_{k-1})<\omega_{\nu,t}(\lambda)\}}$
$+I_{\nu,t}^{\Phi_{n}(\Lambda,Y_{1},\ldots,Y_{n})}(\lambda)1_{\{\Phi_{n}(\Lambda,Y_{1},\ldots,Y_{n})<\omega_{\nu,t}(\lambda)\}}$
$\geq$ $\inf_{Y_{1},\ldots,Y_{n}}\Bigr{(}\sum_{k=1}^{n}K_{\nu,t}(\Phi_{k-1}(\Lambda,Y_{1},\ldots,Y_{k-1}),\lambda,Y_{k})1_{\{\Phi_{k-1}(\Lambda,Y_{1},\ldots,Y_{k-1})<\omega_{\nu,t}(\lambda)\}}$
$+I_{\nu,t}^{\Phi_{n}(\Lambda,Y_{1},\ldots,Y_{n})}(\lambda)1_{\{\Phi_{n}(\Lambda,Y_{1},\ldots,Y_{n})<\omega_{\nu,t}(\lambda)\}}\Bigr{)},$
where in the first line we kept the last term and replaced $F^{+}_{\nu,t}$ by $I_{\nu,t}$ since it is multiplied by zero. We can then take the infimum over $Y_{n}$. By Lemma 2.12,
$\inf_{Y_{n}}\{K_{\nu,t}(\Phi_{n-1}(\Lambda,Y_{1},\ldots,Y_{n-1}),\lambda,Y_{n})+I_{\nu,t}^{\Phi_{n}(\Lambda,Y_{1},\ldots,Y_{n})}(\lambda)1_{\{\Phi_{n}(\Lambda,Y_{1},\ldots,Y_{n})<\omega_{\nu,t}(\lambda)\}}\}\qquad$
$\qquad=I_{\nu,t}^{\Phi_{n-1}(\Lambda,Y_{1},\ldots,Y_{n}-1)}(\lambda)1_{\{\Phi_{\ell}(\Lambda,Y_{1},\ldots,Y_{n-1})<\omega_{\nu,t}(\lambda)\}}$
and proceeding inductively we deduce that
$F_{\nu,t}^{+}(\Lambda,\lambda)\geq I_{\nu,t}^{\Lambda}(\lambda).$
$\bullet$ We show that it is impossible that $\Phi_{n}(\Lambda,Y_{1},\ldots,Y_{n})\leq\omega_{\nu,t}(\lambda)$ if $n$ is large enough. We first notice that by definition for every $\Lambda,Y$,
$\frac{y_{0}}{\Phi(\Lambda,Y)-\Lambda}=\sum_{i=1}^{p}\frac{y_{i}}{\eta_{i}-\Phi(\Lambda,Y)}{.}$ (2.61)
We can assume without loss of generality that $\Lambda$ is such that $\eta_{i}-\Phi(\Lambda,Y)>\gamma$ for some $\gamma>0$; otherwise if $\Phi(\Lambda,Y)$ is close to $\ell_{\nu}$, then $\Phi(\Lambda,Y)$ would be greater than $\omega_{\nu,t}(\lambda)$, placing us in the previously discussed scenario. Hence, by (2.61),
$\Phi(\Lambda,Y)\geq\Lambda+\gamma y_{0}\,.$
We therefore have two cases:
  • Either $y_{0}\geq\delta/\gamma$, and then
    $\Phi(\Lambda,Y)\geq\Lambda+\delta\,.$ (2.62)
  • Or $y_{0}\leq\delta/\gamma$ and denote by $\tilde{Y}$ the element of $\Delta^{p+1}$ such that $\tilde{y}_{0}=0$ and $\tilde{y}_{i}=y_{i}(1-y_{0})^{-1}$ for $i\in\{1,\ldots,p\}$. The continuity of $K_{\nu,t}(\Lambda,\lambda,.)$ implies that for $\delta$ small enough, uniformly on $Y,\Lambda,\lambda$,
    $K_{\nu,t}(\Lambda,\lambda,Y)=K_{\nu,t}(\Lambda,\lambda,\tilde{Y})+O(\delta)\geq\min_{Y\in\Delta^{p}}\{K_{\nu,t}(\Lambda,\lambda,(0,Y))\}+O(\delta).$
    By Lemma 2.9, the minimum of $K_{\nu,t}(\Lambda,\lambda,\cdot)$ over elements of the form $(0,Y)$ is taken at $Y=Y(\omega^{*}_{\nu,t}(\lambda))$ and therefore
    $\min_{Y\in\Delta^{p}}\{K_{\nu,t}(\Lambda,\lambda,(0,Y))\}\geq I_{\nu,t}^{\omega^{*}_{\nu,t}(\lambda))}(\lambda).$ (2.63)
    Notice that the right-hand side does not depend on $\Lambda$ and moreover by (2.55), $I_{\nu,t}^{\omega^{*}_{\nu,t}(\lambda))}(\lambda)>0$. We hereafter choose $\delta$ small enough such that $\kappa(\lambda):=\min_{Y\in\Delta^{p}}\{K_{\nu,t}(\Lambda,\lambda,(0,Y))\}+O(\delta)$ is positive and independent of $\Lambda$.
Returning to (2.60), and recalling that $(Y_{1}^{\varepsilon},\ldots,Y_{n}^{\varepsilon})$ is a minimizer, and that $F_{\nu,t}^{+}$ is lower semicontinuous in the variable $\Lambda$ while $K_{\nu,t}$ and $\omega_{\nu,t}$ are continuous, we obtain the following: if again $(Y_{1},\cdots,Y_{n})$ denotes a limiting point of the optimizing sequence $(Y_{1}^{\varepsilon},\ldots,Y_{n}^{\varepsilon})$,
$F_{\nu,t}^{+}(\Lambda,\lambda)$$\geq$ $\sum_{k=1}^{n}K_{\nu,t}(\Phi_{k-1}(\Lambda,Y_{1},\ldots,Y_{k-1}),\lambda,Y_{k})1_{\{\Phi_{k-1}(\Lambda,Y_{1},\ldots,Y_{k-1})<\omega_{\nu,t}(\lambda)\}}$
$+F_{\nu,t}^{+}(\Phi_{n}(\Lambda,Y_{1},\ldots,Y_{n}),\lambda^{-})1_{\{\Phi_{n}(\Lambda,Y_{1},\ldots,Y_{\ell})<\omega_{\nu,t}(\lambda)\}}.$
From the above discussion, we see from (2.62) that if there are $\ell$ indices such that $Y_{\ell}(0)\geq\delta/\gamma$ then $\Phi_{n}(\Lambda,Y_{1},\ldots,Y_{n}),\lambda)\geq\Lambda+\ell\delta$ which implies that $\Phi_{n}(\Lambda,Y_{1},\ldots,Y_{n})>\omega_{\nu,t}(\lambda)$ if $\ell$ is greater than some constant, which contradicts our assumption. Therefore, we conclude that there exists $\ell_{0}$ finite such that for $n-\ell_{0}$ indices $i$, $Y_{i}(0)\leq y_{0}/\gamma$. But then, since $K_{\nu,t}$ and $F_{\nu,t}^{+}$ are non-negative, (2.63) implies
$F_{\nu,t}^{+}(\Lambda,\lambda)\geq(n-\ell_{0})\kappa(\lambda),$
which is impossible for $n$ large enough since $F_{\nu,t}^{+}(\Lambda,\lambda)$ is finite. Therefore, we deduce that there exists $n$ large enough, such that for each $k\geq n$, $\Phi_{k}(\Lambda,Y_{1},\ldots,Y_{k})>\omega_{\nu,t}(\lambda)$. We can argue as in the first step to conclude that
$F_{\nu,t}^{+}(\Lambda,\lambda)\geq I_{\nu,t}^{\Lambda}(\lambda).$
Using the upper bound (2.29), this concludes the proof of Lemma 2.11. ∎
3. Large deviations in the general case
In this section, we suppose that the sequence of matrices $(B_{N})$ satisfies Assumption 1.1. The measure $\nu$ is now allowed to have a continuous support.
3.1. Properties of the rate function
We can now prove the approximation and convexity properties. Recall that $\eta_{\nu}$ is the transport map such that $\nu=\eta_{\nu}\#1_{[0,1]}\mathrm{d}x$.
Lemma 3.1 (Properties of the rate function).
  1. (1)
    (Continuity) Let $\nu$ be a probability measure on $\mathbb{R}$ with bounded support. Let $(\nu_{n})_{n\geq 1}$ be a sequence of probability measures on $\mathbb{R}$ such that $(\eta_{\nu_{n}})$ converges uniformly towards $\eta_{\nu}$. Then for every $x\in(-\infty,{\ell_{\nu,t}})$,
    $\lim_{n\to\infty}I_{\nu_{n},t}^{\Lambda}(x)=I_{\nu,t}^{\Lambda}(x).$
  2. (2)
    (Convexity) Let $\nu$ be a probability measure on $\mathbb{R}$ with support bounded from below. The map $I_{\nu,t}^{\Lambda}$ is strictly convex and achieves its minimum uniquely, at $\rho_{\nu,t}^{\Lambda}:=\Lambda+tG_{\mu}(\Lambda)$ if $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$ or $\ell_{\nu,t}$ if $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$.
Proof.
Step 1: proof of continuity
It is easy to see that if $\eta_{\nu},\lambda$ and $\Lambda$ are uniformly bounded by $M$ and $\eta_{\nu_{n}}-\eta_{\nu}$ is uniformly bounded by $\varepsilon$, there exists a finite constant $C$ which only depends on $M$ and $t$ such that $\ell_{\nu,t}-\ell_{\nu_{n},t}$ are $\varepsilon$ close, $G_{\nu}(x)-G_{\nu_{n}}(x)$ and $H_{\nu,t}(x)-H_{\nu_{n},t}(x)$ is of order $\kappa^{-1}\varepsilon$ if $x<\ell_{\nu,t}-\kappa$, so that since $H^{\prime}_{\nu,t}(x)$ is bounded below by a positive constant, $\omega_{\nu,t}(x)-\omega_{\nu_{n},t}(x)$ and $\omega_{\nu,t}^{*}(x)-\omega_{\nu_{n},t}^{*}(x)$ are of order $\varepsilon$. See also Lemma A.3. By the continuity of $S_{\nu}(x)$ for $x<\ell_{\nu}$ (since $\nu$ has compact support), we deduce the convergence of $I_{\nu_{n},t}^{\Lambda}$ to $I_{\nu,t}^{\Lambda}$ away from $\ell_{\nu,t}$.
Step 2: proof of convexity
Let $\nu$ be a probability measure on $\mathbb{R}$ with support bounded from below. Assume first that $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$ so that $\gamma(\Lambda,x)=\Lambda$. A direct derivation shows that
$\partial_{x}I_{\nu,t}^{\Lambda}(x)=\frac{1}{2t}(\omega_{\nu,t}(x)-\Lambda).$
This implies that for all $x<\ell_{\nu,t}$,
$\partial_{x}^{2}I_{\nu,t}^{\Lambda}(x)=\frac{1}{2t}\omega_{\nu,t}^{\prime}(x)>0\,.$ (3.1)
Indeed the second term is clearly positive since
$\frac{\omega_{\nu,t}^{\prime}(x)}{2t}=\frac{1}{1+tG_{\nu}^{\prime}(\omega_{\nu,t}(x))}>0$
on $(-\infty,\ell_{\nu,t})$ since $\omega_{\nu,t}(\ell_{\nu,t})$ satisfies
$\omega_{\nu,t}(\ell_{\nu,t})=\inf\Bigr{\{}x:\int\frac{\mathrm{d}\nu(\lambda)}{(\lambda-x)^{2}}\geq\frac{1}{t}\Bigr{\}}.$
It follows that $I_{\nu,t}^{\Lambda}$ is strictly convex when $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$. Moreover, since $\omega_{\nu,t}(\rho_{\nu,t}^{\Lambda})=\Lambda$, we find $\partial_{x}I_{\nu,t}^{\Lambda}(\rho_{\nu,t}^{\Lambda})=0$. Thus, $I_{\nu,t}^{\Lambda}$ attains its minimum uniquely at $\ell_{\nu,t}^{\Lambda}=\rho_{\nu,t}^{\Lambda}$.
Now assume that $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$. For all $x<\rho_{\nu,t}^{\Lambda}$, we have still $\gamma(\Lambda,x)=\Lambda$ and so as above,
$\partial_{x}I_{\nu,t}^{\Lambda}(x)=\frac{1}{2t}(\omega_{\nu,t}(x)-\Lambda)\leq 0.$
Moreover, for all $x\in[\rho_{\nu,t}^{\Lambda},\ell_{\nu,t})$, we have $\gamma(\Lambda,x)=\omega_{\nu,t}^{*}(x)$ and we find
$\partial_{x}I_{\nu,t}^{\Lambda}(x)=\frac{1}{2t}(\omega_{\nu,t}(x)-\omega^{*}_{\nu,t}(x))\leq 0.$ (3.2)
It follows that for all $x\in(\rho_{\nu,t}^{\Lambda},\ell_{\nu,t})$,
$\partial_{x}^{2}I_{\nu,t}^{\Lambda}(x)=\frac{1}{2t}\Bigr{(}\frac{1}{1+tG_{\nu}^{\prime}(\omega_{\nu,t}(x))}-\frac{1}{1+tG_{\nu}^{\prime}(\omega_{\nu,t}^{*}(x))}\Bigr{)}>0,$
since $\omega_{\nu,t}(x)<\omega_{\nu,t}(\ell_{\nu,t})<\omega_{\nu,t}^{*}(x)$. Together with (3.1), this proves that $I_{\nu,t}^{\Lambda}$ is strictly convex and decreasing. Thus it reaches its minimum uniquely at $\ell^{\Lambda}_{\nu,t}=\ell_{\nu,t}$. ∎
3.2. Proof of Theorem 1.2
The fact that $I_{\nu,t}^{\Lambda}$ is a good rate function follows from the same arguments as in Proposition 2.1, which did not use that $\nu$ was discrete.
We next construct two sequences of matrices with discrete spectrum that bound $B_{N}$ from below and from above. To do so, we first assume that $\nu$ has no atoms. We let $B_{N}^{0}=B_{N}-\Lambda w_{0}w_{0}^{T}$ be the matrix $B_{N}$ without its outlier. First, we take $\varepsilon>0$ and denote $D_{\varepsilon}^{-}(x)=\ell_{\nu}+\lfloor(x-\ell)\varepsilon^{-1}\rfloor\varepsilon$. We set $B_{N,\varepsilon}^{-}=D_{\varepsilon}^{-}(B_{N}^{0})+\Lambda w_{0}w_{0}^{T}$. This matrix has discrete spectrum (since $B_{N}$ is uniformly bounded) with values $\Lambda$ with multiplicity $1$ and $\ell_{\nu}+q\varepsilon$ with multiplicity denoted by $I^{N}_{q}$. We observe that $q\leq N_{\varepsilon}=\sup_{N}|a_{N}-\ell_{\nu}|/\varepsilon$. Note that $I^{N}_{q}/N$ converges towards $\alpha_{q}=\nu([\ell_{\nu}+q\varepsilon,\ell_{\nu}+(q+1)\varepsilon))$ except possibly if $\nu$ has an atom at $\ell_{\nu}+(q+1)\varepsilon$. Because we assumed that $\nu$ has finitely many atoms, we can choose $\varepsilon$ as small as wished and so that $\nu$ has no atoms at $\ell_{\nu}+q\varepsilon$ for all $q\leq N(\varepsilon)$. Moreover, $B_{N}-B_{N,\varepsilon}^{-}$ has operator norm bounded by $2\varepsilon$ for $N$ large enough (since $a_{2}$ goes to $\ell_{\nu}$). Similarly, let $B_{N,\varepsilon}^{+}$ be the diagonal matrix with eigenvalues $B_{N,\varepsilon}^{+}=D_{\varepsilon}^{+}(B_{N}^{0})+\Lambda w_{0}w_{0}^{T}$ if $D_{\varepsilon}^{+}(x)=\ell_{\nu}+\lceil(x-\ell)\varepsilon^{-1}\rceil\varepsilon$.
Clearly
$B_{N,\varepsilon}^{-}\leq B_{N}\leq B_{N,\varepsilon}^{+},$
so that for every real number $x$
$\mathbb{P}\left(\lambda_{1}(B_{N,\varepsilon}^{-}+G_{N})\geq x\right)\leq\mathbb{P}\left(\lambda_{1}(B_{N}+G_{N})\geq x\right)\leq\mathbb{P}\left(\lambda_{1}(B_{N,\varepsilon}^{+}+G_{N})\geq x\right).$ (3.3)
Next, we can apply Theorem 2.1 to $B_{N,\varepsilon}^{-}+G_{N}$ and $B_{N,\varepsilon}^{+}+G_{N}$ We also observe that
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\delta_{\lambda_{i}(B_{N,\varepsilon}^{\pm})}=\nu_{\varepsilon}^{\pm}:=D_{\varepsilon}^{\pm}(\eta_{\nu})\#1_{[0,1]}\mathrm{d}x.$
Consequently we can use Proposition 2.1 to deduce from (3.3) that
$-\inf_{[x,+\infty)}I_{\nu_{\varepsilon}^{-},t}^{\Lambda}\leq\liminf_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}\left(\lambda_{1}(B_{N}+G_{N})\geq x\right)\leq\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}\left(\lambda_{1}(B_{N}+G_{N})\geq x\right)\\ \leq-\inf_{[x,+\infty)}I_{\nu_{\varepsilon}^{+},t}^{\Lambda}.$
We now observe that $\eta_{\nu_{\varepsilon}^{\pm}}-\eta_{\nu}$ converges uniformly towards zero so that $I_{\nu_{\varepsilon}^{\pm},\Lambda}-I_{\nu,t}^{\Lambda}$ goes to zero by the second point of Lemma 3.1. Therefore, since $I_{\nu_{\varepsilon}^{\pm},\Lambda}$ goes to infinity like $x^{2}/4t$, we also deduce that
$\lim_{\varepsilon\rightarrow 0}\inf_{[x,+\infty)}I_{\nu_{\varepsilon}^{\pm},\Lambda}=\inf_{[x,+\infty)}I_{\nu,t}^{\Lambda}$
and the same when the infimum is taken on $(x,+\infty)$. Therefore, letting $\varepsilon$ going to zero, (3.4) implies that
$-\inf_{(x,+\infty)}I_{\nu,t}^{\Lambda}\leq\liminf_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}\left(\lambda_{1}(B_{N}+G_{N})\geq x\right)\leq\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}\left(\lambda_{1}(B_{N}+G_{N})\geq x\right)\\ \leq-\inf_{[x,+\infty)}I_{\nu,t}^{\Lambda}.$ (3.4)
Let $\delta>0$ such that $x+\delta\leq\ell$. It follows that since $I_{\nu,t}^{\Lambda}$ is strictly increasing above ${\ell_{\nu,t}^{\Lambda}}$ by Lemma 3.1, for $\kappa>0$ small and $N$ large enough
$P\left(\lambda_{1}(B_{N}+G_{N})\geq x+\delta\right)$$\leq$ $e^{-N\inf_{(-\infty,x+\delta]}(I_{\nu,t}^{\Lambda}-\kappa)}\leq e^{-N\inf_{(-\infty,x]}(I_{\nu,t}^{\Lambda}+\kappa)}$
$\leq$ $e^{-N\kappa/2}P\left(\lambda_{1}(B_{N}+G_{N})\geq x\right)\,.$
Therefore
$P\left(\lambda_{1}(B_{N}+G_{N})\geq x\right)=P\left(\lambda_{1}(B_{N}+G_{N})\in(x,x+\delta)\right)(1+o(1))\,.$
As a consequence, we deduce from (3.4) that for $x\geq\rho_{\nu,t}^{\Lambda}$,
$-\inf_{(x,x+\delta)}I_{\nu,t}^{\Lambda}\leq\liminf_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}\left(\lambda_{1}(B_{N}+G_{N})\in[x,x+\delta)\right)\\ \leq\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}\left(\lambda_{1}(B_{N}+G_{N})\in[x,x+\delta)\right)\leq-\inf_{[x,x+\delta)}I_{\nu,t}^{\Lambda}.$ (3.5)
Since $I_{\nu,t}^{\Lambda}$ is lower semicontinuous we have
$\lim_{\delta\to 0}\inf_{(x,x+\delta)}I_{\nu,t}^{\Lambda}=\lim_{\delta\to 0}\inf_{[x,x+\delta)}I_{\nu,t}^{\Lambda}=I_{\nu,t}^{\Lambda}(x).$
Hence, letting $\delta$ going to zero, (3.5) gives the weak large deviations estimate above ${\ell_{\nu,t}^{\Lambda}}$. Similarly we find by the above comparison that
$-\inf_{(-\infty,x]}I_{\nu_{\epsilon}^{+},\Lambda}\leq\liminf_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}\left(\lambda_{1}(B_{N}+G_{N})\leq x\right)\leq\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}\left(\lambda_{1}(B_{N}+G_{N})\leq x\right)\\ \leq-\inf_{(-\infty,x]}I_{\nu_{\varepsilon}^{-},\Lambda}.$
We can also use the convergence of $I_{\nu_{\varepsilon}^{\pm},\Lambda}$ towards $I_{\nu^{\pm},\Lambda}$ and the fact that $I_{\nu,t}^{\Lambda}$ is strictly decreasing below ${\ell_{\nu,t}^{\Lambda}}$ to deduce the weak large deviation principle below ${\ell_{\nu,t}^{\Lambda}}$. This proves the weak LDP everywhere. This result can be upgraded to a full LDP using exponential tightness as in the proof of Proposition 2.1 and the fact that $I_{\nu,t}^{\Lambda}$ is a good rate function. Lemma 3.1 allows to complete the proof of Theorem 1.2.
4. Comparison with prior results
In this section we show that our rate function agrees with [Mai07, McK21].
4.1. The case of the GOE with an outlier
We first consider the case where $B_{N}=\Lambda e_{1}e_{1}^{T}$ for $\Lambda<0$ and compare our result with [Mai07]. Let $G_{N}$ be a GOE of variance $t>0$. We study the smallest eigenvalue of $G_{N}+\Lambda e_{1}e_{1}^{T}$. Note that $\sigma_{t}=\sigma_{t}\boxplus\delta_{0}$. Therefore, the characteristic associated to the Burgers equation, see Remark A.1, is
$H_{0,t}:x\mapsto x+\frac{t}{x}.$
The characteristic has a shock at $x=-\sqrt{t}$ and the corresponding value of $H_{0,t}$ is $H_{0,t}(x)=-2\sqrt{t}$. The support of $\sigma_{t}$ is therefore $[-2\sqrt{t},2\sqrt{t}]$. Let $\omega_{0,t}$ be the inverse of $H_{0,t}$ on $(-\infty,-2\sqrt{t}]$, namely
$\omega_{0,t}:\lambda\in(-\infty,-2\sqrt{t}]\mapsto\frac{\lambda+\sqrt{\lambda^{2}-4t}}{2}.$
With the normalization of [Mai07], taking $\beta=1$ gives a LDP for the smallest eigenvalue of $G_{N}+\Lambda e_{1}e_{1}^{T}$ where $G_{N}$ is GOE of variance $t=1/2$. In the sequel we take $t=1/2$. Denote $\rho_{\nu,t}^{\Lambda}:=\Lambda+\frac{1}{2\Lambda}$. Let $J_{\Lambda}$ be the rate function found in [Mai07], given if $\Lambda\leq-\frac{1}{\sqrt{2}}$ by
$J_{\Lambda}(x):=\begin{cases}M_{\Lambda}(x)&\text{if $x\leq-\sqrt{2}$}\\ +\infty&\text{if $x>-\sqrt{2}$},\end{cases}$
where
${M_{\Lambda}}(x):=-\frac{1}{2}\int_{x}^{\rho_{\nu,t}^{\Lambda}}\sqrt{z^{2}-2}\mathrm{d}z-\Lambda(x-\rho_{\nu,t}^{\Lambda}))+\frac{1}{4}(x^{2}-\rho_{\nu,t}^{\Lambda})$
and if $\Lambda\geq-\frac{1}{\sqrt{2}}$ by
$J_{\Lambda}(x):=\begin{cases}N_{\Lambda}(x)&\text{if $x\leq\rho_{\nu,t}^{\Lambda}$}\\ \int_{x}^{-\sqrt{2}}\sqrt{z^{2}-2}\mathrm{d}z&\text{if $\rho_{\nu,t}^{\Lambda}\leq x\leq-\sqrt{2}$}\\ +\infty&\text{if $x>-\sqrt{2}$},\end{cases}$
where
${N_{\Lambda}}(x):=-\frac{1}{2}\int_{x}^{-\sqrt{2}}\sqrt{z^{2}-2}\mathrm{d}z-\Lambda x+\frac{1}{4}x^{2}+\frac{1}{4}+\frac{1}{4}\log 2+\frac{\Lambda^{2}}{2}+\frac{1}{2}\log\Lambda.$
Lemma 4.1.
We have
$J_{\Lambda}=I_{0,1/2}^{\Lambda}$
where we denoted in short $I_{0,1/2}^{\Lambda}$ for $I_{\delta_{0},1/2}^{\Lambda}$.
Proof.
For all $x<-\sqrt{2}$, let $\gamma_{1}(x)\leq\gamma_{2}(x)$ be the two solutions in $\mathbb{R}^{-}$ of the equation $\gamma+\frac{1}{2\gamma}=x$, namely
$\gamma_{1}(x)=\frac{x-\sqrt{x^{2}-2}}{2}\quad\text{and}\quad\gamma_{2}(x)=\frac{x+\sqrt{x^{2}-2}}{2}.$
Let $x<-\sqrt{2}$. Assume that $\Lambda\leq-\frac{1}{\sqrt{2}}$, or $\Lambda\geq-\frac{1}{\sqrt{2}}$ and $x<-\sqrt{2}$. One can note that
$J_{\Lambda}^{\prime}(x)=-\frac{1}{2}\sqrt{x^{2}-2}+\frac{x}{2}-\Lambda.$
Moreover
$\partial_{x}I_{0,1/2}^{\Lambda}(x)=(\omega_{\nu,1/2}(x)-\Lambda)=-\frac{1}{2}\sqrt{x^{2}-2}+\frac{x}{2}-\Lambda=J_{\Lambda}^{\prime}(x).$
Assume that $\Lambda\geq-\frac{1}{\sqrt{2}}$ and $x>\rho_{\nu,1/2}^{\Lambda}$. By Equation (3.2),
$\partial_{x}I_{0,1/2}^{\Lambda}(x)=\omega_{\nu,1/2}(x)-\omega_{\nu,1/2}^{*}(x)=-\sqrt{x^{2}-2}=J_{\Lambda}^{\prime}(x).$
Since $J_{\Lambda}$ and $I_{0,1/2}^{\Lambda}$ are $\mathcal{C}^{1}$ on $(-\infty,-\sqrt{2})$ and $\inf J_{\Lambda}=\inf I_{0,1/2}^{\Lambda}=0$, we conclude that $J_{\Lambda}=I_{0,1/2}^{\Lambda}$. ∎
4.2. The no outlier case
Let us now compare our result with [McK21] where no outlier is permitted. Again we take $\beta=1$ in [McK21]. In the normalization of [McK21], the variance of the non-diagonal entry distribution is $1$. In our setting (1.3), this corresponds to taking $t=1$ (instead of $t=1/2$ in the normalization of [Mai07]).
We apply Theorem 1.2 to $\Lambda:=\ell_{\nu}$. The rate function of the LDP of [McK21] is expressed variationally as
$J_{\nu}:x\mapsto\sup_{\theta\geq 0}J(x,\theta),$ (4.1)
where for all $x\in\mathbb{R}$, $\theta\in\mathbb{R}^{+}$,
$J(x,\theta):=J(\theta,\nu\boxplus\sigma_{t},x)-J(\theta,\nu,x)-\theta^{2},$
where for all $x\in\mathbb{R}$, $\theta\in\mathbb{R}^{+}$ and probability measure $\mu$ on $\mathbb{R}$,
$J(\theta,\mu,x):=\begin{cases}\frac{1}{2}\int_{0}^{2\theta}R_{\mu}(s)\mathrm{d}s&\text{if $0\leq 2t\theta\leq G_{\mu}(x)$}\\ \theta x-\frac{1}{2t}(1+\log(2\theta))-\frac{1}{2}\int\log|x-y|\mathrm{d}\mu(y)&\text{if $2\theta>G_{\mu}(x)$},\end{cases}$
where $R_{\mu}$ is the Voiculescu $R$-transform, which is given by $R_{\mu}:y\mapsto K_{\mu}(y)-\frac{1}{y}$ with $K_{\mu}$ as in (1.12).
Lemma 4.2.
We have
${J_{\nu}=I_{\nu,1}^{\ell_{\nu}}.}$
Proof.
Let $\rho_{\nu,{1}}^{\Lambda}:=\ell_{\nu}+G_{\nu}(\ell_{\nu})\in\mathbb{R}\cup\{-\infty\}$. For all $x\leq\ell_{\nu,1}$, let $\gamma_{1}(x)\leq\gamma_{2}(x)$ be the two solutions of the equation $H_{\nu,1}(\gamma)=x$ on $(-\infty,\ell_{\nu})$.
First observe that for all $x<\ell_{\nu,1}$,
$J_{\nu}^{\prime}(x)=\partial_{x}J_{\nu}(x,\theta(x))+\partial_{\theta}J_{\nu}(x,\theta)\theta^{\prime}(x)=\partial_{x}J_{\nu}(x,\theta(x))=\theta(x)-\frac{1}{2}G_{\nu\boxplus\sigma_{t}}(x)=\theta(x)-\frac{1}{2}G_{\nu}(\gamma_{1}(x)).$ (4.2)
where we have applied Lemma A.4 in the last equality.
Let $x\in(\rho_{\nu,{1}}^{\Lambda},\ell_{\nu,1})$. As shown in [McK21, (2.14)], for $x\in(\rho_{\nu,t}^{\Lambda},\ell_{\nu,1})$, the optimal $\theta(x)$ in (4.1) is the unique solution of the equation
$2\theta(x)+K_{\nu}(2\theta)=x,\quad\theta(x)\in\Bigr{(}\frac{1}{2}G_{\nu}(\ell_{\nu}),\frac{1}{2}G_{\nu\boxplus\sigma_{t}}(\ell_{\nu,t})\Bigr{)},$
where $K_{\nu}$ is as in (1.12). One can easily show
$\theta(x)=\frac{1}{2}G_{\nu}(\gamma_{2}(x)).$
Substituting this into (4.2), we find that for all $x\in(\rho_{\nu,t}^{\Lambda},\ell_{\nu,1})$,
$J_{\nu}^{\prime}(x)=G_{\nu}(\gamma_{2}(x))-G_{\nu}(\gamma_{1}(x))=\frac{1}{2}(\gamma_{1}(x)-\gamma_{2}(x))=\partial_{x}I_{\nu,1}^{\ell_{\nu}}(x),$
where we have used (3.2) in the last equality.
If $x\leq\rho_{\nu,t}^{\Lambda}$, then by [McK21],
$\theta(x)=\frac{1}{2}(x-\ell_{\nu}),$
which implies that for all $x<\rho_{\nu,t}^{\Lambda}$,
$J_{\nu}^{\prime}(x)=\frac{1}{2}(x-\Lambda-G_{\mu}(\gamma_{1}(x)))=\frac{1}{2}(\gamma_{1}(x)-\Lambda)=\partial_{x}I_{\nu,1}^{\ell_{\nu}}(x).$
Since $J_{\nu}$ and ${I_{\nu,1}^{\ell_{\nu}}}$ are $\mathcal{C}^{1}$ on $(-\infty,\ell_{\nu,1})$ and $\inf J_{\nu}=\inf{I_{\nu,1}^{\ell_{\nu}}}=0$, this proves that $J_{\nu}=I_{\nu,1}^{\ell_{\nu}}$. ∎
Appendix A Reminder on the free convolution
We recall some classical results on free convolution from [Bia97] that will be useful throughout this article. Recall the definitions of $v_{t}(u)$ and $\Omega_{\mu,t}$ in (1.2) and (1.1). We have the following:
Lemma A.1 (Biane).
Let $\mu$ be a probability measure on $\mathbb{R}$. Let $t>0$. The function $H_{\mu,t}$ is a homeomorphism, which is conformal from $\Omega_{\mu,t}$ onto $\mathbb{C}^{+}$. We let $\omega_{\mu,t}:{\mathbb{C}^{+}\cup\mathbb{R}}\to\overline{\Omega_{\mu,t}}$ be its inverse, called the subordination function. Moreover, for every $z\in\Omega_{\mu,t}$,
$G_{\mu\boxplus\sigma_{t}}(z+tG_{\mu}(z))=G_{\mu}(z).$ (A.1)
Assume that the support of $\mu$ is bounded from below. The left edge of $\mu\boxplus\sigma_{t}$ is given by $\ell_{\mu,t}=H_{\mu,t}(\omega)$ where
$\omega:=\inf\Bigr{\{}\omega\leq\ell_{\mu}:\int\frac{\mathrm{d}\mu(\lambda)}{(\lambda-\omega)^{2}}\geq\frac{1}{t}\Bigr{\}}.$ (A.2)
For every $x<\omega_{\mu,t}(\ell_{\mu,t})$, we have
$\omega_{\mu,t}(x)=x-tG_{\mu\boxplus\sigma_{t}}(x).$ (A.3)
Remark A.1.
Lemma A.1 can be reformulated in the langage of PDE’s. The Stieltjes transform of $\mu_{t}:=\mu\boxplus\sigma_{t}$ indeed satisfies a complex Burgers equation: for any $z$ outside the support of $\mu_{t}$ we have
$\partial_{t}G_{\mu_{t}}(z)-G_{\mu_{t}}(z)\partial_{z}G_{\mu_{t}}(z)=0.$ (A.4)
This equation can be solved via the method of characteristics. The characteristic starting from $z\in\mathbb{C}^{+}$ is given by the straight line $t\mapsto z+tG_{\mu}(z)$. Equation (A.1) reflects the fact that the solution is constant along characteristics (after restriction to suitable $z$). Let $\omega\in(-\infty,\ell_{\mu}].$ One may note that $z_{t}$ is injective on $(-\infty,\omega)$ provided
$\int\frac{\mathrm{d}\mu(\lambda)}{(\lambda-\omega)^{2}}<\frac{1}{t}.$
From (A.2), the left edge of $\mu_{t}$ can be interpreted as the first value at which the solution develops a shock at time $t$.
Property A.2.
By Lemma A.1, the map $H_{\mu,t}|_{(-\infty,\ell_{\mu})}$ is strictly concave, strictly increasing on the set $(-\infty,\omega_{\mu,t}(\ell_{\mu,t}))$ and strictly decreasing on the set $(\omega_{\mu,t}(\ell_{\mu,t}),+\infty)$. In addition $H_{\mu,t}$ goes to $\infty$- ∞ when $x$ goes to $\infty$- ∞ and to $H_{\mu,t}(\ell_{\mu})\in\mathbb{R}\cup\{-\infty\}$ when $x$ goes to $\ell_{\mu}$.
It follows that, for all $\lambda\in(H_{\mu,t}(\ell_{\mu}),\ell_{\mu,t})$, the equation
$H_{\mu,t}(\gamma)=\lambda,\quad\gamma<\ell_{\mu}$
admits two solutions $\gamma_{1}<\gamma_{2}$. Moreover if $\lambda\in(H_{\mu,t}(\ell_{\mu}),\ell_{\mu,t})$, we have $\gamma_{1}=\omega_{\mu,t}(\lambda)<\omega_{\mu,t}(\ell_{\mu,t})$ and $\gamma_{2}=\omega_{\nu,t}^{*}(\lambda)>\omega_{\mu,t}(\ell_{\mu,t})$.
Throughout this paper, we used several smoothness results that we summarize and prove here:
Lemma A.3.
  1. (1)
    Let $M$ be a positive finite real number. Then $\mu{\mapsto}\mu\boxplus\sigma_{t}$ is smooth for the weak topology on $\mathcal{P}([-M,M])$.
  2. (2)
    Moreover, for $\eta_{1}<\cdots,\eta_{p}$, denoting $\nu_{\alpha}=\sum\alpha_{i}delta_{\eta_{i}}$, we have for every $t>0$ and $\alpha,\alpha^{\prime}\in\Delta^{p}$,
    $d(\nu_{\alpha}\boxplus\sigma_{t},\nu_{\alpha^{\prime}}\boxplus\sigma_{t})\leq\sum|\alpha_{i}-\alpha_{i}^{\prime}|\,.$
  3. (3)
    For every probability measure $\nu\in\mathcal{P}([-M,M])$, $x\mapsto\omega_{\nu,t}(x)$ is continuous and increasing on $(-\infty,\ell_{\nu,t}]$ with values in $(-\infty,\omega_{\nu,t}(\ell_{\nu,t})]$, and $x\mapsto\omega_{\nu,t}^{*}(x)$ is continuous and increasing on $(-\infty,\ell_{\nu,t})$ with values in $[\omega_{\nu,t}(\ell_{\nu,t}),\ell_{\nu}]$.
  4. (4)
    For every probability measure $\nu\in\mathcal{P}([-M,M])$, let $\eta_{\nu}$ be such that $\nu=\eta_{\nu}\#1_{[0,1]}\mathrm{d}x$. Then, for every $\nu,\nu^{\prime}\in\mathcal{P}([-M,M])$,
    $d(\nu\boxplus\sigma_{t},\nu^{\prime}\boxplus\sigma_{t})\leq\|\eta_{\nu}-\eta^{\prime}_{\nu}\|_{\infty}\mbox{ and }|\ell_{\nu,t}-\ell_{\nu^{\prime},t}|\leq\|\eta_{\nu}-\eta_{\nu^{\prime}}\|_{\infty}\,.$
    Moreover if $x<\ell_{\nu,t}$, $\omega_{\nu,t}(x)-\omega_{\nu^{\prime},t}(x)$ and $\omega^{*}_{\nu,t}(x)-\omega^{*}_{\nu^{\prime},t}(x)$ go to zero when $\|\eta_{\nu}-\eta_{\nu^{\prime}}\|_{\infty}$ goes to zero.
Proof.
$\bullet$ The first point is clear since it is enough to show that $\mu\mapsto G_{\mu\boxplus\sigma_{t}}(z)$ is continuous for $z\in[-M^{\prime},M^{\prime}]^{c}$ for some $M^{\prime}>M+2\sqrt{t}$ finite (since the analytic function $G_{\mu\boxplus\sigma_{t}}$ is uniquely defined on its domain of analyticity from its value on the set $z\in[-M^{\prime},M^{\prime}]^{c}$). But this is clear because then $\mu\mapsto G_{\mu}(z)$ is uniformly continuous and close to $z^{-1}$ on $[-M^{\prime},M^{\prime}]^{c}$ so that $H_{\mu,t}(z)$ is close to $z$ and smoothly invertible with smooth inverse $\omega_{\mu,t}(z)$ and then for $z$ large enough
$\mu\mapsto G_{\mu\boxplus\sigma_{t}}(z)=G_{\mu}(\omega_{\mu,t}(z))$
is uniformly continuous in $\mu$.
$\bullet$ To prove the second point, we note that $\nu_{\alpha}\boxplus\sigma_{t}$ is the distribution of the spectrum of $B_{\alpha}+S_{t}$ where $B_{\alpha}$ is distributed according to $\nu_{\alpha}$, free with the semicircular variable $S_{t}$ with variance $t$. But by definition of $\nu_{\alpha}$, one can couple $B_{\alpha}$ and $B_{\alpha^{\prime}}$ so that $B_{\alpha}-B_{\alpha^{\prime}}$ has rank bounded by $\sum|\alpha_{i}-\alpha_{i}^{\prime}|$. The result follows.
$\bullet$ The function $H_{\nu,t}$ is smooth on $(-\infty,\ell_{\nu})$, strictly increasing on $(-\infty,\omega)$ and strictly decreasing on $(\omega,\ell_{\nu})$ with ${\omega\leq\ell_{\nu}}$ the point where
$\int\frac{{\mathrm{d}\nu(x)}}{(\omega-x)^{2}}=\frac{1}{t},$
if it is strictly smaller than $\ell_{\nu}$, $\omega=\ell_{\nu}$ otherwise. Because the derivative of $H_{\nu,t}$ is strictly positive on $(-\infty,\omega)$, the implicit function theorem shows that $H_{\nu,t}$ has an inverse $\omega_{\nu,t}$ with values in $(-\infty,\omega)$ which is continuous. Similarly, because the derivative of $H_{\nu,t}$ is strictly negative on $(\omega,\ell_{\nu})$, $H_{\nu,t}$ has a second continuous inverse with values in $(\omega,\ell_{\nu})$. By [Bia97], we know that $\ell_{\nu,t}=H_{\nu,t}(\omega)$ so $\omega=\omega_{\nu,t}(\ell_{\nu,t})$.
$\bullet$ For the last point, we can prove the bound on $d(\nu\boxplus\sigma_{t},\nu^{\prime}\boxplus\sigma_{t})$ again by coupling them as the distribution of $\eta_{\nu}(\Lambda)+S_{t}$ and $\eta_{\nu^{\prime}}(\Lambda)+S_{t}$ where $\Lambda$ is distributed according to $1_{[0,1]}\mathrm{d}x$ is free from $S_{t}$. The result follows. Moreover, $\ell_{\nu}=\inf\{\eta_{\nu}\}$ is continuous in $\eta_{\nu}$. Furthermore $\ell_{\nu,t}$ is the smallest eigenvalue of $X_{t}(\nu)=\eta_{\nu}(\Lambda)+S_{t}$ and $\|X_{t}(\nu)-X_{t}(\nu^{\prime})\|_{\infty}\leq\|\eta_{\nu}-\eta_{\nu^{\prime}}\|_{\infty}$ so that $|\ell_{\nu,t}-\ell_{\nu^{\prime},t}|\leq\|\eta_{\nu}-\eta_{\nu^{\prime}}\|_{\infty}$. The last statement regarding the continuity of $\omega_{\nu,t}(x)$ and $\omega^{*}_{\nu,t}(x)$ can be deduced from the following observations: First, we can assume $x<\min\{\ell_{\nu,t},\ell_{\nu^{\prime},t}\}$, which ensures that both $\omega_{\nu,t}(x)$ and $\omega^{*}_{\nu,t}(x)$ are away from $\ell_{\nu}$. Additionally, the inverse $H_{\nu,t}$ has a non vanishing derivative on this range. Therefore, by applying the implicit function theorem and considering that $H_{\nu,t}-H_{\nu^{\prime},t}$ is small, we obtain the desired result. ∎
The PDE interpretation discussed in the Remark A.1 allows one to compute the logarithmic potential of the measure $\mu\boxplus\sigma_{t}$ through the Hopf-Lax formula.
Lemma A.4.
Let $\mu$ be a probability measure on $\mathbb{R}$ with support bounded from below and $t>0$.
  1. (1)
    Let $x\in\mathbb{R}$ and $\omega:=\omega_{\mu,t}(x)\in\mathbb{C}^{+}\cup\mathbb{R}$. Let $(z_{s})_{s\in[0,t]}$ given for all $s\in[0,t]$ by
    $z_{s}=\omega+sG_{\mu}(\omega).$ (A.5)
    If $x\leq\ell_{\mu,t}$, then for all $s\in[0,t]$,
    $x\leq z_{s}\leq\ell(\mu\boxplus\sigma_{s}).$ (A.6)
  2. (2)
    Let $x\in\mathbb{R}$. Assume that $x\leq\ell_{\mu,t}$. We have
    $\int\log|\lambda-x|\mathrm{d}(\mu\boxplus\sigma_{t})(\lambda)=\int\log|\lambda-\omega_{\mu,t}(x)|\mathrm{d}\mu(\lambda)+\frac{(\omega_{\mu,t}(x)-x)^{2}}{2t}.$ (A.7)
Proof.
Let $x\in\mathbb{R}$ and $\omega:=\omega_{\nu,t}(x)$. Consider
$z_{s}=\omega+sG_{\mu}(\omega),\quad 0\leq s\leq t.$
Assume first that $x\leq\ell(\mu\boxplus\sigma_{t})$. First by [GM20, Lemma 11], we have
$\omega_{\mu,t}((-\infty,\ell(\mu\boxplus\sigma_{t})])\subset(-\infty,\ell(\mu)]$
and therefore $\omega\leq\ell(\mu)$. This proves that $z_{0}\leq\ell(\mu)$. Let $s\in(0,t)$. Recall that $\mu\boxplus\sigma_{t}=(\mu\boxplus\sigma_{s})\boxplus\sigma_{t-s}$. We have
$\omega_{\mu\boxplus\sigma_{s},t-s}(x)=z_{s}.$
Indeed,
$z_{s}+(t-s)G_{\mu\boxplus\sigma_{s}}(z_{s})=z_{s}+(t-s)G_{\mu}(\omega)=\omega+sG_{\mu}(\omega)+(t-s)G_{\mu}(\omega)=x.$
Therefore, applying the previous result to $t^{\prime}=t-s$ and $\mu^{\prime}=\mu\boxplus\sigma_{s}$, we obtain
$z_{s}\leq\ell(\mu\boxplus\sigma_{s}).$
Because $\omega\leq\ell(\mu)$, $G_{\mu}(\omega)$ is non-positive and therefore $z_{s}$ is decreasing which shows that $z_{s}\geq z_{t}=H_{\mu,t}(B)=x$ for all $0\leq s\leq t$. This completes the proof of (A.6).
Fix $x\in\mathbb{R}$ and let $\omega:=\omega_{\mu,t}(x)$. For all $s\geq 0$, let $m_{s}:=G_{\mu\boxplus\sigma_{s}}$. It is well-known, see for instance [Bia97], that for all $z\in\mathbb{C}^{+}\cup\mathbb{R}$ outside the support of $\mu\boxplus\sigma_{s}$,
$\partial_{s}m_{s}(z)-m_{s}(z)\partial_{z}m_{s}(z)=0$
and that $m_{s}(z_{s})=m_{0}(z_{0})=G_{\mu}(\omega)$. Define
$F:(s,z)\in\{(s,z)\in\mathbb{R}^{+}\times\mathbb{R}:z<\ell_{\mu,s}\}\mapsto\int\log|\lambda-z|\mathrm{d}(\mu\boxplus\sigma_{s})(\lambda)-\frac{(z-\omega)^{2}}{2s}.$
Let $(z_{s})_{s\in[0,t]}$ be as in (A.5). By [BABM24, Lemma 5.4], for all $z\in\mathbb{C}^{+}\cup(-\infty,\ell_{\mu,t})$, we have
$\partial_{s}F(s,z)=-\frac{1}{2}m_{s}(z)^{2}+\frac{(z-\omega)^{2}}{2s^{2}}.$
As proved in item (1), for all $s\in[0,t)$, $z_{s}<\ell_{\mu,s}$. Hence, since $m_{s}(z_{s})=m_{0}(z_{0})=G_{\mu}(\omega)$ and $z_{s}-\omega=sG_{\mu}(\omega)$,
$\partial_{s}F(s,z_{s})=0\,.$
It follows that
$\frac{\mathrm{d}}{\mathrm{d}s}F(s,z_{s})=0,$
which gives $F(0,\omega)=F(t,x)$, or equivalently
$\int\log|\lambda-x|\mathrm{d}(\mu\boxplus\sigma_{t})(\lambda)=\int\log|\lambda-\omega_{\mu,t}(x)|\mathrm{d}\mu(\lambda)+\frac{(\omega_{\mu,t}(x)-x)^{2}}{2t},$
which proves (A.7). ∎
Appendix B A priori properties on the rate function
Lemma B.1.
Let $(B_{N})$ satisfying Assumptions 1.1 with $\nu$ and $\Lambda$ be as in (1.4) and (1.5). Let $P$ be the law of the eigenvalues of $B_{N}+G_{N}$ where $G_{N}$ is a $N\times N$ GOE variance matrix with variance $t$. Recall
$\ell_{\nu,t}^{\Lambda}:=\begin{cases}\rho_{\nu,t}^{\Lambda}&\text{if $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$}\\ \ell_{\nu,t}&\text{if $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$}.\end{cases}$
Let $x<y<\ell_{\nu,t}^{\Lambda}$ or $\ell_{\nu,t}^{\Lambda}<y<x$. Then, for every $\delta\in(0,|\ell^{\Lambda}_{\nu,t}-y|)$:
$\limsup_{N\rightarrow\infty}\frac{1}{N}\log P(|\lambda_{1}-x|\leq\delta)\leq\limsup_{N\rightarrow\infty}\frac{1}{N}\log P(|\lambda_{1}-y|\leq\delta)\,.$
Proof.
We may and shall assume that $\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-x|\leq\delta)$ is finite since otherwise we are done. The proof follows the proof of [CDG23, Lemma 5.2] based on [CDG23, Lemma 8.1] where a coupling is constructed. Because we restrict ourselves to Gaussian entries in this article, we propose a simpler version based on the symmetric Ornstein-Uhlenbeck
${\mathrm{d}G_{N}(s)=\mathrm{d}H_{N}(s)-\frac{1}{t}G_{N}(s)\mathrm{d}s}$
[…]
${\mathrm{d}(G_{N}(s)-\bar{G}_{N}(s))^{2}=-2(G_{N}(s)-\bar{G}_{N}(s))^{2}\mathrm{d}t}$
implies that
$\|G_{N}(s)-\bar{G}_{N}(s)\|_{\infty}\leq\|G_{N}\|_{\infty}e^{-s}$ (B.1)
where $\|.\|_{\infty}$ denotes the operator norm. We will later take $s=N$. Moreover we work on the set where $\|G_{N}\|_{\infty}\leq N$ so that $\|G_{N}(t)-\bar{G}_{N}(t)\|_{\infty}$ is of order $Ne^{-N}\leq\delta/2$ if $t$ is of order $N$. We finally set $X_{N}(s):=B_{N}+G_{N}(s)$ and $\bar{X}_{N}(s):=B_{N}+\bar{G}_{N}(s)$.
[…]
$\mathbb{P}(|\mathbb{E}[\lambda_{1}(\bar{X}_{N}(N))]-\lambda_{1}(\bar{X}_{N}(N))|\geq\eta)\leq 2e^{-\frac{N\eta^{2}}{4}}{.}$ (B.2)
We deduce that for $N\gg\log\delta^{-1}$ large enough,
$Q_{N}:=\mathbb{P}(|\lambda_{1}(X_{N}(N))-{\ell}_{\nu,t}^{\Lambda}|\leq\delta\big{|}|\lambda_{1}(X_{N}(0))-x|\leq\delta,\|G_{N}\|_{\infty}\leq N)\geq 1/2{.}$
Indeed, by (B.1) which implies that $\|X_{N}(N)-\bar{X}_{N}(N)\|_{\infty}$ is of order $Ne^{-N}$, for $N$ large enough
$\begin{split}Q_{N}&\geq\frac{\mathbb{P}(\{|\lambda_{1}(\bar{X}_{N}(N))-{\ell}_{\nu,t}^{\Lambda}|\leq\delta/2\}\cap\{|\lambda_{1}(X_{N}(0))-x|\leq\delta,\|G_{N}\|_{\infty}\leq N\})}{\mathbb{P}(|\lambda_{1}(X_{N}(0))-x|\leq\delta,\|G_{N}\|_{\infty}\leq N\})}\\ &=\mathbb{P}(\{|\lambda_{1}(\bar{X}_{N}(N))-{\ell}_{\nu,t}^{\Lambda}|\leq\delta/2\}){.}\end{split}$
where we used that $\bar{X}_{N}$ is independent from $X_{N}(0)=G_{N}$. Finally, we have seen that the above right-hand side goes to $1$ as $N$ goes to infinity. The fact that $Q_{N}$ is bounded from below by $1/2$ implies that
$\begin{split}&\mathbb{P}(|\lambda_{1}(X_{N}(0))-x|\leq\delta,\|G_{N}\|_{\infty}\leq N)\\ &\leq 2\mathbb{P}(\{|\lambda_{1}(X_{N}(N))-{\ell}_{\nu,t}^{\Lambda}|\leq\delta/2\}\cap\{|\lambda_{1}(X_{N}(0))-x|\leq\delta,\|G_{N}\|_{\infty}\leq N\})\\ &\leq 2\mathbb{P}(\{|\lambda_{1}(X_{N}(N))-{\ell}_{\nu,t}^{\Lambda}|\leq\delta/2\}\cap\{|\lambda_{1}(X_{N}(0))-x|\leq\delta,\|G_{N}\|_{\infty}\leq N\}\cap\Omega_{N})+\mathbb{P}(\Omega_{N}^{c})\end{split}$
where
${\Omega_{N}=\bigcap_{0\leq p\leq N^{2}-1}\Bigr{\{}\Bigr{\|}X_{N}(\frac{p+1}{N})-X_{N}(\frac{p}{N})\Bigr{\|}_{\infty}\leq\delta\Bigr{\}}.}$
Since $X_{N}(\frac{p+1}{N})-X_{N}(\frac{p}{N})$ is a GOE matrix with variance $t$, we find by (B.2) that $\mathbb{P}(\Omega_{N}^{c})$ is bounded above by $2N^{2}e^{-c(\delta)N^{2}}$ for some $c(\delta)>0$ provided $\delta>0$ is independent of $N$. Moreover, since $y$ lies in between $x$ and ${\ell}_{\nu,t}^{\Lambda}$, we see that
$\{|\lambda_{1}(X_{N}(N))-{\ell}_{\nu,t}^{\Lambda}|\leq\delta/2\}\cap\{|\lambda_{1}(X_{N}(0))-x|\leq\delta,\|G_{N}\|_{\infty}\leq N\}\cap\Omega_{N}\\ \subset{\bigcup_{1\leq p\leq N^{2}}\Bigr{\{}|\lambda_{1}(X_{N}(\frac{p}{N}))-y|\leq\delta\Bigr{\}}.}$
Hence, we deduce that
$\begin{split}\mathbb{P}(|\lambda_{1}(X_{N}(0))-x|\leq\delta,\|G_{N}\|_{\infty}\leq N)&\leq 2\mathbb{P}(\cup_{1\leq p\leq N^{2}}\{|\lambda_{1}(X_{N}(\frac{p}{N}))-y|\leq\delta\})+4N^{2}e^{-c(\delta)N^{2}}\\ &\leq 2N^{2}\mathbb{P}(\{|\lambda_{1}(X_{N}(0))-y|\leq\delta\})+4N^{2}e^{-c(\delta)N^{2}},\end{split}$
where we used that $X_{N}(\frac{p}{N})$ has the same law as $X_{N}(0)$. Finally, since again by concentration of measure $\mathbb{P}(\|G_{N}\|_{\infty}\geq N)\leq e^{-cN^{2}}$, we conclude that for some $c^{\prime}(\delta)>0$ and all $N\in\mathbb{N}$
${\mathbb{P}(|\lambda_{1}(X_{N}(0))-x|\leq\delta)\leq 2N^{2}\mathbb{P}(\{|\lambda_{1}(X_{N}(0))-y|\leq\delta\})+4N^{2}e^{-c^{\prime}(\delta)N^{2}}.}$
The result follows as long as $\mathbb{P}(|\lambda_{1}(X_{N}(0))-x|\leq\delta)$ is much larger than $e^{-cN^{2}}$ which is clear since we assumed that lim $\limsup\frac{1}{N}\log\mathbb{P}(|\lambda_{1}(X_{N}(0))-x|\leq\delta)$ is finite.
Lemma B.2.
The functions $\lambda\mapsto F_{\nu,t}^{\pm}(\Lambda,\lambda)$ are decreasing on $(-\infty,{\ell}^{\Lambda}_{\nu,t})$ and increasing on $({\ell}^{\Lambda}_{\nu,t},\ell_{\nu,t})$. Moreover they vanish at ${\ell}^{\Lambda}_{\nu,t}$. These functions are bounded below by $c(\lambda-{\ell}^{\Lambda}_{\nu,t})^{2}$ for some constant $c>0$.
Proof.
The first point is a slight generalization of the previous proof where we take the supremum over vectors $\mathsf{M}\in\mathcal{B}^{N}_{\delta}(\alpha)$. Moreover, because $G_{N}\mapsto\lambda_{1}(B_{N}+G_{N})$ is Lipschitz with Lispchitz constant $1$, we deduce as in [AGZ10, Theorem 2.3.5] that since $G_{N}$ has Gaussian entries with variance $1/N$, for every $\delta>0$,
$\mathbb{P}(|\lambda_{1}(B_{N}+G_{N})-\mathbb{E}[\lambda_{1}(B_{N}+G_{N})]|\geq\varepsilon)\leq 2e^{-\frac{\varepsilon^{2}}{4}N}\,.$
By Theorem 1.1, we know that if $N_{i}/N$ goes to $\alpha_{i}$, $\mathbb{E}[\lambda_{1}(B_{N}+G_{N})]$ converges towards ${\ell}^{\Lambda}_{\nu,t}$. Note that this limit is continuous in $\nu$ provided $\Lambda<\ell_{\nu}$. As a consequence, for $\varepsilon>0$ and $\delta$ small enough so that
${\sup_{\|\alpha-\alpha^{\prime}\|_{\infty\leq\delta}}\Bigr{|}\ell^{\Lambda}_{\sum\alpha_{i}\delta_{\eta_{i}},t}-\ell^{\Lambda}_{\sum\alpha^{\prime}_{i}\delta_{\eta_{i}},t}\Bigr{|}\leq\varepsilon},$
we find that for $|x-\ell^{\Lambda}_{\nu,t}|\geq 2\varepsilon$,
$\sup_{\mathsf{M}\in\mathcal{B}_{\delta}^{N}}\mathbb{P}(|\lambda_{1}(X_{N})-x|\leq\varepsilon)\leq\sup_{\mathsf{M}\in\mathcal{B}_{\delta}^{N}}\mathbb{P}(|\lambda_{1}(X_{N})-\mathbb{E}[\lambda_{1}(X_{N})]|\geq|x-\ell^{\Lambda}_{\nu,t}|/2)\leq 2e^{-\frac{|x-\ell^{\Lambda}_{\nu,t}|^{2}}{4^{3}}N}.$
Since $\Lambda\mapsto\ell^{\Lambda}_{\nu,t}$ is continuous, it follows by first taking the limit when $N$ goes to infinity and then the supremum over $\Lambda$ in a small ball, that
${F^{+}_{\nu,t}(\Lambda,\lambda)\geq\frac{|\lambda-\ell^{\Lambda}_{\nu,t}|^{2}}{4^{3}}\,}.$
The same result holds for $F^{-}_{\nu}$. Finally, $F^{\pm}_{\nu,t}$ has to vanish somewhere. Indeed, by exponential tightness (2.2) we know that for some finite $M$, and $N$ large enough, we must have
$P(\lambda_{1}(X_{N})\in[-M,M])\geq 1/2\,.$
Moreover $M$ is chosen independently of the choice of $\alpha\in\Delta^{p}$ governing the distribution of the eigenvalues of $B_{N}$. Covering $[-M,M]$ with balls of width $\delta$ and using the previous monotonicity, we deduce that ${\inf F^{\pm}_{\nu,t}(\Lambda,\cdot)}$ equals 00, with the infimum achieved at ${\ell^{\Lambda}_{\nu,t}}$. ∎